### 16 Mar 2026 v1

# Hierarchies of direct powers, ultrapowers and cumulative powers

Abstract.
In this paper we investigate cumulative hierarchies of functions on structures, or cumulative powers, and study their properties. Particularly, we show how they extend the preservation phenomena of reduced powers, direct powers and ultrapowers by offering a characterization result of the fragment of first-order theory it preserves, and elucidate the connection between the three sorts of constructions. More precisely, we show how both direct powers and ultrapowers may be obtained from cumulative powers as a quotient by appropriate equivalence relations. We address how embeddability lifts from generating structures to their cumulative powers, direct powers and ultrapowers, and under what conditions ultrapowers embed into corresponding cumulative powers or direct powers. We further offer an application of the framework to show a straightforward way of constructing a surreal field.
1991 Mathematics Subject Classification:
03C20, 03C98, 03C95, 03C68.
1. Introduction
Functional constructions on mathematical structures play a major role in model theory in the form of direct powers and ultrapowers. These constructions are essentially non-cumulative. Despite there always being a trivial embedding from the generating structure into its direct power or ultrapower, there is nevertheless no direct interaction between the elements from the generating structure and the resulting ones.
Consider the collection of all functions from a given index set $U$ into the naturals, $\mathbb{N}^{I}$. There is a straightforward manner of extending any operation $\mathrm{F}$ on $\mathbb{N}$ to a $\mathrm{F}^{\prime}$ on $\mathbb{N}^{I}$ that behaves similarly to the original operation, as in the corresponding direct power – that is, pointwisely, so that $\mathrm{F}^{\prime}(f_{1},...,f_{n})=g$ iff for all $j\in I$, $\mathrm{F}\big(f_{1}(j),...,f_{n}(j)\big)=g(j)$. Likewise, we might define the constant function $\overline{0}$ as the identity element of the pointwise addition, and its image under the extended successor function is precisely the identity element of the extended multiplication. We may extrapolate this construction, and given a family of index sets $\{I_{n}\}_{n\in\omega}$, inductively defining $\mathbb{N}_{n}$ by setting $\mathbb{N}_{0}=\mathbb{N}$ and $\mathbb{N}_{n+1}=\mathbb{N}_{n}^{I_{n}}$, we may keep extending those operations and identity elements accordingly. Since there is a natural embedding between any two structures so defined, we may even take $\mathbb{N}_{\omega}$ to be the natural limit of embeddings, and given a proper class sized family of index sets $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$, keep transifinitely extending this hierarchy. Each of the $\mathbb{N}_{\beta}$ inherit nice properties of $\mathbb{N}$, and given the preservation results of direct powers by Alfred Horn [13], it may be checked that they are indeed semirings. We may do the same thing with $\mathbb{Z}$, and similarly get a hierarchy of rings $\mathbb{Z}_{\beta}$.
Suppose now we take a cumulative counterpart of the above structures. Starting with $\mathbb{N}_{0}=\mathbb{N}$, we take $\mathbb{N}_{\beta^{+}}=\mathbb{N}_{\beta}^{I_{\beta}}\cup\mathbb{N}_{\beta}$. Then, the operations and constants may no longer be pointwisely defined, but given a definition according to which each stage of the hierarchy is semantically conservative with respect to its predecessors, for limit ordinals $\lambda$ we may take $\mathbb{N}_{\lambda}$ to be the simple union of all the preceding stages. Call such construction a cumulative (direct) power. Is there a natural way of extending addition and multiplication to those structures? If so, what first-order properties are retained by cumulative powers?
The present paper has four different goals, interrelated as a broad study of cumulative powers and hierarchical constructions of functions. The first is to answer the above questions. To do so, in Section 2, we define a framework for working with hierarchies of cumulative powers of arbitrary structures. The main feature of the construction, offered in Definition 2.3, is a hereditary definition of the extended operations and relations of higher levels, so that they are reduced to the lower levels at which they are originally defined. The resulting structures have elements from different levels of such hierarchy, such as higher-order functions and non-functional elements, which nevertheless interact in a natural manner. We then show how a substantial fragment of first-order theory is preserved by the constructions. In fact, we offer the precise fragment of first-order theory preserved by cumulative powers. Call a sentence preserved by direct powers a direct power sentence. Let a formula be non-collapsible if its positive equalities do not force the collapse of different levels of the hierarchy of cumulative powers (as we show, a syntactic definition of non-collapsible formulas may be offered in Definition 2.17). Then:
Corollary 2.27.
Let $\varphi\in\mathcal{L}^{\sigma}$ be a sentence. Then:
  • (1)
    $\varphi$ is preserved by finite cumulative powers iff it is a non-collapsible direct power sentence;
  • (2)
    $\varphi$ is preserved by arbitrary cumulative powers iff it is a constant-free non-collapsible direct power sentence.
This extends well known results of preservation of first-order properties by reduced products by [4], [14], and [10], of direct products by [19], and the classical result for ultrapowers.
Given the construction of cumulative powers, the second aim of this paper naturally arises, which is how the new construction relates to direct powers and ultrapowers. As we argue in Section 3, the main difference between cumulative powers and direct powers is that in the latter equality is also defined hereditarily, whereas in the former, equality is defined as identity. This asymmetry ultimately accounts for the difference between the preservation results for direct powers and the one presented in Corollary 2.27. This fact is consolidated by showing, in Theorem 3.6, how quotienting a cumulative power of a structure by an equivalence relation, defined by an equality modulo hereditary identity, results in a structure which is isomorphic to its relative direct power.
In Section 4, we show a cumulative power may be further refined by quotienting it by yet another equivalence relation, given by an adequately defined ultrafilter, so that the resulting structure is isomorphic to a relative ultrapower, and therefore preserves any first-order property of a finite language. In this way, we may see how cumulative powers serve as a generalisation of both ultrapower and direct power constructions, as the choice of the equivalence relation which quotients it defines the first-order faithfulness to the original structure. The preservation results may be thus summarized as follows. Let $\mathcal{L}_{\mathcal{F}}$ be the fragment of first-order logic preserved by cumulative powers, $\mathcal{L}_{\Pi/F}$ the fragment preserved by reduced powers, $\mathcal{L}_{\Pi}$ the fragment preserved by direct powers, and $\mathcal{L}_{\Pi/\mathcal{U}}$ the fragment preserved by ultrapowers. Then, we have the following graph of containment of preserved first-order formulas:
$\mathcal{L}_{\mathcal{F}}$$\mathcal{L}_{\Pi/F}$$\mathcal{L}_{\Pi}$$\mathcal{L}_{\Pi/\mathcal{U}}\subseteq\subseteq\subseteq$
Cumulative powers, therefore, provide a structure from which we may take both direct powers and ultrapowers to arise by imposing appropriate equivalence relations, providing thus a natural hierarchy of preservation of first-order theory represented by the following flow chart. Let $I$ be an index set, $\mathcal{A}$ a structure, $\mathcal{F}_{1}(\mathcal{A})$ the cumulative power generated by them, and $\mathcal{U}$ an ultrafilter over $I$. Then:
[…]
The third aim of the paper is to study how embeddability is lifted from structures to its cumulative powers, and resulting quotient structures. For natural reasons, part of that question is inevitably reduced to criteria of embedding between ultrapowers. [^1] Since the matter between direct powers is straightforward, as in Theorem 3.8. Let $\leq_{\mathrm{RK}}$ denote the Rudin-Keisler ordering of ultrafilters [17]. A partial answer is then offered by the following result:
Theorem 4.20.
Let $u:\mathcal{A}\to\mathcal{B}$ be injective, $h:J\to I$ be a mapping, and $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters respectively over $I$ and $J$. Let also $H:X\mapsto h^{-1}[X]$ and define $w:A^{I}\to B^{J}$ such that, for each $f\in A^{I}$, $w(f)\in B^{J}$ is the function such that $\{j\in J\mid w(f)(j)=u(a)\}=H(\{j\in I\mid f(j)=a\})$. [^2] As we later note, $w(f)$ is indeed well-defined, that is, its description is unique. Define $e:\mathcal{A}^{I}/\mathcal{U}_{I}\to\mathcal{B}^{J}/\mathcal{U}_{J};[f]_{\mathcal{U}_{I}}\mapsto[w(f)]_{\mathcal{U}_{J}}$. Then $e$ is an embedding iff $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$.
Given the equivalence results between quotiented cumulative powers and ultrapowers, with Corollary 4.22 gives us the desired result.
In the same vein, we address the related question of when there is a fragment of the cumulative power which is isomorphic to a given relative ultrapower. Similarly to the previous case, the question reduces to a matter of the existence of an embedding from an ultrapower into its relative direct power. To present the answer, we need the following notion.
Given a signature $\sigma$, a $\sigma$-structure $\mathcal{A}$, a function $\mathrm{F}\in\sigma$, and $a\in A$, we define $S^{a\downarrow}_{\mathrm{F}}$ as the set of all tuples from which $a$ may be obtained by means of the function $\mathrm{F}^{\mathcal{A}}$; analogously, $S^{a\uparrow}_{\mathrm{F}}$ as the set of all the tuples containing $a$ for which $\mathrm{F}^{\mathcal{A}}$ is defined; and similarly, $S^{a}_{\mathrm{R}}$ as the set of all the tuples containing $a$ for which $\mathrm{R}^{\mathcal{A}}$ obtains. We then let $\tau_{a}$ be the sum of $|S^{a\downarrow}_{\mathrm{F}}|$, $|S^{a\uparrow}_{\mathrm{F}}|$ and $|S^{a}_{\mathrm{R}}|$ for all $\mathrm{F},\mathrm{R}\in\sigma$, and $\tau_{\mathcal{A}}$ be the greatest among the $\tau_{a}$ for all $a\in A$.
Theorem 4.31.
There is a choice function $e$ on the equivalence classes of $A^{I}/\mathcal{U}$ such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff $\mathcal{U}$ is $\tau_{\mathcal{A}^{I}/\mathcal{U}}^{+}$-complete.
A consequence of the above results is that, for $\mathbb{N}_{\mathrm{PA}}$ the naturals with the successor function as its only operation, as we show, it is always the case that $\mathbb{N}_{\mathrm{PA}}^{I}/\mathcal{U}$ embeds into $\mathbb{N}_{\mathrm{PA}}^{I}$, no matter the choice of index set and ultrafilter – where the non-trivial case is when $\mathcal{U}$ is free. However, a different implication is that for any infinite field $\mathcal{A}$, there is a choice function on the equivalence classes of $\mathcal{A}^{I}/\mathcal{U}$ which embeds it into $\mathcal{A}^{I}$ iff those structures are isomorphic.
As the last goal, we offer an application of the hierarchical constructions to real closed fields. We show that, at each ordinal stage of either a hierarchy generated by ultrapowers or quotiented cumulative powers, a suitable choice of ultrafilters produces hyperreal fields, whereas the hierarchy as whole, under the assumption of Global Choice – following results by Conway [5] and Ehrlich [7] –, forms a real closed field isomorphic to the field of surreal numbers.
This paper offers, therefore, cumulative function hierarchies as novel independent objects of study, with systematic analyses of their logical behavior and their relations with more usual functional constructions. The subject may be of interest for studying algebraic extensions of structures, as well as their model theoretic invariances. The background theory in which the paper is developed may be taken to be ZFC, as many proofs make use of Choice. However, some results make the explicit assumption of Global Choice, at which point they may be interpreted as having NBG as their background.
1.1. Conventions
Before we begin, we start with a few remarks on conventions we shall use. For $n\in\omega$, a set $A$ and $a\in A$, we write $\langle x_{1},...,a,...,x_{n-1}\rangle$ for a $n$-tuple of $A^{n}$ in which $a$ occurs in some position. We shall also follow the convention of denoting structures by a calligraphic font, and its corresponding domain, by an italicized font – for example, $\mathcal{A}$ and $A$, respectively. Let $\mathcal{A}$ be a structure and $I$ be a set. We denote by $\mathcal{A}^{I}$ the direct power of $A$ with index set $I$, and by $A^{I}$ its domain, that is, the set of all functions from $I$ into $A$. If $\mathcal{U}$ is an ultrafilter over $I$, we denote by $\mathcal{A}^{I}/\mathcal{U}$ the resulting ultrapower. By the natural embedding of a set or structure into its direct power or ultrapower, we mean the mapping which takes each element of the domain of the set and maps it to either the constant function which assigns to every element of the index set the element of the domain in question, in the case of direct powers, or its corresponding equivalence class modulo the ultrafilter, in the case of ultrapowers. We let $\mathbf{On}$ stand for the class of ordinals, $\alpha$, $\beta$, $\delta$, $\lambda$ and so on, stand for ordinals, $\kappa$, $\tau$, $\xi$, $\zeta$, and so on, for cardinals, $\varphi$, $\psi$, $\gamma$ and so on, for formulas, $\Delta$ for an arbitrary string of quantifiers, and $t$ and $s$ for arbitrary terms of a given language. For an ordinal $\beta$ or a cardinal $\kappa$, we denote by $\beta^{+}$ and $\kappa^{+}$ their respective successors.
We take the basic first-order language to be composed by individual variables, negation, conjunction, equality, and the universal quantifier (where the other connectives and quantifiers may be defined as usual). For an ordered set $\mathcal{A}$, we denote by $\mathrm{coi}(\mathcal{A})$ and $\mathrm{cof}(\mathcal{A})$ its coinitiality and cofinality, respectively. For a signature $\sigma$, we denote by $\mathcal{L}^{\sigma}$ the first-order language induced by it. As usual, we use a calligraphic letter $\mathcal{A}$ to denote a structure, and the respective italicized letter $A$ to denote its corresponding domain. Given a constant $\mathrm{c}$, function symbol $\mathrm{F}$ or relation symbol $\mathrm{R}$ of $\sigma$, and a $\sigma$-structure $\mathcal{A}$, we denote by $\mathrm{c}^{\mathcal{A}}$, $\mathrm{F}^{\mathcal{A}}$ and $\mathrm{R}^{\mathcal{A}}$ their interpretations in $\mathcal{A}$, respectively. Whenever we are working with some model, we shall state things like $\mathcal{A}\models\varphi[a_{1},...,a_{n}]$, for $a_{0},...,a_{n}\in A$, to mean $\varphi$ is a formula with at most $n$ free variables $x_{1},...,x_{n}$, and $\varphi$ is true in $\mathcal{A}$ when the $x_{i}$ are assigned the respective values $a_{i}$. Likewise, given a $\sigma$-term $t$, we shall denote by $t^{\mathcal{A}}(a_{1},...,a_{n})$ the element of $\mathcal{A}$ obtained by interpreting $t$ with each of its free variable $x_{i}$ being assigned the value $a_{i}$ – for example, in $\mathbb{N}$, $x[2]+y[3]$ is the element $2+^{\mathbb{N}}3=5$.
2. A hierarchy of cumulative powers
In this section we introduce the main objects of investigation: cumulative hierarchies of functions on algebraic structures. The idea is simple: we iterate the operation of taking all functions from an object to itself.
Definition 2.1.
Let $X$ be a designated set, and $\{I_{\alpha}\}_{\alpha<\mathbf{On}}$ be a family of index sets. We then define:
  • $\mathcal{F}_{0}(X)=X$;
  • $\mathcal{F}_{\beta^{+}}(X)=\big(\mathcal{F}_{\beta}(X)\big)^{I_{\beta}}\cup\mathcal{F}_{\beta}(X)$;
  • $\mathcal{F}_{\lambda}(X)=\bigcup_{\alpha<\lambda}\mathcal{F}_{\alpha}(X)$, for limit $\lambda$.
We let $\mathcal{F}_{\mathbf{On}}(X)=\bigcup_{\alpha\in\mathbf{On}}\mathcal{F}_{\alpha}(X)$ and call $\mathcal{F}_{\beta}(X)$ the $\beta$-th level of $\mathcal{F}_{\mathbf{On}}(X)$.
Notice, at limit levels of the construction, no new objects are created. We can immediately define a notion of level for the elements of the hierarchy.
Definition 2.2 (Level of an element).
Let $a\in\mathcal{F}_{\mathbf{On}}(X)$. Then, define the level $\rho$ of $a$ as $\rho(a)=\beta$ iff $a\in\mathcal{F}_{\beta}(X)\setminus\bigcup_{\alpha<\beta}\mathcal{F}_{\alpha}(X)$.
In other words, for $a\in\mathcal{F}_{\mathbf{On}}(X)$, $\rho(a)$ is the level at which $a$ is generated. For clarity of our presentation, we shall also define $\mathcal{F}_{\beta}^{*}(X)=\mathcal{F}_{\beta}(X)\setminus\bigcup_{\alpha<\beta}\mathcal{F}_{\alpha}(X)$. [^3] Notice $\mathcal{F}_{\beta}^{*}(X)=\varnothing$ iff either $\beta>0$ is a limit ordinal or $\beta=0$ and $X=\varnothing$. Notice $a\in\mathcal{F}_{\beta}^{*}(X)$ iff $a\in\mathcal{F}^{\beta}(X)$ and $\rho(a)=\beta$, so that we may use the former to abbreviate the latter.
When $X$ has a structure (i.e. functions, relations and constants defined on it) we want to extend them to the whole hierarchy generated by it.
Definition 2.3 (Cumulative power hierarchy).
Let $\sigma=\langle\{F_{i}\}_{i\in I},\{R_{i}\}_{i\in J},\{c_{i}\}_{i\in K},\mathrm{ar}\rangle$ be a signature. [^4] Where $\mathrm{ar}:(\{F_{i}\}_{i\in I}\cup\{R_{i}\}_{i\in J})\to\mathbb{N}$ is the arity function. Let also $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of sets, $\mathcal{A}=\langle A,\langle\mathrm{F}_{i}\rangle_{i\in I},\langle\mathrm{R}_{i}\rangle_{i\in J},\langle\mathrm{c}_{i}\rangle_{i\in K}\rangle$ be a $\sigma$-structure, where $A$ is the underlying set, $\mathrm{F}_{i}:A^{\mathrm{ar}(F_{i})}\to A$ a $\mathrm{ar}(F_{i})$-nary function, $\mathrm{R}_{i}\subseteq A^{\mathrm{ar}(R_{i})}$ a $\mathrm{ar}(R_{i})$-ary relation, and $\mathrm{c}_{i}\in A$ the interpretation of the constant $c_{i}$. The cumulative power hierarchy generated by $\mathcal{A}$ is composed by the following structures, for each $\beta\in\mathbf{On}$:
  • $\mathcal{F}_{0}(\mathcal{A})=\mathcal{A}$;
  • $\mathcal{F}_{\beta}(\mathcal{A})=\langle\mathcal{F}_{\beta}(A),\langle\mathrm{F}_{i,\beta}\rangle_{i\in I},\langle\mathrm{R}_{i,\beta}\rangle_{i\in J},\langle\mathbf{c}_{i,\beta}\rangle_{i\in K}\rangle$.
Where $\mathrm{F}_{i,\beta}:(\mathcal{F}_{\beta}(A))^{\mathrm{ar}(F_{i})}\to\mathcal{F}_{\beta}(A)$, $\mathrm{R}_{i,\beta}\subseteq(\mathcal{F}_{\beta}(A))^{\mathrm{ar}(R_{i})}$ and $\mathbf{c}_{i,\beta}\subseteq\mathcal{F}_{\beta}(A)$ are inductively defined in the following way. Let $a_{0},...,a_{n}\in\mathcal{F}_{\beta}(A)$ and $\delta=\mathrm{max}\{\rho(a_{i})\}_{i\leq n}$. For $\alpha\in\mathbf{On}$ and $j\in\bigcup_{\alpha\in\mathbf{On}}I_{\alpha}$, define the hereditary function $\vartheta:(\mathbf{On}\times\mathcal{F}_{\mathbf{On}}(A)\times\bigcup_{\alpha\in\mathbf{On}}I_{\alpha})\to\mathcal{F}_{\mathbf{On}}(A)$ in the following way:
$\vartheta^{\alpha}_{j}(a_{i})=\begin{cases}a_{i}(j),\ \text{if}\ \rho(a_{i})\geq\alpha\ \&\ j\in\mathrm{dom}(a_{i})\\ a_{i},\ \text{otherwise}\end{cases}$
Now, for a successor $\beta>0$, we have:
  • (i)
    The function $\mathrm{F}_{i,\beta}$: for $\delta=max\{\rho(a_{i})\}_{1\leq i\leq\mathrm{ar}(\mathrm{F}_{i})}>0$,
    $\mathrm{F}_{i,\beta}(a_{1},...,a_{\mathrm{ar}(\mathrm{F}_{i})})=b$ iff $\delta=\rho(b)$ and $\forall j\in I_{\delta-1}\Big(\mathrm{F}_{i,\delta-1}\big(\vartheta_{j}^{\delta}(a_{1}),...,\vartheta_{j}^{\delta}(a_{ar(F_{i})})\big)=\vartheta^{\delta}_{j}(b)\Big)$;
  • (ii)
    The relation $\mathrm{R}_{i,\beta}$: for $\delta=max\{\rho(a_{i})\}_{1\leq i\leq\mathrm{ar}(R_{i})}>0$,
    $\mathrm{R}_{i,\beta}a_{1}...a_{ar(R_{i})}$ iff $\forall j\in I_{\delta-1}\big(\mathrm{R}_{i,\delta-1}\vartheta^{\delta}_{j}(a_{1})...\vartheta^{\delta}_{j}(a_{\mathrm{ar}(\mathrm{R}_{i})})\big)$;
  • (iii)
    The set $\mathbf{c}_{i,\beta}$:
    • $a\in\mathbf{c}_{i,1}$ iff $\forall j\in I_{0}\big(\vartheta^{1}_{j}(a)\in\{\mathrm{c}_{i}\}\big)$;
    • $a\in\mathbf{c}_{i,\beta}$ iff $\forall j\in I_{\rho(a)-1}\big(\vartheta^{\rho(a)}_{j}(a)\in\mathbf{c}_{i,\rho(a)-1}\big)$.
If a $\mathcal{F}_{\beta}(\mathcal{A})$ is constructed in this way, we say it is generated by $\mathcal{A}$. We denote the cumulative power hierarchy generated in this manner by $\mathcal{F}_{\mathbf{On}}(\mathcal{A})=\bigcup_{\alpha\in\mathbf{On}}\mathcal{F}_{\alpha}(\mathcal{A})$.
For a set $A$, we use $\mathcal{F}_{\beta}(A)$ to denote both the set of functions defined in Definition 2.1, so that, in general, we shall use a calligraphic letter $\mathcal{A}$ for the structured set, and a regular $A$ otherwise. Furthermore, let $a\in\mathcal{F}_{\beta}(A)$. For a successor $\alpha>\rho(a)$, we denote the constant function $a:I_{\alpha-1}\to\mathcal{F}_{\alpha-1}(A);j\mapsto a$ by $\overline{a}^{\alpha}$. If the ordinal $\alpha$ is clear by context, we omit the superscript and write $\overline{a}$.
Despite the $\vartheta$ function being defined on proper classes, we do not need proper classes to define the structures above. For each $\beta\in\mathbf{On}$, the function $\vartheta\mathord{\upharpoonright}_{\beta}$ is not a proper class, and suffices for defining the structure $\mathcal{F}_{\beta}(\mathcal{A})$. Despite it not being explicit, the above definition implies that, for a limit $\lambda$, $\mathcal{F}_{\lambda}(A)$, $\mathrm{F}_{i,\lambda}$, $\mathrm{R}_{i,\lambda}$, and $\mathbf{c}_{i,\lambda}$ are just the union of the respective sets of the previous levels of the construction, so that it is implicitly defined that $\mathcal{F}_{\lambda}(\mathcal{A})=\bigcup_{\alpha<\lambda}\mathcal{F}_{\alpha}(\mathcal{A})$.
Notice the general scheme of the definition above: for a given atom $\varphi$ and elements $a_{1},...,a_{n}\in\mathcal{F}_{\beta}(A)$, $\varphi(a_{1},...,a_{n})$ is the case iff for any $j\in I_{\delta-1}$, $\varphi\big(\vartheta^{\delta}_{j}(a_{1}),...,\vartheta^{\delta}_{j}(a_{n})\big)$ holds, where $\delta=\linebreak\max\{\rho(a_{i})\}_{1\leq i\leq n}$. [^5] Although, in the case of the extended functions, one more constraint is imposed – and equality, of course, does not accord to that. Let us call such sort of definition a hereditary definition. We shall shortly see what may be said of structures defined hereditarily.
The main difference between the direct power of a structure and its respective cumulative power is the presence of elements from the generating structure itself in the latter. Given the hereditary definition of the relations and functions in cumulative powers, one might think this new construction may be reduced to direct powers with the addition of indiscernible elements, such as an element of the generating structure and its corresponding constant function. We may see at once that would be a mistake: given any index set, consider $\mathcal{F}_{1}(\mathbb{N})$. At first, it would seem that $2$ and the constant function $\overline{2}$ are indiscernible. However, we may see that while $\mathcal{F}_{1}(\mathbb{N})\models 2+5=7$, we have $\mathcal{F}_{1}(\mathbb{N})\not\models\overline{2}+5=7$. Similarly, $\mathcal{F}_{1}(\mathbb{N})\models\overline{2}+5=\overline{7}$, while $\mathcal{F}_{1}(\mathbb{N})\not\models 2+5=\overline{7}$.
For simplicity, henceforth we shall consider a signature with one $n$-ary function $F$, one $n$-ary relation $R$, and a single constant $c$, whenever that is harmless. It is straightforward how subsequent results apply to larger signatures.
Notice that if $\sigma$ contains constants, then $\mathcal{F}_{\beta}(\mathcal{A})$ so defined is no longer a $\sigma$-structure, as there is no adequate interpretation for the constants. In fact, each constant $\mathrm{c}$ of $\sigma$ has now a predicate associated to it, $\mathbf{c}_{\beta}$. However, for a finite initial segment of the cumulative power hierarchy, we may define:
Definition 2.4 ($\sigma$-structure of the cumulative power hierarchy and hereditary constant).
For $n<\omega$, let $\mathcal{A}$ be a $\sigma$-structure. For a constant $c$ of $\sigma$, its $n$-th hereditary constant (relative to $\mathcal{A}$) is the element $\mathrm{c}_{n}\in\mathcal{F}^{*}_{n}(\mathcal{A})$ recursively defined as $\mathrm{c}_{0}=\mathrm{c}^{I_{0}}$, and $\mathrm{c}_{n+1}=\overline{\mathrm{c}_{n}}$. That allows us to define the $\sigma$-structure $\mathcal{F}_{n}^{+}(\mathcal{A})=\langle\mathcal{F}_{n}(\mathcal{A}),\mathrm{c}_{n}\rangle$.
The reason for the restriction on the levels of the cumulative power hierarchy is that, for limit levels, we find no natural way of defining an interpretation for the constants of $\sigma$.
Observation 2.5.
It is straightforward to check that:
  • (1)
    for $\beta\in\mathbf{On}$, the extended structures $\mathcal{F}_{\beta}(\mathcal{A})$ are well-defined for two reasons: first, by assumption we have $a,a_{i},b\in\mathcal{F}_{\beta}(A)$, and therefore, in each case $\delta\leq\beta$, so that $\vartheta^{\delta}_{j}(a_{i})\in\bigcup_{\alpha<\beta}\mathcal{F}_{\alpha}(A)$; second, for any $c\in\mathcal{F}^{\mathbf{On}}(A)$, $\rho(c)$ is not a limit ordinal. Furthermore, the extended operations, functions, and relations for limit ordinals follow straightly from the above definition.
  • (2)
    The $\vartheta$ function defined above allows us to give a single definition for the extended operations and relations for every successor ordinal, but it can be made more explicit. For example, an equivalent definition of a binary function is (for $\delta=\mathrm{max}\{\rho(a),\rho(b)\}$):
[…]
    Notice that means the definition of $\mathrm{F}_{\beta}(a,b)$ when $\rho(a)\neq\rho(b)$ is such that $\mathrm{F}_{\beta}(a,b)=\mathrm{F}_{\gamma}(a,b)$, for every $\gamma\in[\delta,\beta]$. The intuitive idea in the definitions, represented by the function $\vartheta$, consists in going back to where the largest element is defined and use the operation there, which by assumption is already defined. In the definition of the relations, an equivalent definition would need to break down into the levels of each member of any $n$-tuple. For example, for a binary $\mathrm{R}$, an equivalent definition is:
[…]
    We may see the amount of clauses in a definition like that grows proportionally to the number of permutations of a $n$-tuple, so that using the $\vartheta$ function allows for a more economic definition. Likewise, the definition of the set defined by the constants may be easily calculated to be:
    • $\mathbf{c}_{i,1}=\mathrm{c}_{i}^{I_{0}}\cup\{\mathrm{c}_{i}\}$; [^6] That is, the collection of all functions from $I_{0}$ into $\{\mathrm{c}_{i}\}$ plus the singleton itself.
    • $\mathbf{c}_{i,\beta^{+}}=\mathbf{c}_{i,\beta}^{I_{\beta}}\cup\mathbf{c}_{i,\beta}$; [^7] Notice $\mathbf{c}_{i,\beta}$, for $\beta>0$, is a set.
    • $\mathbf{c}_{i,\lambda}=\bigcup_{\beta\in\lambda}\mathbf{c}_{i,\beta}$.
  • (3)
    for $\alpha<\beta$, $\mathrm{F}_{\beta}\mathord{\upharpoonright}_{\mathcal{F}_{\alpha}(A)}=\mathrm{F}_{\alpha}$. Each structure inherits the elements of the domain of all preceding structures, plus all of the functions with domain and codomain in those elements.
  • (4)
    for any $\beta\in\mathbf{On}$, $\rho\big(\mathrm{F}_{\beta}(a_{1},...,a_{n})\big)=max\{\rho(a_{1}),...,\rho(a_{n})\}$.
  • (5)
    for any $a\in\mathcal{F}^{*}_{\beta}(A)$ and $j\in\mathrm{dom}(a)$, $\rho\big(a(j)\big)<\rho(a)$.
  • (6)
    for any $\beta\in\mathbf{On}$, $j\in\bigcup_{\alpha\in\mathbf{On}}I_{\alpha}$ and $a\in\mathcal{F}_{\beta}(A)$, $\rho\big(\vartheta^{\beta}_{j}(a)\big)\leq\rho(a)$.
Proposition 2.6.
For $\delta=max\{\rho(a_{i})\}_{i\leq n}$ and $\alpha\leq\delta\leq\beta$ :
  • (1)
    if $j\in I_{\delta}$, then $\vartheta^{\alpha}_{j}\big(\mathrm{F}_{\beta}(a_{0},...,a_{n})\big)=\big(\mathrm{F}_{\beta}(a_{0},...,a_{n})\big)(j)=\mathrm{F}_{\beta}\big(\vartheta^{\alpha}_{j}(a_{0}),...,\vartheta^{\alpha}_{j}(a_{n})\big)$;
  • (2)
    if $j\not\in I_{\delta}$, then $\vartheta^{\delta}_{j}\big(\mathrm{F}_{\beta}(a_{0},...,a_{n})\big)=\mathrm{F}_{\beta}(a_{0},...,a_{n})$.
Proof.
Straightforward by Definition 2.3.888We offer a quick argument for a binary function: by Observation 2.5, $\rho\big(\mathrm{F}_{\beta}(a,b)\big)=\delta$. Thus, $\vartheta^{\alpha}_{j}\big(\mathrm{F}_{\beta}(a,b)\big)=\big(\mathrm{F}_{\beta}(a,b)\big)(j)$. Without loss of generality, we may suppose $\delta=\rho(a)$. Then, $\vartheta^{\alpha}_{j}(a)=a(j)$. If $\rho(b)<\rho(a)$, then $\vartheta^{\alpha}_{j}(b)=b$, and so $\vartheta^{\alpha}_{j}\big(\mathrm{F}_{\beta}(a,b)\big)=\mathrm{F}_{\beta}\big(a(j),b\big)=\mathrm{F}_{\beta}\big(\vartheta^{\alpha}_{j}(a),\vartheta^{\alpha}_{j}(b)\big)$. If now $\rho(a)=\rho(b)$, then $\vartheta^{\alpha}_{j}(b)=b(j)$, so that $\vartheta^{\alpha}_{j}\big(\mathrm{F}_{\beta}(a,b)\big)=\big(\mathrm{F}_{\beta}(a,b)\big)(j)=\mathrm{F}_{\beta}\big(a(j),b(j)\big)=\mathrm{F}^{\beta}\big(\vartheta^{\alpha}_{j}(a),\vartheta^{\alpha}_{j}(b)\big)$.
Corollary 2.7.
For any $\beta$ and $j\in I_{\beta}$, $\vartheta^{\beta^{+}}_{j}:\mathcal{F}_{\beta^{+}}(A)\to\mathcal{F}_{\beta}(A)$ is an endomorphism (but not an automorphism).
Proposition 2.8.
If $\mathcal{A}$ is closed under $\mathrm{F}$, then $\mathcal{F}_{\beta}(\mathcal{A})$ is closed under $\mathrm{F}_{\beta}$.
Proof.
By induction on $\beta$. The base case is trivial, and for the inductive step it suffices to notice the definition of $\mathrm{F}_{\beta}$ works iff the $\mathrm{F}_{\alpha}$ are defined for all $\alpha<\beta$. The limit case is straightforward, as two elements in the limit case must be members of some iteration at which the operation is defined for them. ∎
2.1. A concrete example
In order to make the construction of the cumulative power hierarchy clearer, we may consider as a concrete case the one generated by the natural numbers.
Definition 2.9 (The cumulative power hierarchy of $\mathbb{N}$).
For $\beta\in\mathbf{On}$ and $\{\mathcal{F}_{\alpha}(A)\}_{\alpha<\mathbf{On}}$ itself as the family of index sets, we let
  • (i)
    $\mathcal{F}_{0}(\mathbb{N})=\mathbb{N}$,
  • (ii)
    $\mathcal{F}_{\beta}(\mathbb{N})=\langle\mathcal{F}_{\beta}(\mathbb{N}),+_{\beta},\cdot_{\beta},\mathrm{s}_{\beta},\mathbf{0}_{\beta}\rangle$,
  • (iii)
    $\mathcal{F}_{\lambda}(\mathbb{N})=\bigcup_{\alpha<\lambda}\mathcal{F}_{\alpha}(\mathbb{N})$, for limit $\lambda$,
  • (iv)
    $\mathcal{F}_{\mathbf{On}}(\mathbb{N})=\bigcup_{\alpha\in\mathbf{On}}\mathcal{F}_{\alpha}(\mathbb{N})$.
To illustrate the definition, let us consider a concrete calculation in $\mathcal{F}_{2}(\mathbb{N})$, whose domain is
[…]
For $f\in\mathcal{F}^{*}_{1}(\mathbb{N})=\mathbb{N}^{\mathbb{N}}$, let $f^{+}$ be defined as $\forall j\in\mathbb{N}\big(g^{+}(j)=\mathrm{s}(g(j))\big)$. Consider the function $\mathrm{s}_{1}$, the extension of the successor function $\mathrm{s}$ to $\mathcal{F}_{1}(\mathbb{N})$:
$\mathrm{s}_{1}(j)=\begin{cases}j^{+},\ \text{if}\ j\in\mathcal{F}^{*}_{1}(\mathbb{N})\\ \mathrm{s}(j)=j+1,\ \text{if}\ j\in\mathbb{N}\end{cases}$
[…]
$\forall j\in\mathcal{F}_{1}(\mathbb{N})\big(a(j)=f(j)+_{1}3\big)$,
which means that
$a(j)=\begin{cases}j^{+}+_{1}3,\ \text{if}\ j\in\mathbb{N}^{\mathbb{N}}\\ \mathrm{s}(1)+3,\ \text{if}\ j\in\mathbb{N}\end{cases}$
Continuing to calculate, we obtain that $j^{+}+_{1}3=b\in\mathbb{N}^{\mathbb{N}}$, where $b$ is such that
$\forall k\in\mathbb{N}\big(b(k)=j^{+}(k)+3\big)$.
Applying now the definition of $\mathrm{s}_{1}$ we obtain that, for $k\in\mathbb{N}$, we have that
$b(k)=j^{+}(k)+3=\mathrm{s}\big(j(k)\big)+3$.
In order to have a visual representation of the this sum we may take an order of $\mathcal{F}_{1}(\mathbb{N})$ such that $\mathbb{N}$ is the initial segment, and hence we may represent an element of $\mathcal{F}_{2}(\mathbb{N})$ by a column vector with the $j$-th coordinate being its value at $j$. Then the first vector below represents $\mathrm{s}_{1}$, and the sum $\mathrm{s}_{1}+_{2}3$ can be depicted as follows.
$\begin{bmatrix}1\\ 2\\ 3\\ \vdots\\ [1,1,5,...]\\ \vdots\\ [1,3,7,...]\\ \vdots\end{bmatrix}+3=$ $\begin{bmatrix}1+3\\ 2+3\\ 3+3\\ \vdots\\ [1,1,5,...]+3\\ \vdots\\ [1,3,7,...]+3\\ \vdots\end{bmatrix}$= $\begin{bmatrix}1+3\\ 2+3\\ 3+3\\ \vdots\\ [1+3,1+3,5+3,...]\\ \vdots\\ [1+3,3+3,7+3,$...]\\ $\vdots\end{bmatrix}$= $\begin{bmatrix}4\\ 5\\ 6\\ \vdots\\ [4,4,8,...]\\ \vdots\\ [4,6,10,$...]\\ $\vdots\end{bmatrix}$
2.2. Preservation and properties of cumulative powers
If one of the index sets defining a cumulative power hierarchy is either empty or a singleton, then the structure generated in its corresponding level is isomorphic to the structure of its preceding level, and therefore uninteresting. Thus, for the rest of the paper, we shall assume any index set to be of cardinality greater than $1$.
The natural question which follows the introduction of cumulative powers amounts to what sentences of the language induced by $\sigma$ are preserved by the construction. For that, some definitions are in order.
As introduced in [13] (though not called as such there), a Horn formula is a formula in prenex conjunctive normal form whose conjuncts are disjunctions of literals of which at most one is a non-negated atom. In other words, a Horn formula is of the form
$\Delta(\bigwedge_{k\leq m}(\bigvee_{i\leq n_{k}}\psi_{k,i})\big)$,
where each $\psi_{k,i}$ is a literal, and for each $k\leq m$ there is at most one $i\leq n_{k}$ such that $\psi_{k,i}$ is a non-negated atom. A Horn sentence is a closed Horn formula.
In [13], Horn showed that any Horn sentence is preserved by direct products. In [4], Chang and Morel showed Horn sentences are also preserved by reduced products produced by the Frechet filter, and as expressed in [8], Chang further generalized the result, showing that, in fact, any proper (that is, defined by proper filters) reduced product preserves Horn sentences.
A class of structures defined by a set of sentences is a elementary class, and if all those sentences are Horn sentences, call it a Horn class. The results above mentioned therefore say that any Horn class is closed by proper reduced (and thus direct) products. Chang and Morel showed in [4] that not every elementary class closed under direct products is a Horn class, but conjectured that any elementary class closed under reduced products is indeed a Horn class. In [14], Keisler proved the conjecture under the assumption the continuum hypothesis holds for at least one cardinal, whereas in [10], Galvin strengthened the result by eliminating that assumption. Briefly:
Proposition 2.10 ([4], [14], and [10]).
A sentence is preserved by proper reduced products iff it is equivalent to a Horn sentence.
Proposition 2.11 ([14], and [10]).
A sentence is preserved by proper reduced powers iff it is equivalent to a disjunction of Horn sentence.
Let the occurrence of an atom in a formula be positive if in the prenex conjunctive normal form of that formula the atom is preceded by an even number of negations, and let it be negative otherwise. As such, one might suspect the fragment preserved by cumulative powers is the class of Horn sentences without positive equalities, for, as we have seen, there is an imbalance in how the interpretation of constants, functions and other relations are defined in cumulative powers (that is, hereditarily), and how equality is defined (that is, as true identity, rather than hereditary identity). In that case, cumulative powers would preserve strictly less first-order sentences than reduced products. However, consider the sentence $\forall x\forall y\exists z(x+y=z)$. As one may check on our previous example of a hierarchy given by the natural numbers, that sentence is preserved and has a positive occurrence of equality. Therefore, the matter is not as simple. Furthermore, as we may comparatively see on the definitions of cumulative powers and direct powers, the former are much closer to latter than to reduced products or reduced powers. Truth of formulas for functional elements of cumulative powers accord to how their truth are defined in direct powers. Therefore, if we want to characterize the fragment of first-order theory preserved by cumulative powers, our route should be to look at what fragment is preserved by direct powers.
In [19], Weinstein showed it is possible to give a syntactic – although not simple – characterization of that fragment. To expose it, some definitions are needed.
For $\varphi,\psi,\gamma\in\mathcal{L}^{\sigma}$, we write $\varphi\times\psi\Rightarrow\gamma$ when, for any $\sigma$-structures $\mathcal{A}$ and $\mathcal{B}$, $a_{0},...,a_{n}\in A$, and $b_{0},...,b_{n}\in B$, if $\mathcal{A}\models\varphi(a_{0},...,a_{n})$ and $\mathcal{B}\models\psi(b_{0},...,b_{n})$, then $\mathcal{A}\times\mathcal{B}\models\gamma(\langle a_{0},b_{0}\rangle,...,\langle a_{n},b_{n}\rangle)$. For $\Phi\subseteq\mathcal{L}^{\sigma}$, we let:
  • $\neg\Phi=\{\neg\varphi\mid\varphi\in\Phi\}$;
  • $\exists x\Phi=\{\exists x\varphi\mid\varphi\in\Phi\}$ (for the particular variable $x$);
  • $\forall x\Phi=\{\forall x\varphi\mid\varphi\in\Phi\}$;
  • $\bigvee\Phi=\{\bigvee_{i\leq n}\varphi_{i}\mid n<\omega\ \&\ \varphi_{i}\in\Phi\}$ (including the empty disjunction $\top$);
  • $\bigwedge\Phi=\{\bigwedge_{i\leq n}\varphi_{i}\mid n<\omega\ \&\ \varphi_{i}\in\Phi\}$ (including the empty conjunction $\bot$);
  • $\bigwedge^{*}\Phi=\{\bigwedge_{i\leq n}\varphi_{i}\mid n<\omega\ \&\ \varphi_{i}\in\Phi\ \&\ \bigwedge_{i\leq n}\varphi_{i}\ \text{is consistent}\}$.
In the cases of $\bigvee\Phi$ and $\bigwedge\Phi$, we make the assumption that there is a convention for listing the formulas of $\Phi$ to avoid redundancy – that is, the members of those sets are as defined modulo logical equivalence. In that way, when $\Phi$ is finite, $\bigvee\Phi$ and $\bigwedge\Phi$ should also be finite, and the mappings $\Phi\mapsto\bigvee\Phi$ and $\Phi\mapsto\bigwedge\Phi$ should be computable. Thus, we may let $\bigwedge^{*}\Phi$ be such that $\bigwedge^{*}\Phi\subseteq\bigwedge\Phi$. [^9] Even though there may not be a computable way to determine $\bigwedge^{*}\Phi$ as a subset of $\bigwedge\Phi$.
Let $\langle x_{n}\rangle_{n<\omega}$ be an enumeration of the variables of $\mathcal{L}^{\sigma}$. Given a finite set of atomic formulas $\Phi$, we define the following finite sets of atomic formulas:
  • $\Phi_{0}=\bigvee\bigwedge(\Phi\cup\neg\Phi)$,
  • $\Phi_{n+1}=\bigvee\bigwedge\exists x_{n}\Phi_{n}$, when $n$ is even,
  • $\Phi_{n+1}=\bigvee\bigwedge\forall x_{n}\Phi_{n}$, when $n$ is odd,
  • $\Phi^{*}_{0}=\bigvee\bigwedge^{*}(\Phi\cup\neg\Phi)$,
  • $\Phi^{*}_{n+1}=\bigvee\bigwedge^{*}\exists x_{n}\Phi^{*}_{n}$, when $n$ is even,
  • $\Phi^{*}_{n+1}=\bigvee\bigwedge^{*}\forall x_{n}\Phi^{*}_{n}$, when $n$ is odd.
Notice $\Phi^{*}_{n}\subseteq\Phi_{n}$ for every $n<\omega$, and that every formula in $\mathcal{L}^{\sigma}$ is equivalent to a formula in $\Phi^{*}_{n}$ for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$, equivalent to a formula in $\Phi_{n}$ for some $\Phi$ and $n$.
Proposition 2.12 ([19]).
Let $\Phi$ be a finite autonomous set of formulas closed under $\bigwedge$ and $\bigvee$ up to equivalence. Let $x$ be any variable and let $\varphi$, $\varphi^{\prime}$, and $\psi$ be formulas in $\bigvee\bigwedge\exists xF$. Fix $Q\in\{\exists,\forall\}$. Suppose for every disjunct $\varphi^{\prime}$ of $\varphi$ and every disjunct $\psi^{\prime}$ of $\psi$, there is a disjunct $\gamma^{\prime}$ of $\gamma$ such that for every conjunct $Qx\gamma^{\prime\prime}$ of $\gamma^{\prime}$, there are conjuncts $Qx\varphi^{\prime\prime}$ of $\varphi^{\prime}$ and $qx\psi^{\prime\prime}$ of $\psi^{\prime}$ such that $\varphi^{\prime\prime}\times\psi^{\prime\prime}\Rightarrow\gamma^{\prime\prime}$. Then, $\varphi\times\psi\Rightarrow\gamma$. Moreover, if $\varphi$ and $\psi$ are elements of $\bigvee\bigwedge^{*}Qx\Phi$, then the supposition is necessary and sufficient for $\varphi\times\psi\Rightarrow\gamma$.
Given the above proposition, Weinstein defines the primitive recursive predicate $R\subseteq(\mathcal{L}^{\sigma})^{3}$ such that $R(\varphi,\psi,\gamma)$ iff there is $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$ such that $\varphi,\psi,\gamma\in\Phi_{n}$ and:
  • if $n=0$, then $\varphi\times\psi\Rightarrow\gamma$;
  • if $n$ is odd, then for every disjunct $\varphi^{\prime}$ of $\varphi$ and every disjunct $\psi^{\prime}$ of $\psi$, there is a disjunct $\gamma^{\prime}$ of $\gamma$ such that for every conjunct $\exists x_{n-1}\gamma^{\prime\prime}$ of $\gamma^{\prime}$ there are conjuncts $\exists x_{n-1}\varphi^{\prime\prime}$ of $\varphi^{\prime}$ and $\exists x_{n-1}\psi^{\prime\prime}$ of $\psi^{\prime}$ such that $R(\varphi^{\prime\prime},\psi^{\prime\prime},\gamma^{\prime\prime})$ holds;
  • if $n>0$ is even, then for every disjunct $\varphi^{\prime}$ of $\varphi$ and every disjunct $\psi^{\prime}$ of $\psi$, there is a disjunct $\gamma^{\prime}$ of $\gamma$ such that for every conjunct $\forall x_{n-1}\gamma^{\prime\prime}$ of $\gamma^{\prime}$ there are conjuncts $\forall x_{n-1}\varphi^{\prime\prime}$ of $\varphi^{\prime}$ and $\forall x_{n-1}\psi^{\prime\prime}$ of $\psi^{\prime}$ such that $R(\varphi^{\prime\prime},\psi^{\prime\prime},\gamma^{\prime\prime})$ holds.
Notice $R$ is a computable predicate, as, given formulas $\varphi$, $\psi$ and $\gamma$, we can take $\Phi$ to be the set of atomic formulas occurring in $\varphi$, $\psi$ and $\gamma$, and it suffices to check for $n$ that is as large as the greatest between the quantifier ranks of those formulas.
Proposition 2.12 entails the following:
Proposition 2.13 ([19]).
$R(x,y,z)$ satisfies the following:
  • (1)
    for any $\varphi$, $\psi$, and $\gamma$ in some $\Phi_{n}$, if $R(\varphi,\psi,\gamma)$ holds, then $\varphi\times\psi\Rightarrow\gamma$;
  • (2)
    for any $\varphi$, $\psi$, and $\gamma$ in some $\Phi^{*}_{n}$, if $\varphi\times\psi\Rightarrow\gamma$, then $R(\varphi,\psi,\gamma)$ holds.
Finally, it was thus proved that:
Corollary 2.14 ([19]).
A sentence $\varphi\in\mathcal{L}^{\sigma}$ is preserved by direct products iff it is logically equivalent to a sentence $\psi\in\mathcal{L}^{\sigma}$ such that, for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$, $\psi\in\Phi_{n}$ and $R(\psi,\psi,\psi)$.
Notice we may replace products with powers in the above corollary, that is:
Corollary 2.15.
A sentence $\varphi\in\mathcal{L}^{\sigma}$ is preserved by direct powers iff it is logically equivalent to a sentence $\psi\in\mathcal{L}^{\sigma}$ such that, for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$, $\psi\in\Phi_{n}$ and $R(\psi,\psi,\psi)$ (iff it is preserved by direct products).
Proof.
The left to right direction is straightforward. For the other, if $\varphi$ is preserved by direct powers, then particularly $\varphi\times\varphi\Rightarrow\varphi$. But since, as we have seen, every sentence of $\mathcal{L}^{\sigma}$ is equivalent to some sentence in $\Phi^{*}_{n}$ for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$, there is $\psi$ which is logically equivalent to $\varphi$ and $\psi\in\Phi_{n}$ for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$. That also means $\psi\times\psi\Rightarrow\psi$. By Proposition 2.13 (2), we thus obtain $R(\psi,\psi,\psi)$. ∎
Therefore, Weinstein also proved a characterization of the class of sentences preserved by direct powers. Henceforth, we shall call a sentence which is preserved by direct powers a direct power sentence.
To frame the sorts of formulas with positive occurrences of equality (modulo their prenex conjunctive normal form), we need some definitions which are similarly non-intuitive.
Definition 2.16 (Non-collapsible positive equality).
We call the occurrence of a positive equality $\varphi$ in a formula in prenex conjunctive normal form non-collapsible if either:
  • (i)
    $\varphi$ has no variables on one of its sides;
  • (ii)
    or each and every universally quantified or free variable occurring on one side of $\varphi$ also occurs on the other;
  • (iii)
    or each of the sides of $\varphi$ either has variables bound by existential quantifiers, or has the occurrence of a constant.
If a positive equality in a formula in prenex conjunctive normal form is not non-collapsible, we call it collapsible.
Notice, with the exception of when condition (i) above is satisfied, a positive equality’s being non-collapsible is determined by the whole context of the formula in which it occurs (that is, by the quantifiers binding it), rather than just depending on its form alone. For the next results, for a formula $\varphi\in\mathcal{L}^{\sigma}$, let
$\mathrm{noncoll}(\varphi)=\{\psi\mid\psi\ \text{is a non-collapsible positive equality of}\ \varphi\}$.
Definition 2.17 (Non-collapsible formula).
We call a formula non-collapsible if it is in prenex conjunctive normal form $\varphi=\Delta\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}$ and either:
  • (i)
    $\varphi$ has no positive occurrences of equality;
  • (ii)
    or each of the positive equalities in $\varphi$ is non-collapsible;
  • (iii)
    or for each of its conjuncts $\bigvee_{l\leq p_{i}}\psi_{i,l}$, for any $t=s\in\mathrm{noncoll}(\Delta\bigvee_{i\leq p_{i}}\psi_{i,l})$ there is $\gamma\in\linebreak\bigvee\big(\{\psi_{i,l}\}_{l\leq p_{i}}\setminus\mathrm{noncoll}(\Delta\bigvee_{l\leq p_{i}}\psi_{i,l})\big)\setminus\{\top\}$ (that is, some non-empty subdisjunction of $\bigvee_{i\leq p_{i}}\psi_{i,l}$ which does not contain collapsible positive equalities of $\varphi$) such that $\vdash t=s\rightarrow\gamma$.
We may loosen our terminology and call an arbitrary formula non-collapsible if its prenex conjunctive normal form is non-collapsible. We call a formula as such because they do not force the collapse of different levels of the hierarchy, that is, they are formulas whose truth do not require the identification of elements of different levels of the hierarchy. A more detailed explanation of the above definitions are in order. To offer it, we shall consider the context of a cumulative power of first level, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})$. As one may easily see, a literal which is not a positive equality is preserved in both directions between direct powers and their relative cumulative powers. Thus, the satisfaction of condition (i) of Definition 2.17 by a formula entails its preservation. Similarly, the satisfaction of condition (iii) of Definition 2.17 ensures that, in each conjunct of a formula in prenex conjunctive normal form, there is a subformula not containing positive equalities which is satisfied, so that we also obtain preservation. On the other hand, if a sentence satisfies condition (ii) Definition 2.17, there are three possible cases. If it satisfies condition (iii) of Definition 2.16, the presence of an existential quantifier on each side of the equality allows us to choose, for whatever choice of elements of $\mathcal{F}_{\beta}(\mathcal{A})$ on the other side, a suitable element of the same level as those, so that identity may be reached. A constant, on the other hand, does a similar job, since it forces the term in which occurs to be of the largest possible rank inside the cumulative power structure in question. If it satisfies condition (ii) of Definition 2.16, since if there is a universally bound variable on one side, then it also occurs on the other side, the ranks of the interpreted terms will always match, for they will be forced. it is possible to choose suitable elements of $\mathcal{F}_{\beta}(\mathcal{A})$ given any elements picked by the universal quantifiers. The constants, in turn, play a similar role, since they are of the highest rank possible in the relevant model, and similarly their values are of the highest possible rank which can be assumed. Thus, if constants are on one side of a positive equality and there are variables bound by existential quantifiers on the other side, we may match the ranks as required.
We may see that cumulative powers do not, in general, preserve formulas containing positive occurrences of equality such that the existential quantifiers occur before the universal quantifiers bounding its variables. In fact, if universal quantifiers bound all of the variables of one side of equality, then such a sentence is not in general satisfied, because the levels of the chosen elements may not match. Consider for instance $\exists x_{1}...x_{n}\forall y_{1}...y_{m}\big(t(x_{1},...,x_{n})=s(y_{1},...,y_{m})\big)$, a formula which is not non-collapsible – that is, collapsible. Since $\rho\big(t^{\mathcal{F}_{1}(\mathcal{A})}(x_{1},...,x_{n})\big)=\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}$ and $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(y_{1},...,y_{m})\big)=\mathrm{max}\{\rho(y_{i})\}_{1\leq i\leq m}$, and since the universal quantifiers may pick elements in either $A$ or $\mathcal{F}^{*}_{1}(A)$, for a choice $b_{1},...,b_{m}\in A$ we would have $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=0$, which would require $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=0$. However, if $b_{1}\in\mathcal{F}^{*}_{1}(A)$, then $\rho\big(s^{\mathcal{F}_{1}(\mathcal{A})}(b_{1},...,b_{m})\big)=1$, which would mean $\mathrm{max}\{\rho(x_{i})\}_{1\leq i\leq n}=1$. Thus, there are no elements $a_{1},...,a_{n}\in\mathcal{F}_{1}(A)$ satisfying $\forall y_{1}...y_{m}\big(t(a_{1},...,a_{n})=s(y_{1},...,y_{m})\big)$ in $\mathcal{F}_{1}(\mathcal{A})$.
Lemma 2.18.
Let $\varphi\in\mathcal{L}^{\sigma}$ be in prenex conjunctive normal form. Then:
  • (1)
    if $\exists x\varphi$ satisfies condition (ii) of Definition 2.17 because it satisfies either condition (i) or (ii) of Definition 2.16, or if it satisfies conditions (i) or (iii) of Definition 2.17, then $\varphi$ is non-collapsible;
  • (2)
    if $\forall x\varphi$ is non-collapsible, then $\varphi$ is non-collapsible;
  • (3)
    if $\exists x\varphi$ is collapsible, then $\varphi$ is collapsible;
  • (4)
    if $\forall x\varphi$ is collapsible, then $\varphi$ is collapsible.
Proof.
(1) Since $\exists x\varphi$ satisfies condition (i), it has no positive equalities. Taking away a quantifier does not change that fact, and so $\varphi$ still satisfies condition (i).
(2) If condition (i), (ii) or (iii) of Definition 2.16 is satisfied, taking away a universal quantifier does not change their satisfaction. If condition (i) or (iii) of Definition 2.17 is satisfied, the conclusion is also trivial.
(3) and (4) If $\forall x\varphi$ is collapsible, by condition (i), it has positive occurrences of equality. By condition (iii), it contains a positive equality $t=s$, and for any $\gamma\in\bigvee\big(\{\psi_{i}\}_{i\leq k}\setminus\mathrm{noncoll}(\bigvee_{i\leq k}\psi_{i})\big)\setminus\{\top\}$ (where $\bigvee\big(\{\psi_{i}\}_{i\leq k}$ is the disjunction in which $t=s$ occurs), $\not\vdash t=s\rightarrow\gamma$. By condition (ii), each side has occurring variables; its sides do not share at least one universally bound or free variable; there are existentially bound variables on at most one of its sides; and the side that does not have existentially bound variables also has no occurrence of constants (for otherwise it would satisfy condition (iii) of Definition 2.16, and thus condition (ii) of Definition 2.17). Since none of these characteristics are altered by removing a quantifier from $\forall x\varphi$ or $\exists x\varphi$, $\varphi$ is also collapsible. ∎
Lemma 2.19.
Let $\mathcal{F}_{1}^{+}(\mathcal{A}$ be generated by the index set $I$, $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex disjunctive normal form. Then,
if $\forall j\in I\big(\mathcal{A}\models\varphi\big[\vartheta^{1}_{j}(a_{0}),...,\vartheta^{1}_{j}(a_{n})\big]\big)$, then $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n}]$.
Proof.
We show the contrapositive by an induction on the number of quantifiers. Let $\varphi[a_{0},...,a_{n}]=\bigvee_{i\leq k}\bigwedge_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$ contain no quantifiers. If $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\varphi[a_{0},...,a_{n}]$, that is, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\neg\psi_{i,l}[a_{0},...,a_{n}]$, then for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. If $\psi^{\prime}_{l,q}[a_{0},...,a_{n}]$ is a negated atom, we might observe, by our definitions, that means
$\forall j\in I\big(\mathcal{A}\not\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{\delta}_{j}(a_{0}),...,\vartheta^{\delta}_{j}(a_{n})]\big)$.
On the other hand, if $\psi^{\prime}_{l,q}[a_{0},...,a_{n}]$ is an atom, we may see, by our definitions, that means
$\exists j\in I\big(\mathcal{A}\not\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{\delta}_{j}(a_{0}),...,\vartheta^{\delta}_{j}(a_{n})]\big)$.
By the arbitrariness of $l^{\prime}$ and $i^{\prime}$, we have
$\exists j\in I\big(\mathcal{A}\not\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i^{\prime},l}[\vartheta^{\delta}_{j}(a_{0}),...,\vartheta^{\delta}_{j}(a_{n})]\big)$.
The rest follows by induction on the number of quantifiers. ∎
For the following results, we need the following notion. For each index set $I$ and $\sigma$-structure $\mathcal{A}$, we define the mapping $\vartheta^{-1}:\mathcal{F}_{1}(A)\to A^{I}$ such that, for each $a\in\mathcal{F}_{1}(A)$,
$\vartheta^{-1}(a)=\begin{cases}a,\ \text{if}\ \rho(a)=1\\ \overline{a},\ \text{otherwise (that is, if}\ \rho(a)=0\text{)}\end{cases}$
Lemma 2.20.
Let $\mathcal{F}^{+}_{1}(\mathcal{A})$ be generated by the index set $I$. For any literal $\varphi\in\mathcal{L}^{\sigma}$ which is not a positive equality and $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$,
$\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n}]$ iff $\mathcal{A}^{I}\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$.
Proof.
Straightforward from the definitions of cumulative power, direct power, and the $\vartheta^{-1}$ mapping. ∎
Lemma 2.21.
Let $\mathcal{F}^{+}_{1}(\mathcal{A})$ be generated by the index set $I$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex conjunctive normal form. If $\varphi$ is non-collapsible, then for any $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$,
if $\mathcal{A}^{I}\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n}]$.
Proof.
We shall consider separately each condition of Definition 2.17.
Suppose $\varphi$ satisfies condition (i). We proceed with an induction on the complexity of $\varphi$. If $\varphi$ is atomic, then it cannot be an equality, so the case is covered in both directions by Lemma 2.20. For the remaining cases, the non-trivial one is the existential. If $\mathcal{A}^{I}\models\forall x\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then for any $b\in A^{I}$, $\mathcal{F}_{1}^{+}(\mathcal{A})\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),b]$. Since $\vartheta^{-1}$ is surjective, that means for any $c\in\mathcal{F}_{1}(A)$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n}),\vartheta^{-1}(c)]$. Thus, by Lemma 2.18, $\varphi(x_{1},...,x_{n},y)$ is non-collapsible, so by the inductive hypothesis, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi[a_{0},...,a_{n},c]$. By the arbitrariness of $c$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\varphi[a_{0},...,a_{n}]$.
Suppose $\varphi$ satisfies condition (ii) because it satisfies condition (i) of Definition 2.16. We proceed with an induction on the number of quantifiers. Let $\mathcal{A}^{I}\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$ (which is possible if there are constants on each side of the positive equalities). That means for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{F}^{+}_{1}(\mathcal{A})\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$. If $\psi_{i^{\prime},l^{\prime}}$ is not a positive equality, then by Lemma 2.20, $\mathcal{A}^{I}\models\psi_{i^{\prime},l^{\prime}}[a_{0},...,a_{n}]$. Otherwise, by assumption it must be of the form $t(x_{0},...,x_{n})=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$ for some constants $\mathrm{c}_{0},...,\mathrm{c}_{v}\in\sigma$. Then, $\mathcal{A}^{I}\models t[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. One may easily check, by our definitions, that implies $\mathcal{F}_{1}^{+}(\mathcal{A})\models t[a_{0},...,a_{n}]=s(\mathrm{c}_{0},...,\mathrm{c}_{v})$. By the arbitrariness of $i^{\prime}$ and $l^{\prime}$, we therefore have $\mathcal{F}_{1}^{+}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$. The inductive step is straightforward by Lemma 2.18 (1) and (2). If $\varphi$ satisfies condition (ii) of Definition 2.16, as above, the result may be shown by an induction on the number of quantifiers. The base case may be tackled by the same reasoning presented above for when $\varphi$ satisfies condition (i) (that is, going down to the level of literals and applying Lemma 2.20), and the argument for the inductive steps are the same. If now $\varphi$ satisfies condition (iii) of Definition 2.16, we may also deal with the base case and inductive steps in the same way.
Suppose $\varphi$ satisfies condition (iii) of Definition 2.17. We proceed by induction on the number of quantifiers. If $\mathcal{A}^{I}\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then for each $i^{\prime}\leq k$ there is $l^{\prime}\leq p_{i^{\prime}}$ such that $\mathcal{A}^{I}\models\psi_{i^{\prime},l^{\prime}}[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$. As before, if $\psi_{i^{\prime},l^{\prime}}$ is not a positive equality, the conclusion is straightforward. Otherwise, by assumption there is $\gamma\in\bigvee\big(\{\psi_{i^{\prime},l}\}_{l\leq p_{i^{\prime}}}\setminus\mathrm{noncoll}(\bigvee_{l\leq p_{i^{\prime}}}\psi_{i^{\prime},l})\big)\setminus\{\top\}$ such that $\vdash\psi_{i^{\prime},l^{\prime}}\rightarrow\gamma$. That means $\mathcal{A}^{I}\models\gamma[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$. Since $\gamma$ is not a non-collapsible equality and $\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}$ is non-collapsible, $\gamma$ must not be a positive equality. Therefore, by Lemma 2.20, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\gamma[a_{0},...,a_{n}]$. Since, by construction, $\vdash\gamma\rightarrow\bigvee_{l\leq p_{i^{\prime}}}\psi_{i^{\prime},l}$, we get $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigvee_{l\leq p_{i^{\prime}}}\psi_{i^{\prime},l}[a_{0},...,a_{n}]$. By the arbitrariness of $i^{\prime}$ and $l^{\prime}$, we finally obtain $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}[a_{0},...,a_{n}]$. Once again, the inductive step is straightforward by Lemma 2.18 (1) and (2). ∎
Corollary 2.22.
For $\beta\in\mathbf{On}$, let $\mathcal{F}_{\beta}(\mathcal{A})$ be generated by the index set $I$, and $\varphi\in\mathcal{L}^{\sigma}$ be in prenex conjunctive normal form. If $\varphi$ is constant-free and non-collapsible, then for any $a_{0},...,a_{n}\in\mathcal{F}_{1}(A)$,
if $\mathcal{A}^{I}\models\varphi[\vartheta^{-1}(a_{0}),...,\vartheta^{-1}(a_{n})]$, then $\mathcal{F}_{\beta}(\mathcal{A})\models\varphi[a_{0},...,a_{n}]$.
Proof.
By induction on $\beta$, using Lemma 2.21. ∎
Theorem 2.23.
Let $\mathcal{F}^{+}_{1}(\mathcal{A})$ be generated by the index set $I$. For any non-collapsible sentence $\varphi\in\mathcal{L}^{\sigma}$,
if $\mathcal{A}^{I}\models\varphi$, then $\mathcal{F}^{+}_{1}(\mathcal{A})\models\varphi$.
Proof.
Let $\mathcal{A}^{I}\models\varphi$. If $\varphi$ is a $\Sigma_{n}$ sentence $\exists x\Delta\psi(x)$, then there is $a\in A^{I}$ such that $\mathcal{A}^{I}\models\Delta\psi[a]$. Since $\vartheta^{-1}$ is surjective, there is $a^{\prime}\in\mathcal{F}_{1}(A)$ such that $a=\vartheta^{-1}(a^{\prime})$. Thus, by Lemma 2.21, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a^{\prime}]$, and so $\mathcal{F}^{+}_{1}(\mathcal{A})\models\exists x\Delta\psi(x)$. If $\psi$ is a $\Pi_{n}$ sentence $\forall x\Delta\psi(x)$, let $a\in\mathcal{F}_{1}(A)$. If $a\in A^{I}$, then $\mathcal{A}^{I}\models\Delta\psi[a]$, so by the same reasoning as before, by Lemma 2.21 we get $\mathcal{F}^{+}_{1}(\mathcal{A})\models\Delta\psi[a]$, since, in that case, $a=\vartheta^{-1}(a)$. If $a\not\in A^{I}$, that is, if $a\in A$, notice, by Lemma 2.18 (2), $\Delta\psi(x)$ is also non-collapsible. Thus, by Lemma 2.21, $\mathcal{F}_{1}(\mathcal{A})\models\Delta\psi[a]$ iff $\mathcal{A}^{I}\models\Delta\psi[\vartheta^{-1}(a)]$, which we know is the case. Therefore, by the arbitrariness of $a$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\Delta\psi(x)$. ∎
Corollary 2.24.
Let $\beta\in\mathbf{On}$ and $\mathcal{F}_{\beta}(\mathcal{A})$ be generated by the index set $I$. For any constant-free non-collapsible sentence $\varphi\in\mathcal{L}^{\sigma}$,
if $\mathcal{A}^{I}\models\varphi$, then $\mathcal{F}_{\beta}(\mathcal{A})\models\varphi$.
Proof.
By induction on $\beta$, using Corollary 2.22. ∎
Lemma 2.25.
Let $\Delta\bigvee_{i\leq k}\psi_{i}\in\mathcal{L}^{\sigma}$ be a collapsible sentence, where each $\psi_{i}$ is a literal (that is, $\Delta\bigvee_{i\leq k}\psi_{i}$ is a formula in prenex normal form with a single conjunct). Then $\varphi$ is not preserved by cumulative powers.
Proof.
Since $\Delta\bigvee_{i\leq k}\psi_{i}$ is collapsible, the conditions of Definition 2.17 are falsified. As we have seen in the proof of Lemma 2.18, $\Delta\bigvee_{i\leq k}\psi_{i}$ contains a positive equality $t_{1}=t_{2}$ such that: for any $\gamma\in\bigvee\big(\{\psi_{i}\}_{i\leq k}\setminus\mathrm{noncoll}(\Delta\bigvee_{i\leq k}\psi_{i})\big)\setminus\{\top\}$, $\not\vdash t_{1}=t_{2}\rightarrow\gamma$; each side has occurring variables; each of its sides do not share at least one universally bound variables; there are existentially bound variables on at most one of its sides; the side which is not existentially bound also has no occurrence of constants. Since $\Delta\bigvee_{i\leq k}\psi_{i}$ is a sentence, we may calculate it must have the form
$\varphi=\Delta\big((\bigvee_{i<k}\psi_{i})\vee t_{1}(x^{\prime},x_{0},...,x_{n})=t_{2}(x_{0},...,x_{n},y_{0},...,y_{m},\mathrm{c}_{0},...,\mathrm{c}_{r})\big)$,
where $x_{0},...,x_{n},x^{\prime}$ are universally bound, $y_{0},...,y_{m}$ are existentially bound, $\mathrm{c}_{0},...,\mathrm{c}_{r}$ are constants, and no constants occur in $t_{1}$. Since $\big\{\neg\gamma\mid\gamma\in\bigvee\big(\{\psi_{i}\}_{i\leq k}\setminus\mathrm{noncoll}(\bigvee_{i\leq k}\psi_{i})\big)\setminus\{\top\}\big\}\cup\{t_{1}=t_{2}\}$ is consistent, by compactness there is a model $\mathcal{A}$ of $\varphi$ such that (a) $\mathcal{A}\models t_{1}=t_{2}$ and $\mathcal{A}\not\models\gamma$ for every $\gamma\in\bigvee\big(\{\psi_{i}\}_{i\leq k}\setminus\mathrm{noncoll}(\bigvee_{i\leq k}\psi_{i})\big)\setminus\{\top\}$. Suppose
$\mathcal{F}^{+}_{1}(\mathcal{A})\models\forall x\big((\bigvee_{i<k}\psi_{i})\vee t_{1}(x)=t_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})\big)\big)$,
Then for any $a\in\mathcal{F}_{1}(A)$, $\mathcal{F}^{+}_{1}(\mathcal{A})\models\bigvee_{i<k}\psi_{i}[a]\vee t_{1}[a]=t_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})$. Let $i^{\prime}<k$. If $\psi_{i^{\prime}}\in\big(\{\psi_{i}\}_{i\leq k}\setminus\mathrm{noncoll}(\Delta\bigvee_{i\leq k}\psi_{i})\big)\setminus\{\top\}$, then by (a) we have $\mathcal{A}\models\neg\psi_{i^{\prime}}$. By the arbitrariness of $a$, by Lemma 2.19, $\mathcal{F}_{1}^{+}(\mathcal{A})\models\neg\psi_{i^{\prime}}$. Otherwise, $\psi_{i^{\prime}}$ must be a collapsible positive equality. But then, since there is a single universal quantifier, by the same restrictions exposed at the beginning of our proof, $\psi_{i^{\prime}}$ must also be of the form $s_{1}(x)=s_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})$. Suppose $\mathcal{F}^{+}_{1}(\mathcal{A})\models s_{1}[a]=s_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})$. By the arbitrariness of $a$, we may let $a\in A$. But then $\rho\big(s_{1}^{\mathcal{F}^{+}_{1}(\mathcal{A})}[a]\big)=0$, and $\rho\big(s_{2}^{\mathcal{F}^{+}_{1}(\mathcal{A})}(\mathrm{c}_{0},...,\mathrm{c}_{r})\big)=\rho(\overline{\mathrm{c}_{0}})=1$, which means $\mathcal{F}^{+}_{1}(\mathcal{A})\not\models s_{1}[a]=s_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})$. Therefore, by the arbitrariness of $i^{\prime}$, we must have $\mathcal{F}^{+}_{1}(\mathcal{A})\models s_{1}[a]=s_{2}(\mathrm{c}_{0},...,\mathrm{c}_{r})$. But then the same reasoning applies, and we have a contradiction. The inductive steps for the quantifiers are straightforward by Lemma 2.18 (3) and (4). ∎
Theorem 2.26.
Let $\varphi\in\mathcal{L}^{\sigma}$ be a collapsible sentence. Then $\varphi$ is not preserved by cumulative powers.
Proof.
Let $\Delta\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}$ be the prenex conjunctive normal form of $\varphi$. If $\varphi$ is preserved, then for each and every $i^{\prime}\leq k$, $\Delta\bigvee_{l\leq p_{i^{\prime}}}\psi_{i^{\prime},l}$ is preserved. But since $\Delta\bigwedge_{i\leq k}\bigvee_{l\leq p_{i}}\psi_{i,l}$ is collapsible, by Definition 2.17 we may see, for some $i^{\prime\prime}$, $\Delta\bigvee_{l\leq p_{i^{\prime\prime}}}\psi_{i^{\prime\prime},l}$ must be collapsible, and therefore, by Lemma 2.25, not preserved by cumulative powers. ∎
Corollary 2.27.
Let $\varphi\in\mathcal{L}^{\sigma}$ be a sentence. Then:
  • (1)
    $\varphi$ is preserved by finite cumulative powers iff it is a non-collapsible direct power sentence;
  • (2)
    $\varphi$ is preserved by arbitrary cumulative powers iff it is a constant-free non-collapsible direct power sentence.
Proof.
(1) The left to right direction is given by Theorem 2.26. For the other, suppose $\varphi$ is a non-collapsible direct power sentence, then for any $\mathcal{A}$ and index set $I$, if $\mathcal{A}\models\varphi$, then $\mathcal{A}^{I}\models\varphi$. By Theorem 2.23, that means $\mathcal{F}_{1}^{+}(\mathcal{A})\models\varphi$. The preservation for further levels of the cumulative power hierarchy may be easily given by induction on the levels of the construction.
(2) Similarly, by using Corollary 2.22
Therefore, for $n<\omega$, $\mathcal{F}^{+}_{n}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by non-collapsible direct power sentences; and for $\beta\in\mathbf{On}$, $\mathcal{F}_{\beta}(\mathcal{A})$ inherits any first-order property from $\mathcal{A}$ characterized by constant-free non-collapsible direct power sentences – since, for $\beta>0$, $\mathcal{F}_{\beta}(\mathcal{A})$ does not have an interpretation of the constants of $\sigma$. Not only that, those are the precisely the classes of first-order sentences inherited by cumulative powers. In other words, just as the fragment preserved by reduced products is precisely the one given by Horn sentences, and the fragment preserved by direct products is precisely the one characterized by Weinstein, the fragment preserved by (finite) cumulative powers is precisely the one given by non-collapsible direct power sentences.
Notice the restrictions of the results on non-collapsible formulas concern precisely equality. That is related to the definition of the extended functions. In our former example, for $j\in\mathbb{N}$, $\vartheta^{1}_{j}(\overline{3})=\overline{3}(j)=3$, $\vartheta^{1}_{j}(\overline{-1})=-1$ and $\vartheta^{1}_{j}(2)=2$, so $\forall j\in\mathbb{N}\big(\vartheta^{1}_{j}(\overline{3})+\vartheta^{1}_{j}(\overline{-1})=\vartheta^{1}_{j}(2)\big)$. However, $\mathcal{F}_{1}(\mathbb{N})\not\models\overline{3}+\overline{-1}=2$, as $\overline{3}+\overline{-1}=\overline{2}\neq 2$. The restriction has also a deeper meaning. It concerns the difference between the definition of other extended relations and the equality relation. While at each level $\beta$ every other relation is defined by using the $\vartheta$ function, and therefore take into account all of the levels lesser than $\beta$ – so that, for example, if $\rho(a)<\rho(b)=\beta$ for a successor $\beta$, then $\mathrm{R}_{\beta}ab$ iff $\mathrm{R}_{\beta-1}ab(j)$ –, the equality relation is interpreted as real identity, and thus holds between two objects iff they are the same object, regardless of how they are hereditarily related. In a few words, whereas every other relation is hereditarily defined, equality is not.
In Section 3, we shall see that once we also define equality hereditarily, the resulting structure is isomorphic to a direct power.
At last, we may see embeddability can be lifted from the generating structures to their respective $\mathcal{F}$-hierarchies.
Theorem 2.28.
Let $\mathcal{A}$ and $\mathcal{B}$ be $\sigma$-structures, and their cumulative power hierarchies be respectively generated by the families of index sets $\{I_{\kappa}\}_{\kappa\in\mathbf{On}}$ and $\{J_{\kappa}\}_{\kappa\in\mathbf{On}}$. If $\mathcal{A}\hookrightarrow\mathcal{B}$ and for each $\beta\in\mathbf{On}$ $|I_{\beta}|\leq|J_{\beta}|$, then for any $\beta,\gamma\in\mathbf{On}$ such that $\beta\leq\gamma$, $\mathcal{F}_{\beta}(\mathcal{A})\hookrightarrow\mathcal{F}_{\gamma}(\mathcal{B})$.
Proof.
Let $e:A\to B$ be the relevant embedding, and for each $\alpha\in\mathbf{On}$ let $u_{\alpha}:I_{\alpha}\to J_{\alpha}$ an injection. We inductively define the mappings $e_{\alpha}:\mathcal{F}_{\alpha}(\mathcal{A})\to\mathcal{F}_{\alpha}(\mathcal{B})$ in the following way:
  • $e_{0}=e$;
  • for the inductive step, let $k\in u_{\alpha}[I_{\alpha}]$. Then, let $e_{\alpha^{+}}(a)=b$ iff
    $\forall j\in u_{\alpha}[I_{\alpha}]\Big(b(j)=e_{\alpha}\big(a(u^{-1}_{\alpha}(j))\big)\Big)$ and $\forall j\in J_{\alpha}\setminus u_{\alpha}[I_{\alpha}]\Big(b(j)=e_{\alpha}\big(a(u^{-1}_{\alpha}(k))\big)\Big)$.
For a limit $\lambda$, we naturally let $e_{\lambda}(a)=b$ iff there is $\delta<\lambda$ such that $e_{\delta}(a)=b$. One might easily check $e_{\alpha^{+}}$ is injective, and preserves all the functions, relations and predicates $\mathbf{c}_{\alpha^{+}}$ of $\sigma$, so that for each $\alpha\in\mathbf{On}$, $e_{\alpha}$ is an embedding. From there, the conclusion is straightforward. ∎
Notice the embeddings are not unique, as we may see there are at least $|J_{\beta}\setminus u_{\beta}[I_{\beta}]|$ different embeddings from $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ into $\mathcal{F}_{\beta^{+}}(\mathcal{B})$.
3. Direct powers and the hierarchy
As the past results show, the $\mathcal{F}$-hierarchies generated, as a whole, present a structure that resembles those of their generating structure, but nevertheless fail to instantiate the generating structure itself. How can we recover more of the first-order properties of the generating structures? A way of doing that is by restricting the elements of the hierarchy to those of the parallel hierarchy which is not cumulative – in other words, a hierarchy if direct powers. Consider the following definition.
Definition 3.1 (Direct power hierarchy).
Let $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ be a family of index sets. The direct power hierarchy generated by $\mathcal{A}$ and $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ is composed by the structures $\Pi_{\beta}(\mathcal{A})$, where each $\Pi_{\beta^{+}}(\mathcal{A})$ is the direct power of $\Pi_{\beta}(\mathcal{A})$ with index set $I_{\beta}$, and for a limit $\lambda$, $\Pi_{\lambda}(\mathcal{A})$ is the limit of embeddings of the preceding structures. More precisely:
  • $\Pi_{0}(\mathcal{A})=\mathcal{A}$;
  • $\Pi_{\beta^{+}}(\mathcal{A})=\langle\Pi_{\beta}(\mathcal{A})^{I_{\beta}},\langle\mathrm{F}_{i,\beta^{+}}\rangle_{i\in J},\langle\mathrm{R}_{i,\beta^{+}}\rangle_{i\in J},\langle\mathrm{c}_{i,\beta^{+}}\rangle_{i\in K}\rangle$, [^10] For simplicity, we use the same symbol to represent the functions, relations and constants of $\mathcal{F}^{-}_{\beta}(\mathcal{A})$ and $\mathcal{F}_{\beta}(\mathcal{A})$. However, as we shall see, there is no problem in assuming they are the same, as those of $\mathcal{F}^{-}_{\beta}(\mathcal{A})$ coincide with those of $\mathcal{F}_{\beta}(\mathcal{A})$ when restricted to the domain of the former, when for each $\alpha<\beta$ the index set is $\mathcal{F}_{\alpha}(\mathcal{A})$ itself.
    where $\mathrm{F}_{i,\beta^{+}}$, $\mathrm{R}_{i,\beta^{+}}$ and $\mathrm{c}_{i,\beta^{+}}$ are inductively defined in the following way:
    • (a)
      The operation $\mathrm{F}_{i,\beta^{+}}$:
      $\mathrm{F}_{i,\beta^{+}}(a_{0},...,a_{n})=b$ iff $\forall j\in I_{\beta}\Big(\mathrm{F}_{i,\beta}\big(a_{0}(j),...,a_{n}(j)\big)=b(j)\Big)$;
    • (b)
      The relation $\mathrm{R}_{i,\beta^{+}}$:
      $\mathrm{R}_{i,\beta^{+}}a_{0}...a_{n}$ iff $\forall j\in I_{\beta}\big(\mathrm{R}_{i,\beta}a_{0}(j)...a_{n}(j)\big)$.
    • (c)
      The constant $\mathrm{c}_{i,\beta^{+}}$:
      $\mathrm{c}_{i,\beta^{+}}=\overline{\mathrm{c}_{i,\beta}}$;
  • $\Pi_{\lambda}(\mathcal{A})=\bigcup_{\alpha<\lambda}\Pi_{\alpha}(\mathcal{A})/\!\equiv_{\lambda}$, for a limit $\lambda$, is the limit of the elementary chain, with the embeddings inductively defined by
    • $e_{\alpha}^{\alpha}$ is the identity function,
    • $e_{\alpha}^{\beta^{+}}:\Pi_{\alpha}(\mathcal{A})\to\Pi_{\beta^{+}}(\mathcal{A});x\mapsto\overline{e_{\alpha}^{\beta^{+}}(x)}_{\beta}$,
    • $e_{\alpha}^{\lambda}:\Pi_{\alpha}(\mathcal{A})\to\Pi_{\lambda}(\mathcal{A});x\mapsto\llbracket x\rrbracket_{\lambda}$, for a limit $\lambda$,
where $\llbracket\cdot\rrbracket_{\lambda}$ denotes the equivalence class defined by the relation $x\equiv_{\lambda}y$ iff there are $\alpha\leq\beta<\lambda$ such that either $e_{\alpha}^{\beta}(x)=y$ or $e_{\alpha}^{\beta}(y)=x$.
By Corollary 2.15, we know what sentences are preserved by $\Pi_{\beta}(\mathcal{A})$. Suppose we let the same family $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ define both a cumulative power and a direct power hierarchy. Then, clearly $\Pi_{\beta}(\mathcal{A})\subset\mathcal{F}_{\beta}(\mathcal{A})$, which means there is a fragment of $\mathcal{F}_{\beta}(\mathcal{A})$ which is a proper extension of $\mathcal{A}$ that preserves any direct power sentence. Suppose, however, we would like to shape the cumulative power hierarchy itself, with all of its cross-level element functions, rather than a fragment of it, into a structure which preserves more first-order properties from its generating structure. That may be done in varying levels of faithfulness, in terms of satisfaction of $\mathcal{L}^{\sigma}$-sentences. The most faithful option, for successor levels, is to quotient each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an adequate ultrafilter over $I_{\beta}$. But there is an intermediate level by quotienting $\mathcal{F}_{\beta}(\mathcal{A})$ with an appropriate equivalence relation, the product of which is isomorphic to $\Pi_{\beta}(\mathcal{A})$.
Definition 3.2.
Let $\beta\in\mathbf{On}$. For $a,b\in\mathcal{F}_{\beta}(A)$, let $\delta=max\{\rho(a),\rho(b)\}$. We inductively define the equivalence relation $\equiv_{\beta}$ on $\mathcal{F}_{\beta}(\mathcal{A})$ in the following way:
  • (a)
    $a\equiv_{0}b$ iff $a=b$;
  • (b)
    $a\equiv_{\beta^{+}}b$ iff $\forall j\in I_{\delta-1}\big(\vartheta^{\delta}_{j}(a)\equiv_{\delta-1}\vartheta_{j}^{\delta}(b)\big)$;
  • (c)
    $a\equiv_{\lambda}b$ iff for some $\alpha<\lambda$, $a\equiv_{\alpha}b$, for a limit $\lambda$.
Notice the definition makes sense: $\delta\leq\beta$, which means $\delta-1<\beta$, and for any $\beta\in\mathbf{On}$, if $\rho(a)=\beta$, then by our definition, $\beta$ is not a limit ordinal – that is, at limit steps of the construction, no new element is introduced –, so that $\delta-1$ indeed exists.
Notice, furthermore, that following our considerations in Section 2, the above definition amounts to a hereditary definition of equality, that is, equality is defined modulo the $\vartheta$ function. It should therefore be no surprise the resulting structures are isomorphic to direct powers.
Proposition 3.3.
$\equiv_{\beta}$ is an equivalence relation.
Proof.
By induction on $\beta$. The base case is trivial, and the limit case is not hard to check, so we show the inductive step. Reflexivity and symmetry are trivial. For transitivity, let $a\equiv_{\alpha^{+}}b\equiv_{\alpha^{+}}c$. The non-trivial case is the one in which
(i) $\forall j\in I_{\rho(a)-1}\big(a(j)\equiv_{\rho(a)-1}b\big)$ and (ii) $\forall j\in I_{\rho(b)-1}\big(b(j)\equiv_{\rho(a)-1}c\big)$.
Notice that means (iii) $\rho(a)>\rho(b)>\rho(c)$. We therefore proceed by induction on $\rho(a)$. For the base case, we then have to let $\rho(a)=2$. Now, let $k\in I_{1}$, which by (i) means (for $\delta=max\{\rho\big(a(k)\big),\rho(b)\}$)
either (iv) $a(k)=b$, or (v) $\forall j\in I_{0}\Big(\vartheta_{j}^{\delta}\big(a(k)\big)\equiv_{1}\vartheta^{\delta}_{t}(b)\Big)$.
If (iv), then by assumption $a(k)\equiv_{1}c$. If (v), suppose $\rho\big(a(k)\big)>\rho(b)$. By (iii), that means (vi) $\rho\big(a(k)\big)=1$ and $\rho(b)=0$. Thus, $\delta=\rho\big(a(k)\big)=1$, and so (vii) $\vartheta^{1}_{j}\big(a(k)\big)=\big(a(k)\big)(j)$ and $\vartheta^{1}_{j}(b)=b$, so that
$\forall j\in I_{0}\Big(\big(a(k)\big)(j)\equiv_{1}b\Big)$.
But by (vi), we have $\rho\Big(\big(a(k)\big)(j)\Big)=0$, so by Definition 3.2, (vii) and (v) mean
$\forall j\in I_{0}\Big(\big(a(k)\big)(j)=b\Big)$, so that $\forall j\in I_{0}\Big(\big(a(k)\big)(j)\equiv_{1}c\Big)$.
Otherwise, that is, if $\rho(b)>\rho\big(a(k)\big)$, then $\rho\big(a(k)\big)=0$ and $\rho(b)=1$. Thus, $\vartheta^{1}_{j}\big(a(k)\big)=a(k)$ and $\vartheta^{1}_{j}(b)=b(j)$, and so from (v) we have
$\forall j\in I_{0}\big(b(j)\equiv_{1}a(k)\big)$,
but since $\rho(b)=1$, we have $\rho\big(b(j)\big)=0$, so by Definition 3.2,
$\forall j\in I_{0}\big(b(j)=a(k)\big)$.
By (ii), we then get $a(k)\equiv_{1}c$. We conclude $a\equiv_{1}c$. The induction step is straightforward, as the analysis of each case goes to the previous level, where the induction hypothesis may be easily applied. ∎
Proposition 3.4.
For any $\beta^{+}$ and $a\in\mathcal{F}_{\beta^{+}}(A)$ there is $b\in\mathcal{F}_{\beta^{+}}^{*}(A)$ such that $b\equiv_{\beta^{+}}a$.
Proof.
Just consider $\overline{a}\in\mathcal{F}_{\beta^{+}}^{*}(A)$. Clearly, $\overline{a}\equiv_{\beta^{+}}a$. ∎
Thus, when comparing two arbitrary elements with respect to $\equiv_{\beta}$, we may consider them, without loss of generality, to be of the same level, as, by the above theorem, for any two elements of different levels there is always an element of the highest level between the two to which it is equivalent under $\equiv_{\beta}$.
Let $a\in\mathcal{F}_{\beta}(A)$. We define the equivalence classes under $\equiv_{\beta}$ as usual, and denote that of $a$ by $a]_{\beta}$. In the absence of ambiguity, we write simply $[a]$.
Definition 3.5.
Define $\mathfrak{F}_{\beta}(\mathcal{A})=\langle\mathcal{F}^{\beta}(A)/\!\equiv_{\beta},\mathrm{F}^{\equiv}_{\beta},\mathrm{R}^{\equiv}_{\beta},\mathrm{c}^{\equiv}_{\beta}\rangle$ as usual – that is:
  • (a)
    $\mathrm{F}^{\equiv}_{\beta}:(\mathcal{F}^{\beta}(A)/\!\equiv_{\beta})^{n}\ \to(\mathcal{F}^{\beta}(A)/\!\equiv_{\beta});\langle[x_{0}],...,[x_{n}]\rangle\mapsto[\mathrm{F}_{\beta}(x_{0},...,x_{n})]$;
  • (b)
    $\mathrm{c}^{\equiv}_{\beta}=[\mathrm{c}]$;
  • (c)
    $\mathrm{R}^{\equiv}_{\beta}[a_{0}]...[a_{n}]$ iff $\exists b_{0},...,b_{n}\in\mathcal{F}^{\beta}(A)(a_{0}\equiv_{\beta}b_{0}\ \&\ ...\ \&\ a_{n}\equiv_{\beta}b_{n}\ \&\ \mathrm{R}_{\beta}b_{0}...b_{n})$.
We may now easily check that:
Theorem 3.6.
Let a direct power hierarchy and a cumulative power hierarchy of $\mathcal{A}$ both be generated and the family of index sets $\{I_{\alpha}\}_{\mathbf{On}}$. Then, for any $\beta\in\mathbf{On}$, $\Pi_{\beta}(\mathcal{A})\cong\mathfrak{F}_{\beta}(\mathcal{A})$.
Proof.
Just take the isomorphism $e:\Pi_{\beta}(A)\to\mathcal{F}_{\beta}(A)/\!\equiv_{\beta};x\mapsto[x]$. ∎
Therefore, cumulative powers may be seen as generalisations of direct powers, as the latter is obtained from the former by simply quotienting it by an equivalence relation defined by hereditary identity.
Corollary 3.7.
A sentence is preserved by $\mathfrak{F}_{\beta}(\mathcal{A})$ iff it is logically equivalent to a sentence $\psi\in\mathcal{L}^{\sigma}$ such that, for some $\Phi\subseteq\mathcal{L}^{\sigma}$ and $n<\omega$, $\psi\in\Phi_{n}$ and $R(\psi,\psi,\psi)$ (that is, iff it is a direct power sentence).
Proof.
By Corollary 2.15 and Theorem 3.6. ∎
Theorem 3.8.
Let $\beta\in\mathbf{On}$ and $\mathcal{A}$ and $\mathcal{B}$ be $\sigma$-structures. If $\mathcal{A}\hookrightarrow\mathcal{B}$, then $\mathfrak{F}_{\beta}(\mathcal{A})\hookrightarrow\mathfrak{F}_{\beta}(\mathcal{B})$.
Proof.
By Theorem 2.28, there is an embedding $e:\mathcal{F}_{\beta}(A)\to\mathcal{F}_{\beta}(B)$. Thus, we need only consider the embedding $e^{\prime}:\mathfrak{F}_{\beta}(\mathcal{A})\to\mathfrak{F}_{\beta}(\mathcal{B});[x]\mapsto[e(x)]$. As before, the embedding is not unique. ∎
4. Ultrapowers and the hierarchy
As we have previously mentioned, to shape the hierarchy into the same first-order structure of its generating structure, it suffices to adapt the ultrapower construction to the present case. That would entail quotienting each level $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$. The problem, of course, is that there are many elements in $\mathcal{F}_{\beta^{+}}(A)$, for $\beta>0$, which are not functions. More so, for $\beta>1$, there are pairs of elements who are functions, but whose domains differ. So simply quotienting $\mathcal{F}_{\beta^{+}}(\mathcal{A})$ by an ultrafilter over $I_{\beta}$ is not a valid construction. That could be worked around by quotienting $\mathcal{F}^{*}_{\beta^{+}}(A)$ by an ultrafilter over $I_{\beta}$. We follow a different route by defining the equivalence classes modulo an ultrafilter by using equality modulo the $\vartheta$ function, as in the construction of $\mathfrak{F}_{\beta}(\mathcal{A})$.
Definition 4.1.
Let $a,b\in\mathcal{F}_{\beta}(A)$. Consider a family $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ of ultrafilters, where each $\mathcal{U}_{\beta}$ is an ultrafilter over $I_{\beta}$. We inductively define a hierarchy of relations such that:
  • (a)
    $a\sim_{0}b$ iff $a=b$,
For $\beta\geq 0$ and $\delta=max\{\rho(a),\rho(b)\}$,
  • (b)
    $a\sim_{\beta^{+}}b$ iff $\{j\in I_{\beta}\mid\vartheta_{j}^{\delta}(a)\sim_{\beta}\vartheta^{\delta}_{j}(b)\}\in\mathcal{U}_{\beta}$
For a limit $\lambda$, just let $a\sim_{\lambda}b$ iff $a\sim_{\alpha}b$ for some $\alpha<\lambda$, that is, $\sim_{\lambda}\ =\bigcup_{\alpha<\lambda}\sim_{\alpha}$.
Thus, intuitively, $a\sim_{\beta^{+}}b$ when either $a\sim_{\alpha}b$ for $\alpha<\beta^{+}$, or when $a$ and $b$ differ modulo $\sim_{\beta}$ in an amount of points that are ultrafilter-small. Now, we may see that:
Proposition 4.2.
For any cumulative power hierarchy and its associated relations $\equiv_{\beta}$ and $\sim_{\beta}$, for any $\beta\in\mathbf{On}$, $\equiv_{\beta}\ \subseteq\ \sim_{\beta}$.
Proof.
To see so, just notice the resemblance between the inductive definition of both relations. ∎
Therefore, similarly to our previous consideration of Proposition 3.4, when proving results about $\sim_{\beta}$ we may consider different elements of the cumulative power hierarchy to be of the same level without loss of generality. More precisely, whenever we have $a\sim_{\beta^{+}}b$ for some arbitrary $a$ and $b$, we may let $\rho(a)=\rho(b)=\beta^{+}$, which then means (for $\delta=max\{\rho(a),\rho(b)\}$ and $j\in I_{\beta}$) $\vartheta^{\delta}_{j}(a)=a(j)$ and $\vartheta^{\delta}_{j}(b)=b(j)$, so that
$a\sim_{\beta^{+}}b$ iff $\{j\in I_{\beta}\mid a(j)\sim_{\beta}b(j)\}\in\mathcal{U}_{\beta}$.
Proposition 4.3.
$\sim$ is an equivalence relation on $\mathcal{F}_{\beta}(A)$.
Proof.
By induction on $\beta$. The base and limit cases are straightforward, as is the inductive step with respect to reflexivity and symmetry. For transitivity, let $a\sim_{\alpha^{+}}b\sim_{\alpha^{+}}c$. As observed above, we may let $\rho(a)=\rho(b)=\rho(c)=max\{\rho(a),\rho(b),\rho(c)\}=\alpha^{+}$. By Definition 4.1, that means
$\{j\in I_{\alpha}\mid a(j)\sim_{\alpha}b(j)\},\{j\in I_{\alpha}\mid b(j)\sim_{\alpha}c(j)\}\in\mathcal{U}_{\alpha}$,
and therefore
$\{j\in I_{\alpha}\mid a(j)\sim_{\alpha}b(j)\}\cap\{j\in I_{\alpha}\mid b(j)\sim_{\alpha}c(j)\}\in\mathcal{U}_{\alpha}$.
But by induction hypothesis,
$\{j\in I_{\alpha}\mid a(j)\sim_{\alpha}b(j)\}\cap\{j\in I_{\alpha}\mid b(j)\sim_{\alpha}c(j)\}\subseteq\{j\in I_{\alpha}\mid a(j)\sim_{\alpha}c(j)\}$,
and therefore $a\sim_{\alpha^{+}}c$. ∎
Let $a\in\mathcal{F}_{\beta}(A)$. We define its equivalence class with respect to $\sim_{\beta}$ as usual, and denote it by $a]_{\beta}$. Similarly, for $j\in\mathrm{dom}(a)$, we denote by $a(j)]_{\beta}$ the equivalence class of $a(j)$. We may omit the subscripts, writing $[a]$ and $\sim$, when the level of the construction is clear by context. Notice we are using the same notation we have for the equivalence classes under $\equiv_{\beta}$. We hope the context makes it unambiguous to what equivalence class we are referring, especially as we shall, from now on, only consider $\sim_{\beta}$.
Proposition 4.4.
For any $\alpha\leq\beta$, $\sim_{\alpha}\ \subseteq\ \sim_{\beta}$.
Proof.
Directly by Definitions 2.3 and 4.1. ∎
Definition 4.5.
We define the structures $\mathbb{F}_{\beta}(\mathcal{A})=\langle\mathcal{F}_{\beta}(A)/\!\sim_{\beta},\mathrm{F}_{\beta}^{\sim},\mathrm{R}_{\beta}^{\sim},\mathrm{c}^{\sim}_{\beta}\rangle$ in the usual manner:
  • (a)
    $\mathrm{F}^{\sim}_{\beta}:(\mathcal{F}_{\beta}(A)/\!\sim_{\beta})^{n}\to(\mathcal{F}_{\beta}(A)/\!\sim_{\beta});\langle[x_{0}],...,[x_{n}]\rangle\mapsto[\mathrm{F}_{\beta}(x_{0},...,x_{n})]$;
  • (b)
    $\mathrm{c}^{\sim}_{\beta}=[\mathrm{c}]$;
  • (c)
    $\mathrm{R}^{\sim}_{\beta}[a_{0}]...[a_{n}]$ iff $\exists b_{0},...,b_{n}\in\mathcal{F}_{\beta}^{*}(A)(a_{0}\sim_{\beta}b_{0}\ \&\ ...\ \&\ a_{n}\sim_{\beta}b_{n}\ \&\ \mathrm{R}_{\beta}b_{0}...b_{n})$.
Lemma 4.6.
Let $a_{0},...,a_{n},b\in\mathcal{F}_{\beta^{+}}^{*}(A)$, and $\mathcal{U}_{\beta}$ be the ultrafilter used for the definition of $\sim_{\beta^{+}}$. Then:
  • (a)
    $\mathrm{F}^{\sim}_{\beta^{+}}([a_{0}],...,[a_{n}])=[b]$ iff $\{j\in I_{\beta}\mid\mathrm{F}^{\sim}_{\beta}\big([a_{0}(j)],...,[a_{n}(j)]\big)=[b(j)]\}\in\mathcal{U}_{\beta}$;
  • (b)
    $\mathrm{R}^{\sim}_{\beta^{+}}[a_{0}]...[a_{n}]$ iff $\{j\in I_{\beta}\mid\mathrm{R}^{\sim}_{\beta}[a_{0}(j)]...[a_{n}(j)]\}\in\mathcal{U}_{\beta}$.
Proof.
(a): By definition, $\mathrm{F}^{\sim}_{\beta^{+}}([a_{0}],...,[a_{n}])=[b]$ iff $b\sim_{\beta^{+}}\mathrm{F}^{\sim}_{\beta^{+}}(a_{0},...,a_{n})$, iff
$\{j\in I_{\beta}\mid\mathrm{F}_{\beta}\big(a_{0}(j),...,a_{n}(j)\big)\sim_{\beta}b(j)\}\in\mathcal{U}_{\beta}$, [^11] Notice, in the context of our previous observations, we are under the assumption $\rho(a_{i})=\rho(b_{i})=\beta^{+}$.
but
$\{j\in I_{\beta}\mid\mathrm{F}_{\beta}\big(a_{0}(j),...,a_{n}(j)\big)\sim_{\beta}b(j)\}=$
$\{j\in I_{\beta}\mid\big[\mathrm{F}_{\beta}\big(a_{0}(j),...,a_{n}(j)\big)\big]=[b(j)]\}=$
$\{j\in I_{\beta}\mid\mathrm{F}^{\sim}_{\beta}\big([a_{0}(j)],...,[a_{n}(j)]\big)=[b(j)]\}$.
(b): By definition, $\mathrm{R}^{\sim}_{\beta^{+}}[a_{0}]...[a_{n}]$ iff there are $b_{0},...,b_{n}\in\mathcal{F}_{\beta^{+}}^{*}(A)$ such that $a_{i}\sim_{\beta^{+}}b_{i}$ and $\mathrm{R}_{\beta^{+}}b_{0}...b_{n}$. By Definition 2.3, that is the case iff
$\{j\in I_{\beta}\mid\mathrm{R}_{\beta}b_{0}(j)...b_{n}(j)\}=\mathcal{F}_{\beta}(A)\in\mathcal{U}_{\beta}$,
and
$\{j\in I_{\beta}\mid a_{i}(j)\sim_{\beta}b_{i}(j)\}\in\mathcal{U}_{\beta}$.
Since $n<\omega$, that means the intersection of all of the above sets is in $\mathcal{U}_{\beta}$, so by definition, that is the case iff (since an ultrafilter is closed by containment)
$\{j\in I_{\beta}\mid\mathrm{R}^{\sim}_{\beta}[a_{0}(j)]...[a_{n}(j)]\}\in\mathcal{U}_{\beta}$.
Then, it is straightforward to adapt Łos’s theorem to the present case, so that we know:
Theorem 4.7.
For any $\varphi\in\mathcal{L}^{\sigma}$ and $a_{1},...,a_{n}\in\mathcal{F}_{\beta^{+}}^{*}(A)$,
$\mathbb{F}_{\beta^{+}}(\mathcal{A})\models\varphi\big[[a_{1}],...,[a_{n}]\big]$ iff $\{j\in I_{\beta}\mid\mathbb{F}_{\beta}(\mathcal{A})\models\varphi\big[[a_{1}(j)],...,[a_{n}(j)]\big]\}\in\mathcal{U}_{\beta}$.
Proof.
Straightforward by Definition 4.5 and Lemma 4.6. ∎
We thus obtain:
Corollary 4.8.
For any $\alpha\leq\beta$, $\mathbb{F}_{\alpha}(\mathcal{A})\preceq\mathbb{F}_{\beta}(\mathcal{A})$, and therefore $\mathcal{A}\preceq\mathbb{F}_{\beta}(\mathcal{A})$.
Therefore, quotienting any stage of the hierarchy by the $\sim$ relation produces a structure elementary equivalent to the one generating the hierarchy.
Suppose now we have $\mathcal{A}\hookrightarrow\mathcal{B}$. We have shown embeddability to lift to the structures presented in Sections 2 and 3, so that $\mathcal{F}_{\beta}(\mathcal{A})\hookrightarrow\mathcal{F}_{\beta}(\mathcal{B})$ and $\mathfrak{F}_{\beta}(\mathcal{A})\hookrightarrow\mathfrak{F}_{\beta}(\mathcal{B})$ for $\beta\in\mathbf{On}$. Clearly, it is not as straightforward a fact that embeddability is lifted to the present structures. When is it the case that $\mathbb{F}_{\beta}(\mathcal{A})\hookrightarrow\mathbb{F}_{\beta}(\mathcal{B})$? It is not hard to see the question is related to the issue of the criteria of embeddability between ultrapowers.
It is well known that $\mathcal{A}^{I}/\mathcal{U}$ is isomorphic to $\mathcal{A}$ iff either $\mathcal{A}$ is finite or $\mathcal{U}$ is principal. Therefore, it is trivial that if $\mathcal{A}\hookrightarrow\mathcal{B}$ and either $\mathcal{A}$ and $\mathcal{B}$ are finite or $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ are principal, then $\mathcal{A}^{I}/\mathcal{U}_{I}\hookrightarrow\mathcal{B}^{J}/\mathcal{U}_{J}$. It is less clear, however, under what conditions embeddability may be lifted to the ultrapowers given both structures are infinite and the ultrapower of the embedding structure is non-isomorphic to its generating structure.
Definition 4.9 (Rudin-Keisler ordering [17]).
Let $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. The Rudin-Keisler ordering of ultrafilters $\leq_{\mathrm{RK}}$ is defined by $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ when there is $h:J\to I$ such that $X\in\mathcal{U}_{I}$ iff $h^{-1}[X]\in\mathcal{U}_{J}$. Two ultrafilters are Rudin-Keisler equivalent, written $\mathcal{U}_{I}\approx_{\mathrm{RK}}\mathcal{U}_{J}$, when $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ and $\mathcal{U}_{J}\leq_{\mathrm{RK}}\mathcal{U}_{I}$. [^12] In fact, that is only a preorder. The Rudin-Keisler (partial) ordering is defined on types of ultrafilters, which are defined as the equivalence classes of ultrafilters under the defined equivalence. For briefness, we use the concept as defined.
In the present context, the Rudin-Keisler ordering of ultrafilters is particularly relevant given the following result.
Proposition 4.10 ([2], p. 87).
Let $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. Then $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ iff for any structure $\mathcal{A}$, $\mathcal{A}^{I}/\mathcal{U}_{I}\preceq\mathcal{A}^{J}/\mathcal{U}_{J}$.
Since it is trivial to check that: [^13] To see so, just notice the identity map on $\mathcal{A}^{I}/\mathcal{U}$ is an embedding into $\mathcal{B}^{I}/\mathcal{U}$.
Lemma 4.11.
Let $\mathcal{A}\hookrightarrow\mathcal{B}$ and $\mathcal{U}$ be an ultrafilter over $I$. Then, $\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{B}^{I}/\mathcal{U}$.
We therefore obtain:
Theorem 4.12.
Let $\mathcal{A}\hookrightarrow\mathcal{B}$ and $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. If $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$, then $\mathcal{A}^{I}/\mathcal{U}_{I}\hookrightarrow\mathcal{B}^{J}/\mathcal{U}_{J}$.
Proof.
By Proposition 4.10, $\mathcal{A}^{I}/\mathcal{U}_{I}\preceq\mathcal{A}^{J}/\mathcal{U}_{J}$, and by Lemma 4.11, $\mathcal{A}^{J}/\mathcal{U}_{J}\hookrightarrow\mathcal{B}^{J}/\mathcal{U}_{J}$, so that $\mathcal{A}^{I}/\mathcal{U}_{I}\hookrightarrow\mathcal{B}^{J}/\mathcal{U}_{J}$. ∎
Therefore, we have a sufficient criterion for the embeddability of $\mathcal{A}$ into $\mathcal{B}$ to be lifted to an embeddability of $\mathcal{A}^{I}/\mathcal{U}_{I}$ into $\mathcal{B}^{J}/\mathcal{U}_{J}$.
Lemma 4.13.
Let $\lambda$ be a limit and $\mathbb{F}_{\beta}(\mathcal{A})\hookrightarrow\mathbb{F}_{\beta}(\mathcal{B})$ for every $\beta<\lambda$. Then, $\mathbb{F}_{\lambda}(\mathcal{A})\hookrightarrow\mathbb{F}_{\lambda}(\mathcal{B})$.
Proof.
[…]
Corollary 4.14.
Let $\mathcal{A}\hookrightarrow\mathcal{B}$ (respectively, $\mathcal{A}\preceq\mathcal{B}$), $\beta\in\mathbf{On}$, and their cumulative power hierarchies be respectively generated by $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ and $\{J_{\alpha}\}_{\alpha\in\mathbf{On}}$. For each $\alpha\in\mathbf{On}$, let $|I_{\alpha}|\leq|J_{\alpha}|$, and for each ultrafilter $\mathcal{U}_{I_{\beta}}$ and $\mathcal{U}_{J_{\beta}}$ respectively defining $\sim^{\mathcal{A}}_{\beta+1}$ and $\sim^{\mathcal{B}}_{\beta+1}$, let $\mathcal{U}_{I_{\beta}}\leq_{\mathrm{RK}}\mathcal{U}_{J_{\beta}}$. Then, $\mathbb{F}_{\beta}(\mathcal{A})\hookrightarrow\mathbb{F}_{\beta}(\mathcal{B})$ (respectively, $\mathbb{F}_{\beta}(\mathcal{A})\preceq\mathbb{F}_{\beta}(\mathcal{B})$).
Proof.
For a successor step, Theorems 2.28 and 4.12 can be adapted to show the result, and for the limit step, we use Lemma 4.13. The case of the elementary embedding further follows from Theorem 4.7. ∎
Despite the criterion provided by the Rudin-Keisler ordering being sufficient for embeddability to be raised to ultrapowers, it is unclear what are necessary and sufficient conditions for the raising of embeddability. For example, in the proof of Theorem 4.12, $\mathcal{A}^{I}/\mathcal{U}_{I}\preceq\mathcal{A}^{J}/\mathcal{U}_{J}$ is used for the result, whereas a weaker condition, such as $\mathcal{A}^{I}/\mathcal{U}_{I}\hookrightarrow\mathcal{A}^{J}/\mathcal{U}_{J}$, would have sufficed. Nevertheless, we may find stronger criteria of embeddability between ultrapowers which do not require the embeddability of their generating structures.
For the upcoming results, we shall assume $|A|>1$, for otherwise embeddability between ultrapowers is trivial.
Definition 4.15.
Let $I$ and $J$ be sets, $H:\mathcal{P}(I)\to\mathcal{P}(J)$, and $X,Y\in\mathcal{P}(I)$. We say $H$ is:
  • (a)
    multiplicative if $H(X)\cap H(Y)=H(X\cap Y)$;
  • (b)
    additive if $H(X)\cup H(Y)=H(X\cup Y)$;
  • (c)
    subtractive if $H(I\setminus X)=H(I)\setminus H(X)$.
  • (d)
    covering if $H(I)=J$.
Proposition 4.16.
If $H:\mathcal{P}(I)\to\mathcal{P}(J)$ is subtractive and multiplicative, then $H(\varnothing)=\varnothing$.
Proof.
Since $H$ is subtractive and multiplicative, $H(\varnothing)=H(\varnothing\cap I)=H(\varnothing)\cap H(I)=\linebreak H(\varnothing)\cap\big(H(I)\setminus H(\varnothing)\big)=\varnothing$. ∎
Proposition 4.17.
Let $H:\mathcal{P}(I)\to\mathcal{P}(J)$ be additive, multiplicative, subtractive and covering. If $\bigcup_{i<\kappa}X_{i}$ is a partition of $I$, then $\bigcup_{i<\kappa}H(X_{i})$ is a partition of $J$.
Proof.
Since $X_{i}\cap X_{j}=\varnothing$ for $i\neq j$ and $H$ is multiplicative and subtractive, by Proposition 4.16, $H(X_{i})\cap H(X_{j})=H(X_{i}\cap X_{j})=H(\varnothing)=\varnothing$. Since $H$ is additive and covering, $j\in J=H(I)=H(\bigcup_{i<\kappa}X_{i})=\bigcup_{i<\kappa}H(X_{i})$. ∎
Lemma 4.18.
Let $H:\mathcal{P}(I)\to\mathcal{P}(J)$ be additive, multiplicative, subtractive and covering, and $u:A\to B$ be an injective mapping. For each $f\in A^{I}$, let $w(f)\in B^{J}$ be the function such that for each $a\in A$,
$\{j\in J\mid w(f)(j)=u(a)\}=H(\{j\in I\mid f(j)=a\})$.
Then, for $f,g\in A^{I}$,
  • $H(\{j\in I\mid f(j)=g(j)\})=\{j\in J\mid w(f)(j)=w(g)(j)\}$, and
  • $H(\{j\in I\mid f(j)\neq g(j)\})=\{j\in J\mid w(f)(j)\neq w(g)(j)\}$.
Proof.
Notice $w(f)$ is well defined because $u$ is injective, and that by Proposition 4.17, $w(f)$ is indeed unique. Furthermore, we have
$\{j\in I\mid f(j)=g(j)\}=\bigcup_{a\in A}(\{j\in I\mid f(j)=a\}\cap\{j\in I\mid g(j)=a\})$,
so that $H(\{j\in I\mid f(j)=g(j)\})=$
== $H\big(\bigcup_{a\in A}(\{j\in I\mid f(j)=a\}\cap\{j\in I\mid g(j)=a\})\big)$
== $\bigcup_{a\in A}H(\{j\in I\mid f(j)=a\}\cap\{j\in I\mid g(j)=a\})$
== $\bigcup_{a\in A}\big(H(\{j\in I\mid f(j)=a\})\cap H(\{j\in I\mid g(j)=a\})\big)$
== $\bigcup_{a\in A}(\{j\in J\mid w(f)(j)=u(a)\})\cap\{j\in J\mid w(g)(j)=u(a)\})$
== $\{j\in J\mid w(f)(j)=w(g)(j)\}$,
since $\mathrm{img}\big(w(f)\big)=\mathrm{img}\big(w(g)\big)=u[A]$. Similarly, we have
$H(\{j\in I\mid f(j)\neq g(j)\})$ == $H(I\setminus\{j\in I\mid f(j)=g(j)\})$
== $J\setminus H(\{j\in I\mid f(j)=g(j)\})$
== $J\setminus\{j\in J\mid w(f)(j)=w(g)(j)\}$
== $\{j\in J\mid w(f)(j)\neq w(g)(j)\}$.
Lemma 4.19.
Let $h:J\to I$ be a mapping and $H:\mathcal{P}(I)\to\mathcal{P}(J);X\mapsto h^{-1}[X]$. Then $H$ is additive, multiplicative, subtractive and covering.
Proof.
Clearly, $H(I)=h^{-1}[I]=J$, so $H$ is covering. To see it is multiplicative, notice
$H(X\cap Y)$ == $h^{-1}[X\cap Y]$
== $\{j\in J\mid h(j)\in X\cap Y\}$
== $\{j\in J\mid h(j)\in X\}\cap\{j\in J\mid h(j)\in Y\}$
== $h^{-1}[X]\cap h^{-1}[Y]$
== $H(X)\cap H(Y)$
By a similar argument, we may see it is additive and subtractive. ∎
Theorem 4.20.
Let $u:\mathcal{A}\to\mathcal{B}$ be injective, $h:J\to I$ be a mapping, and $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. Let also $H:X\mapsto h^{-1}[X]$ and define $w:A^{I}\to B^{J}$ as in Lemma 4.18. Define $e:\mathcal{A}^{I}/\mathcal{U}_{I}\to\mathcal{B}^{J}/\mathcal{U}_{J};[f]_{\mathcal{U}_{I}}\mapsto[w(f)]_{\mathcal{U}_{J}}$. Then $e$ is an embedding iff $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$.
Proof.
For simplicity, we write $\mathcal{A}^{\prime}$ for $\mathcal{A}^{I}/\mathcal{U}_{I}$ and $\mathcal{B}^{\prime}$ for $\mathcal{B}^{J}/\mathcal{U}_{J}$. Notice by Lemma 4.19, $H$ is multiplicative, additive, subtractive and covering, so we may indeed define $w$ as in Lemma 4.18. Now, for a constant $\mathrm{c}$, $w(\overline{\mathrm{c}^{\mathcal{A}}})=\overline{\mathrm{c}^{\mathcal{B}}}$, so $e(\mathrm{c}^{\mathcal{A}^{\prime}})=e([\overline{\mathrm{c}^{\mathcal{A}}}]_{\mathcal{U}_{I}})=[\overline{\mathrm{c}^{\mathcal{B}}}]_{\mathcal{U}_{J}}=\mathrm{c}^{\mathcal{B}^{\prime}}$. Let now $\mathrm{F}^{\mathcal{A}^{\prime}}([f_{0}]_{\mathcal{U}_{I}},...,[f_{n}]_{\mathcal{U}_{I}})=[g]_{\mathcal{U}_{I}}$. One might check that $w\big(\mathrm{F}_{1}^{\mathcal{A}}(f_{0},...,f_{n})\big)=\mathrm{F}_{1}^{\mathcal{B}}\big(w(f_{0}),...,w(f_{n})\big)$ (where $\mathrm{F}_{1}^{\mathcal{A}}$ and $\mathrm{F}_{1}^{\mathcal{B}}$ are as in Definition 2.3), so we get
$e\big(\mathrm{F}^{\mathcal{A}^{\prime}}([f_{0}]_{\mathcal{U}_{I}},...,[f_{n}]_{\mathcal{U}_{I}})\big)$ == $e\big([\mathrm{F}_{1}^{\mathcal{A}}(f_{0},...,f_{n})]_{\mathcal{U}_{I}}\big)$
== $\big[w\big(\mathrm{F}^{\mathcal{A}}_{1}(f_{0},...,f_{n})\big)\big]_{\mathcal{U}_{J}}$
== $\big[\mathrm{F}_{1}^{\mathcal{B}}\big(w(f_{0}),...,w(f_{n})\big)\big]_{\mathcal{U}_{J}}$
== $\mathrm{F}^{\mathcal{B}^{\prime}}\big([w(f_{0})]_{\mathcal{U}_{J}},...,[w(f_{n})]_{\mathcal{U}_{J}}\big)$
== $\mathrm{F}^{\mathcal{B}^{\prime}}\big(e([f_{0}]_{\mathcal{U}_{I}}),...,e([f_{n}]_{\mathcal{U}_{I}})\big)$.
Therefore, $e$ preserves the operations and constants for any choice of mappings $u$ (as long as it is injective) and $h$. We now proceed to showing the remaining of each direction of the equivalence.
($\Rightarrow$) Suppose $e$ is an embedding. For each $X\subseteq I$, choose a unique pair $f_{X},g_{X}\in A^{I}$ such that $X=\{j\in I\mid f_{X}(j)=g_{X}(j)\}$. Since $A^{I}$ contains every possible permutation, there are always such $f_{X}$ and $g_{X}$. By Lemma 4.18, $H:X\mapsto\{j\in J\mid w(f_{X})(j)=w(g_{X})(j)\}$. That means $X=\{j\in I\mid f_{X}(j)=g_{X}(j)\}\in\mathcal{U}_{I}$ iff $f_{X}]_{\mathcal{U}_{I}}=[g_{X}]_{\mathcal{U}_{I}}$ iff $w(f_{X})]_{\mathcal{U}_{J}}=[w(g_{X})]_{\mathcal{U}_{J}}$ iff $\{j\in J\mid w(f_{X})(j)=w(g_{X})(j)\}=H(X)=h^{-1}[X]\in\mathcal{U}_{J}$. Thus, $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$.
($\Leftarrow$) Suppose $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$, so that there is $h^{\prime}:J\to I$ such that $X\in\mathcal{U}_{I}$ iff $h^{\prime-1}[X]\in\mathcal{U}_{J}$. As observed, without loss of generality, we may let $h=h^{\prime}$, which means $X\in\mathcal{U}_{I}$ iff $H(X)\in\mathcal{U}_{J}$. Let now $\mathrm{R}^{\mathcal{A}^{\prime}}[f_{0}]_{\mathcal{U}_{I}}...[f_{n}]_{\mathcal{U}_{I}}$. Then $\{j\in I\mid\mathrm{R}^{\mathcal{A}}f_{0}(j)...f_{n}(j)\}\in\mathcal{U}_{I}$. For convenience, call that set $Z$, and let $k\in Z$ and $g_{0},...,g_{n}$ be such that $g_{i}(j)=f_{i}(j)$ for any $j\in Z$, and $g_{i}(j)=f_{i}(k)$ otherwise. Then, we get $\{j\in I\mid\mathrm{R}^{\mathcal{A}}g_{0}(j)...g_{n}(j)\}=I$, and $Z\subseteq\{j\in I\mid g_{i}(j)=f_{i}(j)\}\in\mathcal{U}_{I}$. That means we have $\{j\in J\mid\mathrm{R}^{\mathcal{B}}w(g_{0})(j)...w(g_{n})(j)\}=J\in\mathcal{U}_{J}$, and by Lemma 4.18 and our supposition, $H(\{j\in I\mid g_{i}(j)=f_{i}(j)\})=\{j\in J\mid w(g_{i})(j)=w(f_{i})(j)\}\in\mathcal{U}_{J}$, so that
$\{j\in J\mid\mathrm{R}^{\mathcal{B}}w(g_{0})(j)...w(g_{n})(j)\}\cap(\bigcap_{i\leq n}\{j\in J\mid w(g_{i})(j)=w(f_{i})(j)\})\subseteq\{j\in J\mid\mathrm{R}^{\mathcal{B}}w(f_{0})(j)...w(f_{n})(j)\}\in\mathcal{U}_{J}$.
Thus, $\mathrm{R}^{\mathcal{B}^{\prime}}[w(f_{0})]_{\mathcal{U}_{J}}...[w(f_{n})]_{\mathcal{U}_{J}}$, that is, that $\mathrm{R}^{\mathcal{B}^{\prime}}e\big([f_{0}]_{\mathcal{U}_{I}}\big)...e\big([f_{n}]_{\mathcal{U}_{I}}\big)$. The other direction (that is, $\mathrm{R}^{\mathcal{B}^{\prime}}e\big([f_{0}]_{\mathcal{U}_{I}}\big)...e\big([f_{n}]_{\mathcal{U}_{I}}\big)\Rightarrow\mathrm{R}^{\mathcal{A}^{\prime}}[f_{0}]_{\mathcal{U}_{I}}...[f_{n}]_{\mathcal{U}_{I}}$) follows analogously. Suppose now $f]_{\mathcal{U}_{I}}\neq[g]_{\mathcal{U}_{I}}$, so that $\{j\in I\mid f(j)\neq g(j)\}\in\mathcal{U}_{I}$. That means with Lemma 4.18 and our supposition, we have $\{j\in J\mid w(f)(j)\neq w(g)(j)\}=H(\{j\in I\mid f(j)\neq g(j)\})\in\mathcal{U}_{J}$. Therefore, $e([f]_{\mathcal{U}_{I}})=[w(f)]_{\mathcal{U}_{J}}\neq[w(g)]_{\mathcal{U}_{J}}=e([g]_{\mathcal{U}_{I}})$. ∎
Therefore, we obtain the following sufficient criterion for embeddability and isomorphism between ultrapowers:
Corollary 4.21.
Let $|A|\leq|B|$ and $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ be ultrafilters over $I$ and $J$. If $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$, then $\mathcal{A}^{I}/\mathcal{U}_{I}\hookrightarrow\mathcal{B}^{J}/\mathcal{U}_{J}$. Therefore, if $|A|=|B|$ and $\mathcal{U}_{I}\approx_{\mathrm{RK}}\mathcal{U}_{J}$, then $\mathcal{A}^{I}/\mathcal{U}_{I}\cong\mathcal{B}^{J}/\mathcal{U}_{J}$.
Proof.
Since $|A|\leq|B|$, there is an injective $u:A\to B$. Thus, choose any mapping $h:J\to I$, and the rest follows by Theorem 4.20. For the isomorphism, with the fact that $|A|=|B|$ and $\mathcal{U}_{I}\approx_{\mathrm{RK}}\mathcal{U}_{J}$, we obtain from the theorem there is an embedding in both directions. ∎
Notice, if it was the case that any embedding between ultrapowers is of the form defined in Theorem 4.20, then given $|A|\leq|B|$, $\mathcal{U}_{I}\leq_{\mathrm{RK}}\mathcal{U}_{J}$ becomes a necessary and sufficient criterion for the embeddability of $\mathcal{A}^{I}/\mathcal{U}_{I}$ into $\mathcal{B}^{J}/\mathcal{U}_{J}$. Thus, we might raise our first open problem:
  • Open Problem 1: Given that $|A|<|B|$, is any embedding $e:\mathcal{A}^{I}/\mathcal{U}_{I}\to\mathcal{B}^{J}/\mathcal{U}_{J}$ of the form $f]_{\mathcal{U}_{I}}\mapsto[w(f)]_{\mathcal{U}_{J}}$ for some injective $u:\mathcal{A}\to\mathcal{B}$ and mapping $h:J\to I$ be a mapping? If not, is there a necessary and sufficient criterion of embeddability between $\mathcal{A}^{I}/\mathcal{U}_{I}$ and $\mathcal{B}^{J}/\mathcal{U}_{J}$?
Still, the above result provides a stronger criterion of embeddability, which may be extended to the quotiented cumulative powers:
Corollary 4.22.
Let $|A|\leq|B|$ (respectively, $|A|=|B|$) and $\beta\in\mathbf{On}$, $\beta\in\mathbf{On}$, and their cumulative power hierarchies be respectively generated by $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ and $\{J_{\alpha}\}_{\alpha\in\mathbf{On}}$. For each $\alpha<\beta\in\mathbf{On}$, let $|I_{\alpha}|\leq|J_{\alpha}|$, and for each ultrafilter $\mathcal{U}_{I_{\alpha}}$ and $\mathcal{U}_{J_{\alpha}}$ respectively defining $\sim^{\mathcal{A}}_{\alpha+1}$ and $\sim^{\mathcal{B}}_{\alpha+1}$, let $\mathcal{U}_{I_{\alpha}}\leq_{\mathrm{RK}}\mathcal{U}_{J_{\alpha}}$ (respectively, $\mathcal{U}_{I_{\alpha}}\approx_{\mathrm{RK}}\mathcal{U}_{J_{\alpha}}$). Then, $\mathbb{F}_{\beta}(\mathcal{A})\hookrightarrow\mathbb{F}_{\beta}(\mathcal{B})$ (respectively, $\mathbb{F}_{\beta}(\mathcal{A})\cong\mathbb{F}_{\beta}(\mathcal{B})$).
Proof.
By induction on $\beta$. As with Corollary 4.14, for a successor step, we adapt Theorem 4.12, and for the limit step, we use Lemma 4.13, with the case of the elementary embedding following from Theorem 4.7. ∎
As we have noted, the present construction holds many similarities to ultrapowers. Indeed, just as we have shown $\mathfrak{F}_{\beta}(\mathcal{A})$ is isomorphic to a direct power, we shall now show $\mathbb{F}_{\beta}(\mathcal{A})$ is as versatile as an ultrapower. For that, we start with some definitions.
Definition 4.23 (Ultrapower hierarchy).
Let $\mathcal{A}$ be a $\sigma$-structure, $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$ a family of index sets, and $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ a family of ultrafilters, where each $\mathcal{U}_{\alpha}$ is an ultrafilter over $I_{\alpha}$. The ultrapower hierarchy generated by those parameters is defined as (for each $\alpha^{+},\lambda<\beta$):
  • $\Upsilon_{0}(\mathcal{A})=\mathcal{A}$;
  • $\Upsilon_{\alpha^{+}}(\mathcal{A})=\Upsilon_{\alpha}(\mathcal{A})^{I_{\alpha}}/\mathcal{U}_{\alpha}$;
  • $\Upsilon_{\lambda}(\mathcal{A})=\bigcup_{\alpha<\lambda}\Upsilon_{\alpha}(\mathcal{A})/\!\approx_{\lambda}$ is the limit of the elementary chain, for a limit $\lambda$, with the embeddings inductively defined by
    • $e^{\alpha}_{\alpha}$ is the identity function,
    • $e^{\alpha}_{\beta^{+}}:\Upsilon_{\alpha}(\mathcal{A})\to\Upsilon_{\beta^{+}}(\mathcal{A});x\mapsto[\overline{e^{\alpha}_{\beta}(x)}]_{\mathcal{U}_{\beta}}$,
    • $e^{\alpha}_{\lambda}:\Upsilon_{\alpha}(\mathcal{A})\to\Upsilon_{\lambda}(\mathcal{A});x\mapsto[x]_{\mathcal{U}_{\lambda}}$, for a limit $\lambda$,
    where $\cdot]_{\mathcal{U}_{\beta}}$ denotes the equivalence class modulo $\mathcal{U}_{\beta}$ for $\beta^{+}$, and $\cdot]_{\mathcal{U}_{\lambda}}$, for a limit $\lambda$, the equivalence class defined by the relation $x\approx_{\lambda}y$, iff there are $\alpha\leq\beta<\lambda$ such that either $e^{\alpha}_{\beta}(x)=y$ or $e^{\alpha}_{\beta}(y)=x$. [^14] The subscripts shall become relevant when we compare this construction with our former one, so that $\cdot]_{\mathcal{U}_{\alpha}}$ denotes an equivalence class of $\mathcal{A}_{\alpha}$, and $\cdot]_{\alpha}$, one of $\mathbb{F}_{\alpha}(\mathbb{A})$. For that reason, we write $\cdot]_{\mathcal{U}_{\lambda}}$, despite there not being an ultrafilter $\mathcal{U}_{\lambda}$ defining $\mathcal{A}_{\lambda}$. When the ordinal is unambiguous, we may write simply $\cdot$].
For $a\in\bigcup_{\alpha\leq\beta}\Upsilon_{\alpha}(\mathcal{A})$, let $\rho(a)=min\{\alpha\mid\exists x\in\Upsilon_{\alpha}(\mathcal{A})(x\approx a)\}$. We might recall the interpretation of the non-logical symbols in an ultrapower is given by (for $\delta=max\{\rho(a_{1}),...,\rho(a_{n})\}=\linebreak min\{\alpha<\lambda\mid\forall x\in B(x\cap\Upsilon_{\alpha}(\mathcal{A})\neq\varnothing)\}$):
  • $\mathrm{F}_{i}^{\Upsilon_{\beta^{+}}(\mathcal{A})}$ maps $\langle[a_{1}],...,[a_{n}]\rangle\mapsto[f]$, where $f(x)=\mathrm{F}_{i}^{\Upsilon_{\beta}(\mathcal{A})}\big(a_{1}(x),...,a_{n}(x)\big)$ for any $x\in\Upsilon_{\beta}(\mathcal{A})$, and for a limit $\lambda$, $\mathrm{F}_{i}^{\Upsilon_{\lambda}(\mathcal{A})}$ maps $\langle[a_{1}],...,[a_{n}]\rangle\mapsto\big[\mathrm{F}_{i}^{\Upsilon_{\delta}(\mathcal{A})}\big(e^{\rho(a_{1})}_{\delta}(a_{1}),...,e^{\rho(a_{n})}_{\delta}(a_{n})\big)\big]$;
  • $\mathrm{R}_{j}^{\Upsilon_{\beta^{+}}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\{x\in\Upsilon_{\alpha}(\mathcal{A})\mid\mathrm{R}_{j}^{\Upsilon_{\beta}(\mathcal{A})}a_{1}(x)...a_{n}(x)\}\in\mathcal{U}_{\beta}$, and for a limit $\lambda$, $\mathrm{R}_{j}^{\Upsilon_{\lambda}(\mathcal{A})}[a_{1}]...[a_{n}]$ iff $\mathrm{R}_{j}^{\Upsilon_{\delta}(\mathcal{A})}e^{\rho(a_{1})}_{\delta}(a_{1})...e^{\rho(a_{n})}_{\delta}(a_{1})$;
  • $\mathrm{c}_{k}^{\Upsilon_{\beta^{+}}(\mathcal{A})}=[\overline{\mathrm{c}_{k}^{\Upsilon_{\beta}(\mathcal{A})}}]$, and for a limit $\lambda$, $\mathrm{c}_{k}^{\Upsilon_{\lambda}(\mathcal{A})}=[\mathrm{c}^{\mathcal{A}}]$.
By replacing above $\beta$ for $\mathbf{On}$, we define the whole ultrapower hierarchy and the class $\Upsilon_{\mathbf{On}}(\mathcal{A})$. [^15] Notice $\Upsilon_{\mathbf{On}}(\mathcal{A})$ will not be a proper class only if either $\mathcal{A}$ is finite, or if there is a cardinal equal to the number of ultrafilters $\mathcal{U}_{\gamma}$ in $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ such that $\mathcal{U}_{\gamma}$ is non-principal. For any $a]\in\Upsilon_{\mathbf{On}}(\mathcal{A})$, we may say $a]\in\Upsilon_{\beta}(\mathcal{A})$ when $a]\cap\Upsilon_{\beta}(\mathcal{A})\neq\varnothing$. Similarly, for $X\subseteq\Upsilon_{\mathbf{On}}(\mathcal{A})$, we may say $X\subseteq\Upsilon_{\beta}(\mathcal{A})$ when for each $a]\in X$, $a]\cap\Upsilon_{\beta}(\mathcal{A})\neq\varnothing$.
We might note this is a stepwise, and sometimes direct, construction of iterated ultrapowers, and in fact when considering the finite stages of the construction, coincides with finitely iterated ultrapowers. [^16] Developed for finite iterations by [8], and for infinite iterations by [9] ([3], Chapter 6). Notice, however, that despite the use of proper classes for the definition of the ultrapower hierarchy as a whole, the present construction is somewhat less reliant on proper classes. Whereas in the analogous iterated ultrapower construction of the proper class sized $\mathcal{A}^{I_{\mathbf{On}}}/\mathcal{U}_{\mathbf{On}}$, where $I_{\mathbf{On}}=\bigtimes_{\alpha<\mathbf{On}}I_{\alpha}$ and $\mathcal{U}_{\mathbf{On}}=\bigtimes_{\alpha<\mathbf{On}}\mathcal{U}_{\alpha}$, any statement about its elements involves the consideration of its proper equivalence classes, any statement about a set sized $X\subseteq\Upsilon_{\mathbf{On}}(\mathcal{A})$ is equivalent to a statement about $\Upsilon_{\delta}(\mathcal{A})$, where $\delta=min\{\alpha\in\mathbf{On}\mid\forall x\in X(x\cap\Upsilon_{\alpha}(\mathcal{A})\neq\varnothing)\}$ – that is, the first stage at which each element of $X$ has a representative.
Lemma 4.24.
Let $\lambda$ be a limit ordinal. If for all $\beta<\lambda$, $\mathbb{F}_{\beta}(\mathcal{A})\cong\Upsilon_{\beta}(\mathcal{A})$, then $\mathbb{F}_{\lambda}(\mathcal{A})\cong\Upsilon_{\lambda}(\mathcal{A})$.
Proof.
Let $\theta_{\beta}:\mathbb{F}_{\beta}(\mathcal{A})\to\mathcal{F}_{\beta}(A)$ be a choice function, for each $\beta<\lambda$. Then, consider $e^{\prime}:\Upsilon_{\lambda}(\mathcal{A})\to\mathbb{F}_{\lambda}(\mathcal{A});[x]_{\mathcal{U}_{\lambda}}\mapsto[\theta_{\delta}\big(e_{\delta}([x]_{\mathcal{U}_{\delta}})\big)]_{\lambda}$, where $\delta$ is the stage of the ultrapower hierarchy at which $x$ is generated, that is, $x\in\Upsilon_{\delta-1}(\mathcal{A})^{I_{\delta-1}}$, and $e_{\delta}$ is the isomorphism between $\Upsilon_{\delta}(\mathcal{A})$ and $\mathbb{F}_{\delta}(\mathcal{A})$. Then, it is not hard to check $e^{\prime}$ is an isomorphism. ∎
Theorem 4.25.
Let a cumulative power hierarchy be generated by $\mathcal{A}$ and the family of index sets $\{I_{\alpha}\}_{\alpha\in\mathbf{On}}$. Let an ultrapower hierarchy similarly be generated by the same structure and family of index set, and the family of ultrafilters $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ respectively over them. For each $\gamma\in\mathbf{On}$, if for all $\beta<\gamma$, $\sim_{\beta^{+}}$ is defined by $\mathcal{U}_{\beta}$, then $\mathbb{F}_{\gamma}(\mathcal{A})\cong\Upsilon_{\gamma}(\mathcal{A})$.
Proof.
By induction on $\beta$. The base and successor cases are trivial, and the limit case is covered by Lemma 4.24. ∎
Therefore, just as it was the case with direct powers, an ultrapower may be seen to be a structure emerging from quotienting a cumulative power by an appropriately defined equivalence relation.
[…]
The answer to the former question depends on the ultrafilter of choice. For example, if $\mathcal{A}$ is an ordered structure and $\mathcal{U}$ is non-principal, then $\mathcal{A}^{I}/\mathcal{U}$ has elements which dominate the image of the natural embedding. However, no matter the choice of representatives, there is no element in the direct power $\mathcal{A}^{I}$ which assumes a value greater than the values of every representative of the equivalence classes of the image of the natural embedding. On the other hand, if $\mathcal{U}$ is principal, the embedding is trivial.
To try to answer that, we start by considering the functions of a structure. Notice the issue with the direct power not being a field is there being functions with 0 in their image, but which are distinct from $\overline{0}$. So, if one wants to retain multiplicative inverses, one should exclude any function containing elements without a multiplicative inverse in its image, that is, any function whose support is not its whole domain. However, for the desired isomorphism, there should still be a representative of each element not possessing multiplicative inverses. To generalize that for arbitrary structures, we start with the following definitions.
For the next results, we let $\sigma$ be a signature, $\mathcal{A}$ be a $\sigma$-structure, $I$ be an index set, and $\mathcal{U}$ be an ultrafilter over it.
Definition 4.26 (Support).
Define the set $\mathrm{supp}(\mathcal{A})$, the support of $\mathcal{A}$, such that
$a\in\mathrm{supp}(\mathcal{A})$ iff $\mathrm{F}(a,...,x_{\mathrm{ar}(\mathrm{F})}),...,\mathrm{F}(x_{1},...,a,...x_{\mathrm{ar}(\mathrm{F})}),...,\mathrm{F}(x_{1},...,a)$ are defined for all $x_{1},...,x_{\mathrm{ar}(\mathrm{F})}\in A$ and $\mathrm{F}\in\sigma$.
Let $\overline{\mathrm{supp}}(\mathcal{A})=A\setminus\mathrm{supp}(\mathcal{A})$, the anti-support of $\mathcal{A}$. Define
$\mathrm{rep}(\mathcal{A}^{I})=\{\overline{a}\in A^{I}\mid a\in\overline{\mathrm{supp}}(\mathcal{A})\}\cup\{f\in A^{I}\mid\forall j\in I(f(j)\in\mathrm{supp}(\mathcal{A}))\}$.
Similarly, for $\gamma\in\mathbf{On}$ and $f\in\mathcal{F}_{\gamma}(A)$, let
$\mathrm{supp}(f)=\{j\in\mathrm{dom}(f)\mid[f(j)]_{\gamma}\cap\overline{\mathrm{supp}}\big(\mathcal{F}_{\gamma}(\mathcal{A})\big)=\varnothing\}$, [^17] That is, $\mathrm{supp}(f)$ is the set of elements $j$ in the domain of $f$ such that $f(j)\not\sim_{\gamma}b$ for any $b\in\overline{\mathrm{supp}}\big(\mathcal{F}_{\gamma}(\mathcal{A})\big)$.
and define
$\mathrm{rep}_{\beta^{+}}(\mathcal{A})=\{\overline{a}\in\mathcal{F}^{*}_{\beta^{+}}(A)\mid a\in\overline{\mathrm{supp}}\big(\mathcal{F}_{\beta}(A)\big)\}\cup\{f\in\mathcal{F}^{*}_{\beta^{+}}(A)\mid\mathrm{supp}(f)=I_{\beta}\}$.
The above definitions generalize the notions of support, and its counterpart, of a structure, and of a function. $\mathrm{rep}(\mathcal{A})$ is divided following the reasoning that the elements of $\mathcal{A}^{I}/\mathcal{U}$ who are in its support should find representatives among the functions in the support of $\mathcal{A}^{I}$ (which are the functions whose support are their whole domains), and elements not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should find representatives among the constant functions whose images are singletons of elements who are not in the support of their domain – intuitively, reflecting also the fact that each element not in the support of $\mathcal{A}^{I}/\mathcal{U}$ should be equivalent, in a sense, to some element in the anti-support of $\mathcal{A}$. In essence, the definition of $\mathrm{rep}(\mathcal{A})$ captures the intuition the ultrapower should not multiply the elements in the anti-support of its generating structure.
Proposition 4.27 ([3], p. 228).
$\mathcal{U}$ is $\kappa$-complete iff for every partition of $I$ into fewer than $\kappa$ parts, one of the parts belongs to $\mathcal{U}$.
Theorem 4.28.
Every element of $\mathcal{A}^{I}/\mathcal{U}$ has a representative in $\mathrm{rep}(\mathcal{A}^{I})$ iff $\mathcal{U}$ is $\kappa^{+}$-complete, where $\kappa=|\overline{\mathrm{supp}}(\mathcal{A})|$.
Proof.
This proof is similar to the one for Proposition 4.2.4 of [3]. ($\Leftarrow$) Let $f\in A^{I}$. If $\mathrm{supp}(f)=I$, the conclusion is straightforward, so suppose otherwise. Now, the sets $\mathrm{supp}(f)$ and $\{j\in I\mid f(j)=a\}_{a\in\overline{\mathrm{supp}}(\mathcal{A})}$ partition $I$. Since $\mathcal{U}$ is $\kappa^{+}$-complete for $\kappa=|\overline{\mathrm{supp}}(\mathcal{A})|$, by Proposition 4.27 exactly one of those sets is in $\mathcal{U}$. But then either $f]=[\overline{b}]$ for some $b\in\overline{\mathrm{supp}}(\mathcal{A})$, or $\mathrm{supp}(f)\in\mathcal{U}$, so consider $g\in A^{I}$ such that for some $c\in\mathrm{supp}(\mathcal{A})$,
$g(j)=\begin{cases}f(j),\ \text{if}\ j\in\mathrm{supp}(\mathcal{A})\\ c,\ \text{otherwise}\end{cases}$
so that $[g]=[f]$. Either way, we have the conclusion, since $\overline{b},g\in\mathrm{rep}(\mathcal{A}^{I})$.
($\Rightarrow$) Suppose every $f]\in\mathcal{A}^{I}/\mathcal{U}$ has a representative in $\mathrm{rep}(\mathcal{A}^{I})$. Let $\bigcup_{\alpha<\xi}X_{\alpha}$ be a partition of $A$, for $\xi\leq\kappa=|\overline{\mathrm{supp}}(\mathcal{A})|$. Let $\langle a_{\alpha}\rangle_{\alpha<\kappa}$ be an enumeration of $\overline{\mathrm{supp}}(\mathcal{A})$ and $f\in A^{I}$ be such that $f(j)=a_{\zeta}$ iff $j\in X_{\zeta}$. By $f$’s construction, $f]\neq[b]$ for any $b\in\mathrm{supp}(\mathcal{A})$, so by assumption we must have $f]=[\overline{c}]$ for some $c\in\overline{\mathrm{supp}}(\mathcal{A})$. But by construction, $c=a_{\zeta}$ for some $\zeta<\xi$, which means $\{j\in I\mid f(j)=c\}=\{j\in I\mid f(j)=a_{\zeta}\}=X_{\zeta}\in\mathcal{U}$. By the arbitrariness of the partition and Proposition 4.27, $\mathcal{U}$ is $\kappa^{+}$-complete. ∎
Corollary 4.29.
Every element of $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ has a representative in $\mathrm{rep}_{\beta^{+}}(\mathcal{A})$ iff for all $\alpha\leq\beta$ the ultrafilter $\mathcal{U}_{\alpha}$ defining $\sim_{\alpha^{+}}$ is $\kappa^{+}$-complete, where $\kappa=|\overline{\mathrm{supp}}\big(\mathcal{F}_{\alpha}(\mathcal{A})\big)|$.
Proof sketch. By induction on $\beta$, using Theorem 4.28.
Every ultrafilter is finitely complete, which justifies the set $\mathrm{rep}(\mathcal{A})$ always containing some representative of every equivalence class of $\mathcal{A}^{I}/\mathcal{U}$ when $|\overline{\mathrm{supp}}\big(\mathcal{A})\big)|<\aleph_{0}$. However, as it is well known, the existence of $\kappa$-complete ultrafilters, for $\kappa>\aleph_{0}$, is independent of ZFC, which means the existence of such a fragment in $\mathrm{rep}(\mathcal{A})$, within ZFC, is only guaranteed for structures with $|\overline{\mathrm{supp}}(\mathcal{A})|<\aleph_{0}$:
Corollary 4.30.
Let $|\overline{\mathrm{supp}}(\mathcal{A})|\geq\aleph_{0}$. If $\mathcal{A}^{I}/\mathcal{U}$ is non-isomorphic to $\mathcal{A}$ (equivalently, if $\mathcal{U}$ is non-principal), then $\mathrm{rep}(\mathcal{A}^{I})$ contains a representative of every equivalence class of $\mathcal{A}^{I}/\mathcal{U}$ iff there is a measurable cardinal $\kappa>|\overline{\mathrm{supp}}(\mathcal{A})|$.
Nevertheless, the result also implies, within ZFC, that if $\mathcal{A}$ is a field (and, in the presence of additional operations $\{\mathrm{F}_{i}\}_{i\leq\kappa}$, as long as $|A^{\mathrm{ar}(\mathrm{F_{i}})}\setminus\mathrm{dom}(\mathrm{F}_{i})|<\aleph_{0}$), then $\mathrm{rep}(\mathcal{A})$ does contain a representative of each element of $\mathcal{A}^{I}/\mathcal{U}$.
For a $\sigma$-structure $\mathcal{A}$ and $a\in A$, let $\tau_{a}$ be the cardinality of all the tuples either from which $a$ may be obtained by means of an operation, or containing $a$ for which some function is defined, or containing $a$ for which some relation holds – that is, for
$S^{a\downarrow}_{\mathrm{F}}$ == $\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in A^{\mathrm{ar}(\mathrm{F})}\mid\mathrm{F}^{\mathcal{A}}(b_{1},...,b_{\mathrm{ar}(\mathrm{F})})=a\}$,
$S^{a\uparrow}_{\mathrm{F}}$ == $\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in A^{\mathrm{ar}(\mathrm{F})}\mid\mathrm{F}^{\mathcal{A}}(b_{1},...,a,...,b_{\mathrm{ar}(\mathrm{F})-1})=b_{\mathrm{ar}(\mathrm{F})}\}$, and
$S^{a}_{\mathrm{R}}$ == $\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R})-1}\rangle\in A^{\mathrm{ar}(\mathrm{R})-1}\mid\mathrm{R}^{\mathcal{A}}b_{1}...a...b_{\mathrm{ar}(\mathrm{R})-1}\}$,
we let $\tau_{a}=\big|\bigsqcup_{\mathrm{F}\in\sigma}(S^{a\downarrow}_{\mathrm{F}}\sqcup S^{a\uparrow}_{\mathrm{F}})\sqcup\bigsqcup_{\mathrm{R}\in\sigma}S^{a}_{\mathrm{R}}\big|=\sum_{\mathrm{F}\in\sigma}(|S^{a\downarrow}_{\mathrm{F}}|+|S^{a\uparrow}_{\mathrm{F}}|)+\sum_{\mathrm{R}\in\sigma}|S^{a}_{\mathrm{R}}|$. Define $\tau_{\mathcal{A}}$ as the greatest of the $\tau_{a}$ for all $a\in A$, that is, $\tau_{\mathcal{A}}=\mathrm{max}\{\tau_{a}\}_{a\in A}$. We may now answer our initial question.
Theorem 4.31.
There is a choice function $e$ on the equivalence classes of $A^{I}/\mathcal{U}$ such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff $\mathcal{U}$ is $\tau_{\mathcal{A}^{I}/\mathcal{U}}^{+}$-complete.
Proof.
($\Leftarrow$) Call $\mathcal{B}=\mathcal{A}^{I}/\mathcal{U}$ and let $\theta:B\to A^{I}$ be a choice function on the equivalence classes. Let $\langle a_{\alpha}\rangle_{\alpha<|B\setminus[A]_{\mathcal{U}}|}$ be an enumeration of the elements of $B\setminus[A]$, where $[A]$ is the image of the natural embedding of $\mathcal{A}$. We inductively define the following relation. Let $e_{0}=\{\langle[\overline{x}],\overline{x}\rangle\mid x\in A\}$. Then, for each $a\in B$, for $e_{<\beta}=\bigcup_{\alpha<\beta}e_{\alpha}$, let
$a^{\dagger}=\begin{cases}e_{<\beta}(a),\ \text{if}\ a\in\mathrm{dom}(e_{<\beta})\\ \theta(a),\ \text{otherwise}\end{cases}$
In that way, for each $\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in S^{a\downarrow}_{\mathrm{F}}$,
$\big\{j\in I\mid\mathrm{F}^{\mathcal{A}}\big(b_{1}^{\dagger}(j),...,b^{\dagger}_{\mathrm{ar}(\mathrm{F})}(j)\big)=a^{\dagger}(j)\big\}\in\mathcal{U}$;
similarly, for each $\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in S^{a\uparrow}_{\mathrm{F}}$,
$\big\{j\in I\mid\mathrm{F}^{\mathcal{A}}\big(b_{1}^{\dagger}(j),...,a^{\dagger}(j),...,b_{\mathrm{ar}(\mathrm{F})-1}^{\dagger}(j)\big)=b_{\mathrm{ar}(\mathrm{F})}^{\dagger}(j)\big\}\in\mathcal{U}$,
and for each $\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R}-1)}\rangle\in S^{a}_{\mathrm{R}}$,
$\big\{j\in I\mid\mathrm{R}^{\mathcal{A}}b_{1}^{\dagger}(j)...a^{\dagger}(j)...b_{\mathrm{ar}(\mathrm{F})-1}^{\dagger}(j)\big\}\in\mathcal{U}$.
Since $\tau_{\mathcal{B}}\geq\big|\bigsqcup_{\mathrm{F}\in\sigma}(S^{a\downarrow}_{\mathrm{F}}\sqcup S^{a\uparrow}_{\mathrm{F}})\sqcup\bigsqcup_{\mathrm{R}\in\sigma}S^{a}_{\mathrm{R}}\big|$ and $\mathcal{U}$ is $\tau^{+}_{\mathcal{B}}$-complete,
$\bigcap_{\mathrm{F}\in\sigma}\big(\bigcap_{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in S_{\mathrm{F}}^{a\downarrow}}\big\{j\in I\mid\mathrm{F}^{\mathcal{A}}\big(b_{1}^{\dagger}(j),...,b^{\dagger}_{\mathrm{ar}(\mathrm{F})}(j)\big)=a^{\dagger}(j)\big\}\ \cap\ \linebreak\bigcap_{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in S_{\mathrm{F}}^{a\uparrow}}\big\{j\in I\mid\mathrm{F}^{\mathcal{A}}\big(b_{1}^{\dagger}(j),...,a^{\dagger}(j),...,b_{\mathrm{ar}(\mathrm{F})-1}^{\dagger}(j)\big)=b_{\mathrm{ar}(\mathrm{F})}^{\dagger}(j)\big\}\big)\cap\linebreak\bigcap_{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})-1}\rangle\in S_{\mathrm{R}}^{a}}\big\{j\in I\mid\mathrm{R}^{\mathcal{A}}b_{1}^{\dagger}(j)...a^{\dagger}(j)...b_{\mathrm{ar}(\mathrm{R})-1}^{\dagger}(j)\big\}\in\mathcal{U}$.
Call the above set $X^{a}$ and let $k\in X^{a}$. Define then $f_{a}$ such that for each $j\in I$,
$f_{a}(j)=\begin{cases}\theta(a)(j),\ \text{if}\ j\in X^{a}\\ \theta(a)(k),\ \text{otherwise}\end{cases}$
Then, we let $e_{\beta}=\bigcup_{\alpha<\beta}e_{\alpha}\cup\{\langle a,f_{a}\rangle\}$.
From the above construction, it is clear that $e=\bigcup_{\alpha<|B\setminus[A]_{\mathcal{U}}|}e_{\alpha}$ is a function, that it preserves the constants, functions and relations of $\sigma$, and that for any $a\in B$, $[e(a)]=a$.
($\Rightarrow$) Let $e:\mathcal{B}\hookrightarrow\mathcal{A}^{I}$ be the embedding, $\kappa\leq\tau_{\mathcal{B}}$, $\{X_{\alpha}\}_{\alpha<\kappa}\subseteq\mathcal{U}$, and $a\in B$ be the element such that $\tau_{a}=\tau_{\mathcal{B}}$. Let also $u:\{X_{\alpha}\}_{\alpha<\kappa}\to\bigsqcup_{\mathrm{F}\in\sigma}(S^{a\downarrow}_{\mathrm{F}}\sqcup S^{a\uparrow}_{\mathrm{F}})\sqcup\bigsqcup_{\mathrm{R}\in\sigma}S^{a}_{\mathrm{R}}$ be an injection and $\chi:\mathcal{P}(A)\to A$ be a choice function. For each $i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]$, let $f_{i}\in A^{I}$ be such that for each $j\in I$,
$f_{i}(j)=\begin{cases}e(a)(j),\ \text{if}\ j\in u^{-1}(i)\\ \chi\big(A\setminus\{\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(b_{\mathrm{ar}(\mathrm{F})})(j)\big)\}\big),\ \text{otherwise}\end{cases}$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\downarrow}_{\mathrm{F}}}$,
$f_{i}(j)=\begin{cases}e(b_{\mathrm{ar}(\mathrm{F})})(j),\ \text{if}\ j\in u^{-1}(i)\\ \chi\big(A\setminus\{\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(a)(j),...,e(b_{\mathrm{ar}(\mathrm{F})-1})(j)\big)\}\big),\ \text{otherwise}\end{cases}$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\uparrow}_{\mathrm{F}}}$, and
$f_{i}(j)=\begin{cases}e(a)(j),\ \text{if}\ j\in u^{-1}(i)\\ \chi\big(A\setminus\{x\mid\mathrm{R}^{\mathcal{A}}e(b_{1})(j)...x...e(b_{\mathrm{ar}(\mathrm{F})-1})(j)\}\big),\ \text{if that set is non-empty}\\ \chi(A),\ \text{otherwise}\end{cases}$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R})-1}\rangle_{S^{a}_{\mathrm{R}}}$. Since $e$ is an embedding which is a choice function on the equivalence classes, by construction, that means if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{\mathrm{S}^{a\downarrow}_{\mathrm{F}}}$, then
$f_{i}(j)=\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(b_{\mathrm{ar}(\mathrm{F})})(j)\big)$ iff $j\in u^{-1}(i)$,
and thus
$f_{i}]=a=\mathrm{F}^{\mathcal{A}^{I}/\mathcal{U}}\big(b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\big)$;
similarly, if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{\mathrm{S}^{a\uparrow}_{\mathrm{F}}}$, then
$f_{i}(j)=\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(a)(j),...,e(b_{\mathrm{ar}(\mathrm{F})-1})(j)\big)$ iff $j\in u^{-1}(i)$,
and so
$f_{i}]=b_{\mathrm{ar}(\mathrm{F})}=\mathrm{F}^{\mathcal{A}^{I}/\mathcal{U}}\big(b_{1},...,a,...,b_{\mathrm{ar}(\mathrm{F})-1}\big)$;
and if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R})-1}\rangle_{\mathrm{S}^{a}_{\mathrm{R}}}$, then
if $j\in u^{-1}(i)$, then $\mathrm{R}^{\mathcal{A}}e(b_{1})(j)...f_{i}(j)...e(b_{\mathrm{ar}(\mathrm{R})-1})(j)$,
and so
$\mathrm{R}^{\mathcal{A}^{I}/\mathcal{U}}b_{1}...[f_{i}]...b_{\mathrm{ar}(\mathrm{R})-1}$.
For $i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]$ and $j\in I$, let $\varphi\big(e(i)(j)\big)$ stand for
$f_{i}(j)=\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(b_{\mathrm{ar}(\mathrm{F})})(j)\big)$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\downarrow}_{\mathrm{F}}}$,
$f_{i}(j)=\mathrm{F}^{\mathcal{A}}\big(e(b_{1})(j),...,e(a)(j),...,e(b_{\mathrm{ar}(\mathrm{F})-1})(j)\big)$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\uparrow}_{\mathrm{F}}}$, and
$\mathrm{R}^{\mathcal{A}}e(b_{1})(j)...f_{i}(j)...e(b_{\mathrm{ar}(\mathrm{R})-1})(j)$
if $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R})-1}\rangle_{S^{a}_{\mathrm{R}}}$. Let $i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]$. If $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\downarrow}_{\mathrm{F}}}$, we have $f_{i}]=[e(a)]$, and therefore $\{j\in I\mid f_{i}(j)=e(a)(j)\}\in\mathcal{U}$, once again since $e$ is an embedding which is a choice function on the equivalence classes. However,
$k\in\{j\in I\mid f_{i}(j)=e(a)(j)\}$ iff $\bigwedge_{i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]}\varphi\big(e(i)(k)\big)$,
and
$\bigwedge_{i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]}\varphi\big(e(i)(k)\big)$ iff $k\in\bigcap_{i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]}\{j\in I\mid\varphi\big(e(i)(j)\big)\}=\bigcap_{i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]}u^{-1}(i)$.
Therefore, $\bigcap_{i\in u[\{X_{\alpha}\}_{\alpha<\kappa}]}u^{-1}(i)=\bigcap_{\alpha<\kappa}X_{\alpha}\in\mathcal{U}$. If $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle_{S^{a\uparrow}_{\mathrm{F}}}$ (which we may need in case $S^{a\uparrow}_{\mathrm{F}}=S^{a}_{\mathrm{R}}=\varnothing$ for every $\mathrm{F},\mathrm{R}\in\sigma$), the same argument applies since $\{j\in I\mid f_{i}(j)=e(b_{\mathrm{ar}(\mathrm{F})})(j)\}\in\mathcal{U}$. If $i=\langle b_{1},...,b_{\mathrm{ar}(\mathrm{R})-1}\rangle_{S^{a}_{\mathrm{R}}}$ (which we may need in case $S^{a\uparrow}_{\mathrm{F}}=S^{a\downarrow}_{\mathrm{F}}=\varnothing$ for every $\mathrm{F}\in\sigma$) a similar reasoning may be employed, for then $e(a)]=[f_{i}]$. ∎
Proposition 4.32.
Let $\mathcal{A}$ be a $\sigma$-structure whose relations are empty. Let also $\mathrm{m}\in A$ be the element such that $\tau_{\mathcal{A}}=\tau_{\mathrm{m}}$, and $Z\in\mathcal{U}$ be such that $|Z|=\mathrm{min}\{|X|\}_{X\in\mathcal{U}}$. Then:
  • (a)
    $\tau_{\mathcal{A}^{I}}\ =\ \sum_{\mathrm{F}\in\sigma}|S^{\mathrm{m}\downarrow}_{\mathrm{F}}|^{|I|}+|S^{\mathrm{m}\uparrow}_{\mathrm{F}}|^{|I|}\ \leq\ (\tau_{\mathcal{A}})^{|I|}$;
  • (b)
    $\tau_{\mathcal{A}^{I}/\mathcal{U}}=$ $\mathrm{max}\big\{\sum_{\mathrm{F}\in\sigma}\big(\prod_{j\in Z}|S^{a(j)\downarrow}_{\mathrm{F}}|+\prod_{j\in Z}|S^{a(j)\uparrow}_{\mathrm{F}}|\big)\big\}_{a\in A^{I}}\$$\leq(\tau_{\mathcal{A}})^{|Z|}$.
Proof.
(a): Let $a\in\mathcal{A}^{I}$. Then
$S^{a\downarrow}_{\mathrm{F}}$ == $\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in(\mathcal{A}^{I})^{\mathrm{ar}(\mathrm{F})}\mid\mathrm{F}^{\mathcal{A}^{I}}(b_{1},...,b_{\mathrm{ar}(\mathrm{F})})=a\}$
== $\big\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in(\mathcal{A}^{I})^{\mathrm{ar}(\mathrm{F})}\mid\forall j\in I\big(\mathrm{F}^{\mathcal{A}}\big(b_{1}(j),...,b_{\mathrm{ar}(\mathrm{F})}(j)\big)=a(j)\big)\big\}$
== $\big\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in(\mathcal{A}^{I})^{\mathrm{ar}(\mathrm{F})}\mid\forall j\in I\big(\langle b_{1}(j),...,b_{\mathrm{ar}(\mathrm{F})}(j)\rangle\in S^{a(j)\downarrow}_{\mathrm{F}}\big)\big\}$
and likewise for $S^{a\uparrow}_{\mathrm{F}}$. Thus,
[…]
(and likewise for $|S^{a\uparrow}_{\mathrm{F}}|$). Let $\mathrm{m}\in A$ the element such that $\tau_{\mathcal{A}}=\tau_{\mathrm{m}}=\sum_{\mathrm{F}\in\sigma}|S_{\mathrm{F}}^{\mathrm{m}\downarrow}|+|S_{\mathrm{F}}^{\mathrm{m}\uparrow}|$. We may notice that
$\tau_{\mathcal{A}^{I}}$ == $\tau_{\overline{\mathrm{m}}}$
== $|\bigsqcup_{\mathrm{F}\in\sigma}S^{\overline{\mathrm{m}}\downarrow}_{\mathrm{F}}\sqcup S^{\overline{\mathrm{m}}\uparrow}_{\mathrm{F}}|$
== $\sum_{\mathrm{F}\in\sigma}\big(\Pi_{j\in I}|S^{\mathrm{m}\downarrow}_{\mathrm{F}}|+\Pi_{j\in I}|S^{\mathrm{m}\uparrow}_{\mathrm{F}}|\big)$
== $\sum_{\mathrm{F}\in\sigma}|S^{\mathrm{m}\downarrow}_{\mathrm{F}}|^{|I|}+|S^{\mathrm{m}\uparrow}_{\mathrm{F}}|^{|I|}\ \leq\ \big(\sum_{\mathrm{F}\in\sigma}|S^{\mathrm{m}\downarrow}_{\mathrm{F}}|+|S^{\mathrm{m}\uparrow}_{\mathrm{F}}|\big)^{|I|}\ =\ (\tau_{\mathcal{A}})^{|I|}$.
(b): Let $a\in\mathcal{A}^{I}/\mathcal{U}$ and $\theta:\mathcal{A}^{I}/\mathcal{U}\to A^{I}$ be a choice function on the equivalence classes. Then, as we have seen,
[…]
$\big|\big\{\langle b_{1},...,b_{\mathrm{ar}(\mathrm{F})}\rangle\in(A^{I})^{\mathrm{ar}(\mathrm{F})}\mid\{j\in I\mid\langle b_{1}(j),...,b_{\mathrm{ar}(\mathrm{F})}(j)\rangle\in S^{\theta(a)(j)\downarrow}_{\mathrm{F}}\}=X\big\}\big|$ ==
[…]
Thus, we may see, if $X\in\mathcal{U}$,
$\big|\big\{\langle[b_{1}]_{\mathcal{U}},...,[b_{\mathrm{ar}(\mathrm{F})}]_{\mathcal{U}}\rangle\in(A^{I}/\mathcal{U})^{\mathrm{ar}(\mathrm{F})}\mid\{j\in I\mid\langle b_{1}(j),...,b_{\mathrm{ar}(\mathrm{F})}(j)\rangle\in S^{\theta(a)(j)\downarrow}_{\mathrm{F}}\}=X\big\}\big|$ $\leq$
[…]
so that
[…]
and likewise for $S^{a\uparrow}_{\mathrm{F}}$. Thus,
$\tau_{\mathcal{A}^{I}/\mathcal{U}}=\mathrm{max}\big\{\sum_{\mathrm{F}\in\sigma}\big(\prod_{j\in Z}|S^{\theta(a)(j)\downarrow}_{\mathrm{F}}|+\prod_{j\in Z}|S^{\theta(a)(j)\uparrow}_{\mathrm{F}}|\big)\big\}_{a\in A^{I}/\mathcal{U}}$,
and by the arbitrariness of $\theta$, we may conclude
$\tau_{\mathcal{A}^{I}/\mathcal{U}}=\mathrm{max}\big\{\sum_{\mathrm{F}\in\sigma}\big(\prod_{j\in Z}|S^{a(j)\downarrow}_{\mathrm{F}}|+\prod_{j\in Z}|S^{a(j)\uparrow}_{\mathrm{F}}|\big)\big\}_{a\in A^{I}}$.
Therefore, similarly to the proof of (a),
$\tau_{\mathcal{A}^{I}/\mathcal{U}}$ == $\mathrm{max}\big\{\sum_{\mathrm{F}\in\sigma}\big(\prod_{j\in Z}|S^{a(j)\downarrow}_{\mathrm{F}}|+\prod_{j\in Z}|S^{a(j)\uparrow}_{\mathrm{F}}|\big)\big\}_{a\in A^{I}}$
$\leq$ $\big(\sum_{\mathrm{F}\in\sigma}|S^{\mathrm{m}\downarrow}_{\mathrm{F}}|+|S^{\mathrm{m}\uparrow}_{\mathrm{F}}|\big)^{|Z|}$
== $(\tau_{\mathcal{A}})^{|Z|}$.
We therefore may immediately see that:
Proposition 4.33.
If $\mathcal{A}$ is a $\sigma$-structure whose relations are empty and $\mathcal{U}$ is a uniform ultrafilter, then $\tau_{\mathcal{A}^{I}/\mathcal{U}}=\tau_{\mathcal{A}^{I}}$.
Proof.
By Proposition 4.32, for $|Z|=\mathrm{min}\{|X|\}_{X\in\mathcal{U}}$, $\tau_{\mathcal{A}^{I}/\mathcal{U}}=(\tau_{\mathcal{A}})^{|Z|}=(\tau_{\mathcal{A}})^{|I|}=\tau_{\mathcal{A}^{I}}$. ∎
An interesting consequence of Theorem 4.31 is the following result:
Corollary 4.34.
Let $\mathcal{A}$ be a $\sigma$-structure whose relations are empty. If for any $\mathrm{F}\in\sigma$ and $a\in A$, $|S^{a\downarrow}_{\mathrm{F}}|,|S^{a\downarrow}_{\mathrm{F}}|\leq 1$, and for any $a\in A$ there are finitely many $\mathrm{F}\in\sigma$ such that either $|S^{a\downarrow}_{\mathrm{F}}|=1$ or $|S^{a\uparrow}_{\mathrm{F}}|=1$, then there is a choice function $e$ on the equivalence classes of $A^{I}/\mathcal{U}$ such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$.
Proof.
By Proposition 4.32, for a finite $\{\mathrm{F}_{i}\}_{i<n}\subseteq\sigma$,
$\mathrm{max}\big\{\sum_{\mathrm{F}\in\sigma}\big(\prod_{j\in Z}|S^{\theta(a)(j)\downarrow}_{\mathrm{F}}|+\prod_{j\in Z}|S^{\theta(a)(j)\uparrow}_{\mathrm{F}}|\big)\big\}_{a\in A^{I}/\mathcal{U}}\ \leq\ \sum_{i<n}1^{|Z|}+1^{|Z|}\ <\ \aleph_{0}$,
so since every ultrafilter is $\aleph_{0}$-complete, by we have our result. ∎
Let $\mathbb{N}_{\mathrm{PA}}$ be the naturals under the signature $\{0,\mathrm{s}\}$, for $\mathrm{s}$ the successor function, so that multiplication and addition may be later defined. The interest in the above result is that from it we obtain that $\mathbb{N}_{\mathrm{PA}}^{I}/\mathcal{U}\hookrightarrow\mathbb{N}_{\mathrm{PA}}^{I}$, meaning the unordered hypernaturals under that signature always embed into the respective direct power of the naturals.
Lemma 4.35.
Let $\mathcal{A}$ be a field. For any $a\in\mathcal{A}^{I}/\mathcal{U}$, $|S^{a\downarrow}_{+}|=|S^{a\uparrow}_{+}|=|S^{a\downarrow}_{\times}|=|S^{a\uparrow}_{\times}|=\big|A^{I}/\mathcal{U}\big|$, and so $\tau_{\mathcal{A}^{I}/\mathcal{U}}=4\times\big|A^{I}/\mathcal{U}\big|$ (so in case $A$ is infinite, or $I$ is infinite and $\mathcal{U}$ is free, $\tau_{\mathcal{A}^{I}/\mathcal{U}}=\big|A^{I}/\mathcal{U}\big|$).
Proof.
Just notice for any $a\in\mathcal{A}^{I}/\mathcal{U}$,
$|S^{a\downarrow}_{+}|$ == $|\{\langle x,y\rangle\mid x+y=a\}|$ == $\big|A^{I}/\mathcal{U}\big|$ == $|S^{a\downarrow}_{\times}|$, and
$|S^{a\uparrow}_{+}|$ == $|\{\langle x,y\rangle\mid x+a=y\}|$ == $\big|A^{I}/\mathcal{U}\big|$ == $|S^{a\uparrow}_{\times}|$,
so $\tau_{\mathcal{A}^{I}/\mathcal{U}}=|S^{a\downarrow}_{+}|+|S^{a\uparrow}_{+}|+|S^{a\downarrow}_{\times}|+|S^{a\uparrow}_{\times}|=4\times\big|A^{I}/\mathcal{U}\big|$. ∎
Proposition 4.36 ([3], p. 228).
The natural embedding between $\mathcal{A}$ and $\mathcal{A}^{I}/\mathcal{U}$ is an isomorphism iff $\mathcal{U}$ is $|A|^{+}$-complete.
Corollary 4.37.
Let $\mathcal{A}$ be a field such that $|A|\geq\aleph_{0}$. Then, there is a choice function $e$ on the equivalence classes such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff the natural embedding is an isomorphism.
Proof.
By Theorem 4.31, there is a choice function $e$ on the equivalence classes such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff $\mathcal{U}$ is $\tau_{\mathcal{A}^{I}/\mathcal{U}}^{+}$-complete, but since $|A|\geq\aleph_{0}$, by Lemma 4.35, $\tau_{\mathcal{A}^{I}/\mathcal{U}}=|A^{I}/\mathcal{U}|\geq|A|$. Thus, by Proposition 4.36, the natural embedding is an isomorphism. The converse is trivial, since the natural embedding is a choice function on the equivalence classes. ∎
Therefore, no infinite field has an ultrapower non-isomorphic to itself that embeds into its relative direct power. Particularly, for any non-principal $\mathcal{U}$ over $I$, $\mathbb{Q}^{I}/\mathcal{U}\not\hookrightarrow\mathbb{Q}^{I}$ and $\mathbb{R}^{I}/\mathcal{U}\not\hookrightarrow\mathbb{R}^{I}$.
It is well known that, for a totally ordered $\mathcal{A}$ and $\mathcal{U}$ a non-principal ultrafilter over $I$, $\mathcal{A}^{I}/\mathcal{U}$ has elements which are greater than any element in the natural embedding of $\mathcal{A}$ – that is, $\mathcal{A}^{I}/\mathcal{U}$ has elements greater than any of those in $\mathcal{A}$. Therefore, any such ultrapower cannot embed into its respective direct power, since there is no element in the generating structure which is strictly greater than all others, a condition necessary for the existence of such dominating element in the direct power. That means $\mathcal{A}^{I}/\mathcal{U}$ embeds into $\mathcal{A}^{I}$ iff the embedding is an isomorphism. In fact, this phenomenon may be generalised for any concurrent relation.
Definition 4.38 ($\kappa$-concurrent relation).
For a cardinal $\kappa$, a relation $\mathrm{R}\subseteq A\times B$ is $\kappa$-concurrent if for any subset $X\subseteq A$ with $|X|<\kappa$, if for every $a\in X$ there is $b$ such that $\mathrm{R}ab$, then there is $c\in B$ such that $\mathrm{R}ac$ for all $a\in X$.
In other words, $\mathrm{R}$ is $\kappa$-concurrent if its associated model is $\kappa$-saturated in the language induced solely by it. Therefore, the usual notion of concurrent relation coincides with $\aleph_{0}$-concurrency. We let $\kappa_{\mathrm{R}}=\mathrm{max}\{\kappa\mid\mathrm{R}\ \text{is}\ \kappa\text{-concurrent}\}$, that is, the greatest cardinal for which $\mathrm{R}$ is saturated.
Theorem 4.39.
Let $\mathcal{A}$ have a concurrent relation $\mathrm{R}^{\mathcal{A}}$ such that $\kappa_{\mathrm{R}^{\mathcal{A}}}=|A|$. If $\kappa_{\mathrm{R}^{\mathcal{A}}}$ is a limit, then, there is a choice function $e$ on the equivalence classes such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff the natural embedding is an isomorphism.
Proof.
By Theorem 4.31, the antecedent is the case iff $\mathcal{U}$ is $\tau_{\mathcal{A}^{I}/\mathcal{U}}^{+}$-complete. But since $\mathrm{R}^{\mathcal{A}}$ is $\kappa_{\mathrm{R}^{\mathcal{A}}}$-concurrent, for every $\zeta<\kappa_{\mathrm{R}^{\mathcal{A}}}$ there is $a_{\zeta}\in A$ such that $S^{a_{\zeta}}_{\mathrm{R}}=\zeta$. That means, since $\kappa_{\mathrm{R}^{\mathcal{A}}}$ is a limit, $\tau_{\mathcal{A}^{I}/\mathcal{U}}\geq\tau_{\mathcal{A}}\geq\mathrm{sup}\{S^{a_{\zeta}}_{\mathrm{R}}\}_{\zeta<\kappa_{\mathrm{R}^{\mathcal{A}}}}=\mathrm{sup}\{\zeta\}_{\zeta<\kappa_{\mathrm{R}^{\mathcal{A}}}}=\kappa_{\mathrm{R}^{\mathcal{A}}}=|A|$. Therefore $\mathcal{U}$ is $|A|^{+}$-complete, which by Proposition 4.36, is the case iff the natural embedding is an isomorphism. ∎
In case $\mathrm{R}$ is an order, the above result may be strengthened.
Theorem 4.40.
Let $\mathcal{A}$ be partially ordered. If either $\mathrm{coi}(\mathcal{A})=|A|$ or $\mathrm{cof}(\mathcal{A})=|A|$, there is a choice function $e$ on the equivalence classes such that $e:\mathcal{A}^{I}/\mathcal{U}\hookrightarrow\mathcal{A}^{I}$ iff the natural embedding is an isomorphism.
Proof.
Once again, by Theorem 4.31 the antecedent is the case iff $\mathcal{U}$ is $\tau_{\mathcal{A}^{I}/\mathcal{U}}^{+}$-complete. Since either $\mathrm{coi}(\mathcal{A})=|A|$ or $\mathrm{cof}(\mathcal{A})=|A|$, for an $a\in A$, either $|\{x\in A\mid x\leq^{\mathcal{A}}a\}|=|A|$ or $|\{x\in A\mid a\leq^{\mathcal{A}}x\}|\linebreak=|A|$. In either case, that means $\tau_{a}\geq|S^{a}_{\leq}|=|A|$, so $\tau_{\mathcal{A}^{I}/\mathcal{U}}\geq\tau_{\mathcal{A}}\geq\tau_{a}\geq|A|$. Therefore $\mathcal{U}$ is $|A|^{+}$-complete, which as before is the case iff the natural embedding is an isomorphism. ∎
The two above theorems offer the reason for which, in general, ultrapowers by non-principal ultrafilters of ordered structures, or structures with concurrent relations, do not embed into their related direct powers – for example, whereas Corollary 4.34 tells us a non-trivial ultrapower of the unordered naturals under the signature with solely the successor a function as a primitive symbol may embed into its relative direct power, Theorem 4.40 says the same structure with the addition of the usual total order, $\mathbb{N}_{\mathrm{PA},<}=\langle\mathbb{N}_{\mathrm{PA}},<\rangle$, is such that $\mathbb{N}_{\mathrm{PA},<}^{I}/\mathcal{U}\not\hookrightarrow\mathbb{N}_{\mathrm{PA},<}^{I}$ when $\mathcal{U}$ is free.
Theorem 4.41.
Let $\mathcal{B}$ be an ordered field. If $\mathcal{A}\hookrightarrow\mathcal{B}$ is a subfield which is dense in $\mathcal{B}$ and such that $|B|\leq|A|^{+}$, and $\mathcal{U}$ is a tails ultrafilter over $A$, then $\mathcal{B}\hookrightarrow\mathcal{A}^{A}/\mathcal{U}$.
Proof.
Let $\langle r_{\alpha}\rangle_{\alpha<|B|}$ be an enumeration of $B$. For simplicity, assume $r_{0}\in B\setminus A$. For each $r\in B$, let $L_{r}=\{j\in A\mid j<r\}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$, $L_{r}$ has no maximal element, and we may see $\mathrm{cof}(L_{r})\leq\mathrm{cof}(A)$, so we may build an order preserving $h_{r}:A\to L_{r}$ cofinal in its codomain. For each $\beta<|B|$, we employ the following construction:
  • (0th step) let:
    $B_{\alpha,0}$ == $\{r_{0},r_{0}^{-1}\}\cup A$,
    $B_{\alpha,n+1}$ == $B_{0,n}\cup\{x+y\mid x,y\in B_{0,n}\}\cup\{x\cdot y\mid x,y\in B_{0,n}\}$,
    $B_{\alpha}$ == $\bigcup_{i\in\omega}B_{0,i}$
    Let $e_{0,0}=\{h_{r_{0}},h_{r^{-1}_{0}}\}\cup\{h_{m}\mid m\in B\}$, where $h_{m}=\overline{m}$, $h_{r_{0}}:A\to L_{r_{0}}$ is some order preserving mapping such that $h_{r_{0}}(A)$ is cofinal in $L_{r_{0}}$, and $h_{r^{-1}_{0}}:A\to A$ is defined by:
    $h_{r^{-1}_{0}}=\begin{cases}\big(h_{r_{0}}(j)\big)^{-1},\ \text{if}\ \big(h_{r_{0}}(j)\big)^{-1}\neq 0\\ 0,\ \text{otherwise}\end{cases}$
    Let then
    $e_{0,n+1}$ == $e_{0,n}\cup\{h_{x+y}\mid x,y\in B_{0,n}\}\cup\{h_{x\cdot y}\mid x,y\in B_{0,n}\}$,
    where $h_{x+y}=h_{x}+_{1}h_{y}$ and $h_{x\cdot y}=h_{x}\cdot_{1}h_{y}$, for $_{1}$ and $\cdot_{1}$ being the pointwisely defined operations, and $e_{0}=\bigcup_{i\in\omega}e_{0,i}$; [^18] Notice that is well defined, for if $x\in B_{0,n}$, then $h_{x}\in e_{0,n}$; and $h_{x}\in A\times L_{x}\subseteq A^{A}$.
[…]
    $B_{\beta,0}$ == $\{r_{\beta},r_{\beta}^{-1}\}\cup(\bigcup_{\alpha<\beta}B_{\alpha})$,
    $B_{\beta,n+1}$ == $B_{\beta,n}\ \cup\ \{x+y\mid x,y\in B_{\beta,n}\}\cup\{x\cdot y\mid x,y\in B_{\beta,n}\}$,
    $B_{\beta}$ == $\bigcup_{i\in\omega}B_{\beta,i}$
    As before, $e_{\beta,0}=\{h_{r_{\beta}},h_{r_{\beta}^{-1}}\}\cup(\bigcup_{\alpha<\beta}e_{\alpha})$, where $h_{r_{\beta}}$ and $h_{r_{\beta}^{-1}}$ are defined similarly to the previous step, and
    $e_{\beta,n+1}$ == $e_{\beta,n}\cup\{h_{x+y}\mid x,y\in B_{\beta,n}\}\cup\{h_{x\cdot y}\mid x,y\in B_{\beta,n}\}$,
    where $h_{x+y}$ and $h_{x\cdot y}$ are as before. Let $e_{\beta}=\bigcup_{i\in\omega}e_{\beta,i}$.
Notice the construction is consistent in that for any $x\in B$, $h_{x}$ is only defined once: if $r_{\alpha^{+}}+y\in B_{\alpha}$ for some $y\in B_{\alpha}$, then $r_{\alpha^{+}}+y-y=r_{\alpha^{+}}\in B_{\alpha}$; if $r_{\alpha^{+}}\cdot y\in B_{\alpha}$, then $r_{\alpha^{+}}\cdot y\cdot y^{-1}\in B_{\alpha}$; and if $r_{\alpha^{+}}^{-1}\in B_{\alpha}$, then $r_{\alpha^{+}}\in B_{\alpha}$. Therefore, whenever $r_{\alpha^{+}}\not\in B_{\alpha}$, every element $x\in B_{\alpha^{+}}\setminus B_{\alpha}$ is such that $h_{x}\not\in e_{\alpha}$, and therefore $h_{x}$ is not defined in $e_{\alpha}$, so that it may be defined in $e_{\alpha^{+}}$. By the construction, we may see $r_{i}\in B_{\beta}$ iff $h_{r_{i}}\in e_{\beta}$, and for every $r\in B$, $h_{r}\in e_{|B|}$. We note $e:B\to\mathcal{A}^{A}/\mathcal{U};x\mapsto[h_{x}]$ is an embedding. To see it is injective, let $r,r^{\prime}\in B$ and $r\neq r^{\prime}$. If $r,r^{\prime}\in A$, then $h_{r}=\overline{r}$ and $\overline{r^{\prime}}=h_{r^{\prime}}$, which means $\{j\in A\mid h_{r}(j)=h_{r^{\prime}}(j)\}=A\in\mathcal{U}$, and thus $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Otherwise, without loss of generality, let $r\in B\setminus A$ and $r^{\prime}\in A$. We have either $r<r^{\prime}$ or $r<r^{\prime}$. In either case, by the definition of $h_{r^{\prime}}$, $\{j\in A\mid h_{r}(j)=\overline{r}(j)=r=h_{r^{\prime}}(j)\}\not\in\mathcal{U}$ and therefore $e(r)=[h_{r}]=[\overline{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Let now $r,r^{\prime}\in B\setminus A$. Once again, without loss of generality, let $r<r^{\prime}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$ there is $m\in A$ such that $r<m<r^{\prime}$. That means $m\in L_{r^{\prime}}$ but $m\not\in L_{r}$. Since $h_{r^{\prime}}(A)$ is cofinal in $L_{r^{\prime}}$, there is $m^{\prime}\in h_{r^{\prime}}(A)$ such that $m<m^{\prime}$, so since $h_{r^{\prime}}(A)\subseteq L_{r^{\prime}}$, $m^{\prime}\in L_{r^{\prime}}$, which means $r<m<m^{\prime}<r^{\prime}$. Let $t\in A$ be such that $h_{r^{\prime}}(t)=m^{\prime}$. Since $h_{r^{\prime}}$ is order-preserving, for any $t^{\prime}>t$, $h_{r^{\prime}}(t^{\prime})>h_{r}(t^{\prime})$. Thus, we have $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}$ is a tail of $\mathcal{A}$, and therefore $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}\subseteq\{j\in A\mid h_{r}(j)\neq h_{r^{\prime}}(j)\}\in\mathcal{U}$. Hence, $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Certainly, $e(0)=[0]$ and $e(1)=[1]$. Consider $r_{i},r_{j}\in B$. Then, $r_{i},r_{j}\in B_{max\{i,j\}}$, which means $h_{r_{i}},h_{r_{j}}\in e_{max\{i,j\}}$. By the definition of $e_{max\{i,j\}}$, we have $e(r_{i})+e(r_{j})=[h_{r_{i}}]+[h_{r_{j}}]=[h_{r_{i}}+_{1}h_{r_{j}}]=[h_{r_{i}+_{1}r_{j}}]=e(r_{i}+r_{j})$ (and likewise for multiplication). Let now $r_{i}<r_{j}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$, that means there is $a\in L_{r_{j}}$ such that for all $b\in L_{r_{i}}$, $b<a$. Since $h_{r_{j}}(A)$ is cofinal in $L_{r_{j}}$, there is $t\in A$ such that $h_{r_{j}}(t)>a$, and since $h_{r_{j}}$ is order preserving, for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{j}}(t)$. Since $h_{r_{i}}(A)\subseteq L_{r_{i}}$, that means for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{i}}(t^{\prime})$. Therefore, $\{j\in A\mid h_{r_{i}}(j)<h_{r_{j}}(j)\}$ is a tail of $\mathcal{A}$, and therefore in $\mathcal{U}$, which means $e(r_{i})=[h_{r_{i}}]<_{\mathcal{U}}[h_{r_{j}}]=e(r_{j})$. ∎
Let $\mathbb{Q}^{*}$ be some construction of the hyperrationals, and $\infty$ and $\varepsilon$ be its infinite and infinitesimal elements, respectively. It is well known the quotient ring $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$ is isomorphic to $\mathbb{R}$. However, the above result provides a direct embedding of $\mathbb{R}$ into the construction of $\mathbb{Q}^{*}$ in the present framework:
Corollary 4.42.
If $\mathcal{A}\hookrightarrow\mathbb{R}$ is an ordered subfield which is dense in $\mathbb{R}$, $|A|\geq\aleph_{0}$, and $\mathcal{U}$ is a tails ultrafilter over $A$, then $\mathbb{R}\hookrightarrow\mathcal{A}^{A}/\mathcal{U}$. Particularly, that means:
  • (a)
    $\mathbb{R}\hookrightarrow\mathbb{Q}^{\mathbb{Q}}/\mathcal{U}\cong\mathbb{F}_{1}(\mathbb{Q})$;
  • (b)
    $\mathbb{R}\preceq\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}\cong\mathbb{F}_{1}(\overline{\mathbb{Q}})$. [^19] Given $I_{0}=\mathbb{Q}$ (respectively, $I_{0}=\overline{\mathbb{Q}}$), and a adequate choice of ultrafilter defining $\sim_{1}$.
That is, the hyperrational field $\mathbb{Q}^{\mathbb{Q}}/\mathcal{U}$, although not being real closed, contains an isomorphic copy of the reals; and the hyperalgebraic field $\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}$ is a real closed non-Archimidean field extension of the reals. Notice, by the construction in Theorem 4.41, the relevant embeddings are not unique – as should be expected by the usual construction $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$.
5. An application to real closed fields
For an ordered set $\mathcal{A}$ and $f,g\in A^{A}$, say $f$ dominates $g$ if there is $j\in A$ such that for all $k>j$, $f(k)>g(k)$. Given an order $\langle A,<\rangle$ and $L,R\subseteq A$, we write $L<R$, as usual, to denote that every element of $L$ is lesser than every element of $R$. Given $a\in A$, for simplicity we write $a<L$ for $\{a\}<L$. Call $B\subseteq A^{A}$ unbounded if there is no $h\in A^{A}$ such that $h$ dominates every element of $B$. Following [18], define $\mathfrak{b}_{A}$ to be the size of the smallest unbounded subset of $A^{A}$. [12] has shown that $\aleph_{0}<\mathfrak{b}_{\omega}$. We adapt here that proof to the more general case:
Lemma 5.1.
Let $B\subseteq A^{A}$ be unbounded and $\{B_{\tau}\}_{\tau<\mathrm{cof}(\mathcal{A})}$ be a decomposition of $B$. Then, for some $\tau$, $B_{\tau}$ is unbounded.
Proof.
Suppose the latter is false, so that for any $\kappa<\mathrm{cof}(\mathcal{A})$ there is $f_{\kappa}\in A^{A}$ such that for all $g\in B_{\kappa}$, $f_{\kappa}$ dominates $g$. Since each $\kappa<\mathrm{cof}(\mathcal{A})$, there is $l_{\kappa}\in A$ such that $l_{\kappa}>\{f_{\tau}(\kappa)\}_{\tau\leq\kappa}$. Consider $f:A\to A;x\mapsto l_{x}$. Then $f$ dominates every $f_{\kappa}$ for $\kappa<\mathrm{cof}(\mathcal{A})$, and therefore every element of $A^{A}$. That means $A^{A}$ is bounded. ∎
Since $A^{A}$ is unbounded iff $\mathcal{A}$ is unbounded, we obtain:
Corollary 5.2.
If $\mathcal{A}$ is unbounded, then $\mathrm{cof}(\mathcal{A})<\mathfrak{b}_{\mathcal{A}}$.
Definition 5.3 (Tails filter).
A filter $F$ over an ordered set $\langle I,\leq\rangle$ is a tails filter if it contains every tail of the order, that is, if for every $k\in I$, $\{j\in I\mid k\leq j\}\in F$.
By the definition of $\mathfrak{b}_{\mathcal{A}}$, we may see that:
Lemma 5.4.
If $\mathcal{A}$ is unbounded and $\mathcal{U}$ is a tails ultrafilter over $A$, then $\mathfrak{b}_{\mathcal{A}}\leq\mathrm{cof}(\mathcal{A}^{A}/\mathcal{U})$.
By the above lemma and Corollary 5.2, we have:
Corollary 5.5.
If $\mathcal{A}$ is unbounded and $\mathcal{U}$ is a tails ultrafilter over $A$, then $\mathrm{cof}(\mathcal{A})<\mathrm{cof}(\mathcal{A}^{A}/\mathcal{U})$. Therefore, if also $|A|=\mathrm{cof}(\mathcal{A})$, then $|A|<|\mathcal{A}^{A}/\mathcal{U}|$.
For $X\subseteq A$, let $[X]$ denote the image of the natural embedding of $X$ into the ultrapower $\mathcal{A}^{I}/\mathcal{U}$. Say $\mathcal{A}^{I}/\mathcal{U}$ dominates $[X]$ if there is $a\in A^{I}$ such that $a]>[\overline{x}]$ for all $x\in X$ – that is, if there is an upper bound of $[X]$ in $\mathcal{A}^{I}/\mathcal{U}$.
Lemma 5.6.
Let $\mathcal{A}$ be a totally ordered set. Then $\mathcal{A}^{I}/\mathcal{U}$ dominates $[A]$ iff there is a total order $\langle I,<^{\prime}\rangle$ with the same cofinality as $\mathcal{A}$ over which $\mathcal{U}$ is a tails ultrafilter.
Proof.
($\Leftarrow$) Just let $X\subseteq A$ be cofinal with $A$ and consider an order preserving $f:I\to X$. Then $\overline{n}]<[f]$ for any $n\in A$.
($\Rightarrow$) Let $f\in A^{X}$ with $f]>[\overline{x}]$ for all $x\in A$. Equivalently, for all such $x$, $\{j\in A\mid x<f(j)\}\in\mathcal{U}$. Then, it is straightforward to see $I$ must be of the same cardinality as the cofinality of $\mathcal{A}$. Notice if $m,n\in A$ and $m<n$, then
$\{j\in A\mid n<f(j)\}\subset\{j\in A\mid m<f(j)\}$.
Then, $\big\{\{j\in A\mid n<f(j)\}\big\}_{n\in A}$ can be linearly ordered by supersets. Define
$l<^{\prime}k$ iff
$\big\{n\in A\mid k\in\{j\in A\mid n<f(j)\}\big\}\subset\big\{n\in A\mid l\in\{j\in A\mid n<f(j)\}\big\}$
Then $\langle I,<^{\prime}\rangle$ is a total order with the same cofinality as $\mathcal{A}$, and one might easily check $\mathcal{U}$ is a tails ultrafilter over it. ∎
Definition 5.7 (Saturation).
A structure $\mathcal{A}$ is $\kappa$-saturated if for any $X\subseteq A$ with $X<\kappa$, $\mathcal{A}$ realizes any complete type over XX. [^20] A type over $X\subseteq A$ is a set $\Phi$ of formulas with $n$ free variables and parameters in $X$ such that any finite subset of $\Phi$ is satisfiable in $\mathcal{A}$. A type is realized by $\mathcal{A}$ if there is a $n$-tuple in $A^{n}$ satisfying $\Phi$, and complete if for any $\varphi$ with parameters in $X$, either $\varphi\in\Phi$ or $\neg\varphi\in\Phi$.
Due to Hausdorff [11], we also define the following:
Definition 5.8 ($\eta_{\beta}$-set).
For $\beta\in\mathbf{On}$, an ordered set $A$ is an $\eta_{\beta}$-set if for any $L,R\subseteq A$ with $|L|,|R|<\aleph_{\beta}$ and $L<R$ there is $a\in A$ such that $L<a<R$. Likewise, a $\eta_{\mathbf{On}}$-class is a proper class that is a $\eta_{\beta}$-class for every $\beta\in\mathbf{On}$. [^21] That is, such that for any sets $X,Y\subset\mathbf{No}$ such that $X<Y$ there is $a\in\mathbf{No}$ such that $X<a<Y$.
Proposition 5.9 ([3], p. 343).
An unbounded densely ordered set is a $\eta_{\beta}$-set iff it is $\aleph_{\beta}$-saturated.
It is now straightforward to check that:
Lemma 5.10.
If $\mathcal{A}$ is an $\eta_{0}$-set, then $\mathcal{A}^{I}/\mathcal{U}$ is an $\eta_{0}$-set.
Proof.
By Proposition 5.9 and the properties of ultrapowers. ∎
Lemma 5.11.
Let $\mathcal{A}$ be an $\eta_{0}$-set. If $\mathcal{U}$ is a tails ultrafilter over a total order $\langle I,<^{\prime}\rangle$ with the same cofinality as $\mathcal{A}$, then for every $L,R\subseteq A$ such that $L<R$ there is $c\in\mathcal{A}^{I}/\mathcal{U}$ such that $[L]<c<[R]$.
Proof.
Without loss of generality, we may assume $(L,R)$ is either a gap or a cut, for otherwise by Lemma 5.10 we may take an element between the last element of $L$ and the first element of $R$. Suppose first the former. Let $\kappa_{L}$ and $\kappa_{R}$ be the cofinality and coinitiality of $L$ and $R$, respectively. Let first $\kappa_{L}\geq\kappa_{R}$. Then there is an order reversing mapping $v:L\to R$ with an image coinitial in its codomain. By the definition of $L$ and $R$, for any $l\in L$, $l<v(l)$, so by Lemma 5.10 $\big(l,v(l)\big)$ is a non-empty interval. That means we may take a choice function $\theta$ on the set of open intervals $\bigcup_{l\in L}\big\{\big(l,v(l)\big)\big\}$. Let $f:I\to\theta\big(\bigcup_{l\in L}\big\{\big(l,v(l)\big)\big\}\big)$ with an image cofinal in its codomain and such that if $a<^{\prime}b$ then $f(a)\leq f(b)$. Let now $c\in L$. Then, we have
$\overline{c}]<_{\mathcal{U}}[f]$ iff $\{j\in I\mid c<f(j)\}\in\mathcal{U}$.
Since the image of $f$ is cofinal in its codomain, there is $d\in I$ such that $f(d)\geq\theta\big(\big(c,v(c)\big)\big)$. Consider $k\in I$ such that $k>^{\prime}d$. Then, by our definitions, $f(k)\geq f(d)\geq\theta\big(\big(c,v(c)\big)\big)>c$. Since $\mathcal{U}$ is a tails ultrafilter on $\langle I,<^{\prime}\rangle$, by the arbitrariness of $k$ we conclude $\overline{c}]<_{\mathcal{U}}[f]$. By the arbitrariness of $c$, we may see $L]<_{\mathcal{U}}[f]$. By similar considerations, we may also see $f]<_{\mathcal{U}}[R]$. Let now $\kappa_{R}\geq\kappa_{L}$. We proceed analogously: we take an order preserving mapping $v:R\to L$ with an image cofinal in its codomain. Similarly to the previous case, for any $r\in R$, $\big(v(r),r\big)$ is a non-empty interval, so we may take a choice function $\theta$ on $\bigcup_{r\in R}\big\{\big(v(r),r\big)\big\}$, and let $f:I\to\theta\big(\bigcup_{r\in R}\big\{\big(v(r),r\big)\big\}\big)$ with an image cofinal in its codomain such that if $a<^{\prime}b$ then $f(a)\leq f(b)$. We then proceed with the same argument as the previous one, and arrive to the same conclusion. Suppose now $(L,R)$ is a cut. Then either $L$ has a last element or $R$ has a first element. Suppose the former, and let $M\in L$ be the last element. Similarly to the previous case, for any $r\in R$, by Lemma 5.10 $(M,r)$ is non-empty, so we take a choice function $\theta$ on the set $\bigcup_{r\in R}\{(M,r)\}$ and let $f:I\to\theta\big(\bigcup_{r\in R}\big\{\big(M,r\big)\big\}\big)$ have an image cofinal in its codomain, and be such that if $a<^{\prime}b$ then $f(a)\leq f(b)$. By the same argument as before, we obtain $L]<_{\mathcal{U}}[f]<_{\mathcal{U}}[R]$. If, on the other hand, $R$ has a first element $m$, an analogous argument may be employed by taking a choice function on $\bigcup_{l\in L}\{(l,m)\}$. ∎
Theorem 5.12.
Let $\lambda$ be a limit ordinal, $\mathcal{A}$ be an $\eta_{0}$-set, and its ultrapower hierarchy be generated by taking $\{\Upsilon_{\alpha}\}_{\alpha\in\mathbf{On}}$ as the index sets. Let $\aleph_{\lambda}$ be strongly inaccessible. For each $\alpha<\lambda$, suppose there are $\beta,\gamma<\lambda$ such that $|\Upsilon_{\beta}(\mathcal{A})|>|\Upsilon_{\alpha}(\mathcal{A})|$ and $\Upsilon_{\gamma}(\mathcal{A})$ dominates $\Upsilon_{\alpha}(\mathcal{A})]$. Then, $\Upsilon_{\lambda}(\mathcal{A})$ is an $\eta_{\lambda}$-set (equivalently, $\aleph_{\lambda}$-saturated).
Proof.
By the first supposition, $\aleph_{\lambda}\leq|\Upsilon_{\lambda}(\mathcal{A})|$. Let $L,R\subseteq A_{\lambda}$ with $|L|,|R|<\aleph_{\lambda}$ and $L<_{\lambda}R$. Since $\aleph_{\lambda}$ is strongly inaccessible, there is some $\beta<\lambda$ such that $L,R\subseteq\Upsilon_{\beta}(\mathcal{A})$. By the second supposition and Lemma 5.6, there is $\gamma>\beta$ such that $\mathcal{U}_{\gamma}$ is a tails ultrafilter over the totally ordered set $\langle\Upsilon_{\gamma}(\mathcal{A}),<^{\prime}\rangle$ with the same cofinality as $\Upsilon_{\gamma}(\mathcal{A})$. By Lemma 5.11, there is $c\in\Upsilon_{\gamma^{+}}(\mathcal{A})$ such that $L<_{\gamma^{+}}c<_{\gamma^{+}}R$. Once again, $\lambda$ is a limit, so that $\gamma^{+}<\lambda$, and therefore $L<_{\lambda}c<_{\lambda}R$. ∎
In other words, starting with an $\eta_{0}$-set $\mathcal{A}$, if there are enough increases in cardinality throughout the ultrapower hierarchy up to a limit $\lambda$, and if one may always saturate the order of a given level at a subsequent level that is lower than $\lambda$, and given $\aleph_{\lambda}$ is regular, then $\Upsilon_{\lambda}(\mathcal{A})$ is saturated in every cardinality lower than $\aleph_{\lambda}$. Notice, of course, if $|A|=\aleph_{\lambda}$, then the increases in cardinality are needless, and the theorem follows without that assumption. We therefore might see that:
Theorem 5.13.
Let $\mathcal{A}$ be an $\eta_{0}$-set, and for $\beta<\lambda$, let each successor stage $\beta^{+}$ of an ultrapower hierarchy be constructed by taking $\Upsilon_{\beta}(\mathcal{A})$ as the index set and a tails ultrafilter over it. If $\aleph_{\lambda}$ is strongly inaccessible, then, $\mathcal{A}_{\lambda}$ is an $\eta_{\lambda}$-set.
Proof.
Since each $\Upsilon_{\beta^{+}}(\mathcal{A})=\Upsilon_{\beta}(\mathcal{A})^{\Upsilon_{\beta}(\mathcal{A})}/\mathcal{U}_{\beta}$, where $\mathcal{U}_{\beta}$ is a tails ultrafilter over $\Upsilon_{\beta}(\mathcal{A})$, by Corollary 5.5, $\mathrm{cof}\big(\Upsilon_{\beta}(\mathcal{A})\big)<\mathrm{cof}\big(\Upsilon_{\beta^{+}}(\mathcal{A})\big)$, and so $|\Upsilon_{\lambda}(\mathcal{A})|\geq\mathrm{cof}\big(\Upsilon_{\lambda}(\mathcal{A})\big)=\mathrm{sup}\{\mathrm{cof}\big(\Upsilon_{\alpha}(\mathcal{A})\big)\mid\alpha<\lambda\}\geq\aleph_{\lambda}$. If we now let $L,R\subseteq\Upsilon_{\lambda}(\mathcal{A})$ with $L<_{\lambda}R$ and $|L|,|R|<\aleph_{\lambda}$, by Lemma 5.11 (using the same argument we have for Theorem 5.12), we may see $\Upsilon_{\lambda}(\mathcal{A})$ is an $\eta_{\lambda}$-set. ∎
Notice particularly that if $|A|=\aleph_{0}$, under the assumption of GCH, $|\Upsilon_{\lambda}(\mathcal{A})|=\aleph_{\lambda}$, and is therefore unique up to order isomorphism ([11], pp. 180–185). [^22] One might see that by induction, since $|A_{0}|=|A|=\aleph_{0}<\mathrm{cof}(A_{1})\leq|A^{A}|=2^{\aleph_{0}}=\aleph_{1}$.
The above results imply:
Corollary 5.14.
Let $\mathcal{A}$ be an $\eta_{0}$-set and each $\Upsilon_{\beta^{+}}(\mathcal{A})$ be constructed by taking $\Upsilon_{\beta}(\mathcal{A})$ as the index set and a tails ultrafilter over it. Then, $\Upsilon_{\mathbf{On}}(\mathcal{A})$ is an $\eta_{\mathbf{On}}$-class (equivalently, saturated in every cardinal).
Given Theorem 4.25, the above results implies that if $\mathcal{A}$ is an $\eta_{0}$-set, $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\in\mathbf{On}}$ is the family of index sets generating the cumulative power hierarchy, and each ultrafilter $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ defining $\sim_{\alpha^{+}}$ is a tails ultrafilter, then for each strongly inaccessible $\aleph_{\lambda}$, $\mathbb{F}_{\lambda}(\mathcal{A})$ is an $\eta_{\lambda}$-set, and $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ is an $\eta_{\mathbf{On}}$-class. However, we may eliminate the dependence on ultrapowers of that sort of construction by showing that, for any ultrapower hierarchy constructed by taking each successor level as an ultrapower having the preceding level as its index set – as constructed above –, there is a corresponding hierarchy of cumulative powers constructed by taking each preceding level as the index set of its successor such that, once quotiented by the equivalence relations $\sim$ defined by appropriate choices of ultrafilter, its structures are isomorphic to their ultrapower counterparts.
Lemma 5.15.
Let $\equiv$ be an equivalence relation on a set $A$. Then:
  • (a)
    every ultrafilter $\mathcal{U}$ over $A$ induces an ultrafilter $\mathcal{U}^{\prime}$ over $A/\!\equiv$ such that $\mathcal{U}^{\prime}\leq_{\mathrm{RK}}\mathcal{U}$;
  • (b)
    every ultrafilter $\mathcal{U}$ over $A/\!\equiv$ induces a filter base $F$ over $A$ such that $\{x\in A\mid[x]\in X\}\in F$ iff $X\in\mathcal{U}$.
Proof.
(a) Let $\mathcal{U}^{\prime}=\big\{\{[x]\mid x\in X\}\mid X\in\mathcal{U}\big\}$. Since $A\in\mathcal{U}$, we may see $A/\!\equiv\ \in\mathcal{U}^{\prime}$. Likewise, since $\varnothing\not\in\mathcal{U}$, we have $\varnothing\not\in\mathcal{U}^{\prime}$. Let $Z,T\in\mathcal{U}^{\prime}$. Then, $Z=\{[x]\mid x\in X\}$ and $T=\{[x]\mid x\in Y\}$ for some $X,Y\in\mathcal{U}$, which means $X\cap Y\in\mathcal{U}$, which in turn means $\{[x]\mid x\in X\cap Y\}=\linebreak\{[x]\mid x\in X\}\cap\{[x]\mid x\in Y\}\in\mathcal{U}^{\prime}$. Suppose now $Z\in\mathcal{U}^{\prime}$ and $Z\subseteq T\subseteq A/\!\equiv$. Once again, $Z$ must be of the previously described form. Let $\theta:(A/\!\equiv)\to A$ be a choice function. Then $X\subseteq X\cup\theta[T]$, so that $X\cup\theta[T]\in\mathcal{U}$, and therefore, $\{[x]\mid x\in X\cup\theta[T]\}=\{[x]\mid x\in X\}\cup\{[x]\mid x\in\theta[T]\}=Z\cup T=T\in\mathcal{U}^{\prime}$. At last, let $Z\not\in\mathcal{U}^{\prime}$. Then, for any $X$ such that $\{[x]\mid x\in X\}=Z$, $X\not\in\mathcal{U}$. Particularly, that means $\{x\mid[x]\in Z\}\not\in\mathcal{U}$, and therefore $\{x\mid[x]\not\in Z\}\in\mathcal{U}^{\prime}$. But then $\big\{[x]\mid x\in\{y\mid[y]\not\in Z\}\big\}=(A/\!\equiv)\setminus Z\in\mathcal{U}^{\prime}$. Consider now $h:A\to(A/\!\equiv);x\mapsto[x]$, and let $Z\in\mathcal{U}^{\prime}$. Since $Z=\{[x]\mid x\in X\}$ for some $\in\mathcal{U}, XX$ can be partitioned as $\{X_{i}\}_{i<|Z|}$, where $a,b\in X_{i}$ iff $a\equiv b$. Let $a_{i}\in X_{i}$. Since $X_{i}\subseteq[a_{i}]$, $X=\bigcup_{i<|Z|}X_{i}\subseteq\bigcup_{i<|Z|}[a_{i}]=h^{-1}(Z)$, and therefore $h^{-1}(Z)\in\mathcal{U}$. For the other direction, if $h^{-1}(Z)\in\mathcal{U}$ for some $Z\subseteq(A/\!\equiv)$, then by the very definition of $\mathcal{U}^{\prime}$ and $h$, $Z\in\mathcal{U}^{\prime}$. Therefore, $\mathcal{U}^{\prime}\leq_{\mathrm{RK}}\mathcal{U}$.
(b) Let $F^{\prime}=\big\{\{x\in A\mid[x]\in X\}\mid X\in\mathcal{U}\big\}$. Since $A/\!\equiv\ \in\mathcal{U}$ and $\varnothing\not\in\mathcal{U}$, we have $A\in F$ and $\varnothing\not\in F$. Let now $X,Y\in F$. Then, $\{x\in A\mid[x]\in X\},\{x\in A\mid[x]\in Y\}\in\mathcal{U}$, which means $\{x\in A\mid[x]\in X\}\cap\{x\in A\mid[x]\in Y\}=\{x\in A\mid[x]\in X\cap Y\}\in\mathcal{U}$. Thus, $X\cap Y\in F$. That $F$ has the described property follows from its definition and the fact $\mathcal{U}$ is an ultrafilter. ∎
Theorem 5.16.
Let a ultrapower hierarchy $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$, be generated by $\mathcal{A}$ and the families of index sets $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha<\beta}$ and ultrafilters $\{\mathcal{U}_{\alpha}\}_{\alpha<\beta}$ respectively over them. Then, there is a family of ultrafilters $\{\mathcal{U}^{\prime}_{\alpha}\}_{\alpha<\beta}$ respectively over $\{\mathcal{F}_{\alpha}(A)\}_{\alpha<\beta}$ such that for each $\gamma\leq\beta$, if for all $\delta<\gamma$, $\sim_{\delta^{+}}$ is defined by $\mathcal{U}^{\prime}_{\delta}$, then $\mathbb{F}_{\gamma}(\mathcal{A})\cong\Upsilon_{\gamma}(\mathcal{A})$.
Proof.
By induction on $\gamma$. The base case is trivial (just take the same ultrafilter), and the limit case is covered by Lemma 4.24. For the successor case, let $e:\Upsilon_{\alpha}(\mathcal{A})\to\mathbb{F}_{\alpha}(\mathcal{A})$ be the isomorphism, and $\mathcal{U}^{\prime}=\linebreak\{e[X]\mid X\in\mathcal{U}_{\alpha}\}$. One might check $\mathcal{U}^{\prime}$ is an ultrafilter over $\mathbb{F}_{\alpha}(\mathcal{A})$. Consider then $F=\big\{\{x\in\mathcal{F}_{\alpha}(A)\mid[x]\in X\}\mid X\in\mathcal{U}^{\prime}\big\}$. By Lemma 5.15, $F$ is a filter base over $\mathcal{F}_{\alpha}(A)$, and $\{x\in\mathcal{F}_{\alpha}(A)\mid[x]\in X\}\in F$ iff $X\in\mathcal{U}^{\prime}$. Let $\theta:\mathbb{F}_{\alpha}(\mathcal{A})\to\mathcal{F}_{\alpha}(A)$ be a choice function. By the definition of $F$, $\mathcal{F}_{\alpha}(A)\setminus\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\not\in F$, and we may also see $\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\cap X\neq\varnothing$ for any $X\in F$. Thus, $F^{\prime}=F\cup\{\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\}$ is also a filter base. Let $\mathcal{U}^{\prime}_{\alpha}\supseteq F^{\prime}$ be an ultrafilter and let it define $\sim_{\alpha^{+}}$. Consider $w:\Upsilon_{\alpha}(\mathcal{A})^{\Upsilon_{\alpha}(\mathcal{A})}\to\mathcal{F}^{*}_{\alpha^{+}}(A)$ such that for each $f\in\Upsilon_{\alpha}(\mathcal{A})^{\Upsilon_{\alpha}(\mathcal{A})}$, $w(f):x\mapsto\theta\big(e\big(f(e^{-1}([x]_{\alpha}))\big)\big)$. For clarity, let us abbreviate $e\big(f(e^{-1}([j]_{\alpha}))\big)$ as $f^{e}(j)$. We now show that $e^{\prime}:\Upsilon_{\alpha^{+}}(\mathcal{A})\to\mathbb{F}_{\alpha^{+}}(\mathcal{A});[x]_{\mathcal{U}_{\alpha}}\mapsto[w(x)]_{\alpha^{+}}$ is an isomorphism. It is trivial to see that $e^{\prime}$ preserves the interpretations of constants, that is, that for any constant $\mathrm{c}$, $e^{\prime}(\mathrm{c}^{\Upsilon_{\alpha^{+}}(\mathcal{A})})=e^{\prime}([\overline{\mathrm{c}^{\Upsilon_{\alpha}(\mathcal{A})}}]_{\mathcal{U}_{\alpha}})=[w(\overline{\mathrm{c}^{\Upsilon_{\alpha}(\mathcal{A})}})]_{\alpha^{+}}=[\overline{\mathrm{c}^{\mathbb{F}_{\alpha}(\mathcal{A})}}]_{\alpha^{+}}=\mathrm{c}^{\mathbb{F}_{\alpha^{+}}(\mathcal{A})}$. Let $e^{\prime}([f]_{\mathcal{U}_{\alpha}})=[w(f)]_{\alpha^{+}}=[w(g)]_{\alpha^{+}}=e^{\prime}([g]_{\mathcal{U}_{\alpha}})$, meaning $\{j\in\mathcal{F}_{\alpha}(A)\mid w(f)(j)\sim_{\alpha}w(g)(j)\}\in\mathcal{U}^{\prime}_{\alpha}$. Since $\mathcal{U}^{\prime}_{\alpha}$ is closed upwards, we have
$\{j\in\mathcal{F}_{\alpha}(A)\mid\exists k\big(j\sim_{\alpha}k\ \&\ w(f)(k)\sim_{\alpha}w(g)(k)\big)\}\in\mathcal{U}^{\prime}_{\alpha}$.
By the definition of $\mathcal{U}^{\prime}_{\alpha}$, one might check that is the case only if
$\{j\in\mathcal{F}_{\alpha}(A)\mid\exists k\big(j\sim_{\alpha}k\ \&\ w(f)(k)\sim_{\alpha}w(g)(k)\big)\}\in F$.
Thus, by the definition of $F$, we have
$\{[j]_{\alpha}\in\mathbb{F}_{\alpha}(A)\mid w(f)(j)\sim_{\alpha}w(g)(j)\}\in\mathcal{U}^{\prime}$.
Since $w(f)(j)=\theta\big(f^{e}(j)\big)\sim_{\alpha}\theta\big(g^{e}(j)\big)=w(g)(j)$ iff $f^{e}(j)=e\big(f(e^{-1}([j]_{\alpha}))\big)=e\big(g(e^{-1}([j]_{\alpha}))\big)=g^{e}(j)$, we have
$\{[j]_{\alpha}\in\mathbb{F}_{\alpha}(A)\mid f^{e}(j)=g^{e}(j)\}\in\mathcal{U}^{\prime}$.
By the definition of $\mathcal{U}^{\prime}$, that is the case only if
$\{e^{-1}([j]_{\alpha})\in\Upsilon_{\alpha}(\mathcal{A})\mid f^{e}(j)=g^{e}(j)\}\in\mathcal{U}_{\alpha}$,
that is,
$\{e^{-1}([j]_{\alpha})\in\Upsilon_{\alpha}(\mathcal{A})\mid e\big(f(e^{-1}([j]_{\alpha}))\big)=e\big(g(e^{-1}([j]_{\alpha}))\big)\}\in\mathcal{U}_{\alpha}$.
Since $e$ is an isomorphism, its inverse is surjective, and therefore we may see
$\{j\in\Upsilon_{\alpha}(\mathcal{A})\mid e\big(f(j)\big)=e\big(g(j)\big)\}\in\mathcal{U}_{\alpha}$.
But since $e$ is injective, $e\big(f(j)\big)=e\big(g(j)\big)$ iff $f(j)=g(j)$, so that
$\{j\in\Upsilon_{\alpha}(\mathcal{A})\mid f(j)=g(j)\}\in\mathcal{U}_{\alpha}$,
and therefore $f]_{\mathcal{U}_{\alpha}}=[g]_{\mathcal{U}_{\alpha}}$, so $e^{\prime}$ is injective. Let now $\mathrm{R}^{\Upsilon_{\alpha^{+}}(\mathcal{A})}[f_{0}]_{\mathcal{U}_{\alpha}}...[f_{n}]_{\mathcal{U}_{\alpha}}$. That is the case iff
$\{j\in\Upsilon_{\alpha}(\mathcal{A})\mid\mathrm{R}^{\Upsilon_{\alpha}(\mathcal{A})}f_{0}(j)...f_{n}(j)\}\in\mathcal{U}_{\alpha}$, iff
$\{e(j)\in\mathbb{F}_{\alpha}(\mathcal{A})\mid\mathrm{R}^{\Upsilon_{\alpha}(\mathcal{A})}f_{0}(j)...f_{n}(j)\}\in\mathcal{U}^{\prime}$.
Since $e$ is surjective, $j=e^{-1}([k]_{\alpha})$ for some $k\in\mathcal{F}_{\alpha}(A)$, so that is the same as
$\{[k]_{\alpha}\in\mathbb{F}_{\alpha}(\mathcal{A})\mid\mathrm{R}^{\Upsilon_{\alpha}(\mathcal{A})}f_{0}\big(e^{-1}([k]_{\alpha})\big)...f_{n}\big(e^{-1}([k]_{\alpha})\big)\}\in\mathcal{U}^{\prime}$.
Since $e$ is an isomorphism, that is the same as
$\{[k]_{\alpha}\in\mathbb{F}_{\alpha}(\mathcal{A})\mid\mathrm{R}^{\sim}_{\alpha}f_{0}^{e}(k)...f_{n}^{e}(k)\}\in\mathcal{U}^{\prime}$.
By the definition of $F$ and $\mathcal{U}^{\prime}_{\alpha}$, that is the case iff
$\{k\in\mathcal{F}_{\alpha}(A)\mid\exists j\in\mathcal{F}_{\alpha}(A)\big(k\sim_{\alpha}j\ \&\ \mathrm{R}^{\sim}_{\alpha}f_{0}^{e}(j)...f_{n}^{e}(j)\big)\}\in\mathcal{U}^{\prime}_{\alpha}$.
Since $f_{i}^{e}(k)=[\theta\big(f_{i}^{e}(k)\big)]_{\alpha}=[w(f_{i})(k)]_{\alpha}$, that is the same as
$\{k\in\mathcal{F}_{\alpha}(A)\mid\exists j\in\mathcal{F}_{\alpha}(A)\big(k\sim_{\alpha}j\ \&\ \mathrm{R}^{\sim}_{\alpha}[w(f_{0})(j)]_{\alpha}...[w(f_{n})(j)]_{\alpha}\big)\}\in\mathcal{U}^{\prime}_{\alpha}$.
However, if $k\sim_{\alpha}j$, then $e^{-1}([k]_{\alpha})=e^{-1}([j]_{\alpha})$. Thus, by the definition of $w$, for any $f\in\Upsilon_{\alpha}(\mathcal{A})^{\Upsilon_{\alpha}(\mathcal{A})}$, if $k\sim_{\alpha}j$, then $f^{e}(k)=f^{e}(j)$, and so $w(f)(k)=w(f)(j)$. Therefore, the above is the case iff
$\{k\in\mathcal{F}_{\alpha}(A)\mid\mathrm{R}^{\sim}_{\alpha}[w(f_{0})(k)]_{\alpha}...[w(f_{n})(k)]_{\alpha}\}\in\mathcal{U}^{\prime}_{\alpha}$,
so by Lemma 4.6, that is the case iff $\mathrm{R}^{\sim}_{\alpha^{+}}[w(f_{0})]_{\alpha^{+}}...[w(f_{n})]_{\alpha^{+}}$, that is, $\mathrm{R}^{\sim}_{\alpha^{+}}e([f_{0}]_{\mathcal{U}_{\alpha}})...e([f_{n}]_{\mathcal{U}_{\alpha}})$. By a similar argument, we may show $e^{\prime}$ preserves the functions. At last, to see $e^{\prime}$ is surjective, it suffices to show for any $f\in\mathcal{F}_{\alpha^{+}}(A)$ there is $f^{\prime}\in\Upsilon_{\alpha}(\mathcal{A})^{\Upsilon_{\alpha}(\mathcal{A})}$ such that $f\sim_{\alpha^{+}}w(f^{\prime})$. To see that, consider $f^{\prime}$ such that for any $j\in\Upsilon_{\alpha}(\mathcal{A})$, $f^{\prime}(j)=e^{-1}\big([f\big(\theta(e(j))\big)]_{\alpha}\big)$. Then, for any $j\in\mathcal{F}_{\alpha}(A)$,
$w(f^{\prime})(j)=\theta\big(e\big(f^{\prime}(e^{-1}([j]_{\alpha}))\big)\big)=\theta\big(e\big(e^{-1}\big([f\big(\theta(e(e^{-1}([j]_{\alpha})))\big)]_{\alpha}\big)\big)\big)=\theta\big([f(\theta([j]_{\alpha}))]\big)\sim_{\alpha}f\big(\theta([j]_{\alpha})\big)$.
Therefore,
$\{j\in\mathcal{F}_{\alpha}\mid w(f^{\prime})(j)\sim_{\alpha}f(j)\}\supseteq\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\in\mathcal{U}^{\prime}_{\alpha}$,
so since $\mathcal{U}^{\prime}_{\alpha}$ is closed by supersets, $w(f^{\prime})\sim_{\alpha^{+}}f$. ∎
Notice that, under a suitable assumption, the proof of the above result essentially goes both ways – that is, to show a ultrapower hierarchy isomorphic to the hierarchy of $\mathbb{F}_{\beta}(\mathcal{A})$ may be constructed.
Theorem 5.17.
Let $\{\mathbb{F}_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ be generated by a family of ultrafilters $\{\mathcal{U}_{\alpha}\}_{\alpha<\beta}$ respectively over $\{\mathcal{F}_{\alpha}(A)\}_{\alpha<\beta}$. If for each $\alpha<\beta$ there is a choice function $\theta:\mathbb{F}_{\alpha}(\mathcal{A})\to\mathcal{F}_{\alpha}(A)$ such that $\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\in\mathcal{U}_{\alpha}$, then there is a family of ultrafilters $\{\mathcal{U}^{\prime}_{\alpha}\}_{\alpha<\beta}$ respectively over the index sets $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha<\beta}$ such that if an ultrapower hierarchy $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ is generated by those parameters, then for any $\gamma\leq\beta$, $\Upsilon_{\gamma}(\mathcal{A})\cong\mathbb{F}_{\gamma}(\mathcal{A})$.
Proof.
The same argument as for Theorem 5.16. ∎
The above result allows for a narrower condition for the construction of isomorphic ultrapowers.
Lemma 5.18.
Let $A$ be a set, $\equiv$ an equivalence relation over it, $\kappa$ be the size of the smallest equivalence class of $A/\!\equiv$, and $\mathcal{U}$ be an ultrafilter over $A$. If for every choice function $\theta:(A/\!\equiv)\to A$, $\theta[A/\!\equiv]\not\in\mathcal{U}$, then $\mathcal{U}$ is $\kappa$-incomplete.
Proof.
For each equivalence class $a\in A/\!\equiv$, let $\langle x^{a}_{i}\rangle_{i<|a|}$ be an enumeration of its elements. Then there are $\kappa$ choice functions $\theta_{\tau}$ such that $\theta_{\tau};a\mapsto x^{a}_{\tau}$. Since $\theta_{\tau}[A/\!\equiv]\not\in\mathcal{U}$, we have $A\setminus\theta_{\tau}[A/\!\equiv]\in\mathcal{U}$, but $\bigcap_{\tau<\kappa}(A\setminus\theta_{\tau}[A/\!\equiv])=\varnothing\not\in\mathcal{U}$. ∎
Proposition 5.19.
For any $a,b\in A^{I}$, $|[a]_{\mathcal{U}}|=|[b]_{\mathcal{U}}|$, that is, every equivalence class in $\mathcal{A}^{I}/\mathcal{U}$ is equally sized.
Proof.
If $|A|=1$ the result is trivial, so suppose $|A|>1$. We may build an injection from $a]_{\mathcal{U}}$ to $b]_{\mathcal{U}}$, and the other way around, in the following manner. Let $\langle c_{\alpha}\rangle_{\alpha<|[a]_{\mathcal{U}}|}$ be an enumeration of the members of that equivalence class such that $c_{0}=a$. Then $\{j\in I\mid a(j)=c_{\alpha}(j)\}\in\mathcal{U}$. For each $0<\beta<|[a]_{\mathcal{U}}|$, we inductively define $h_{\beta}\in A^{I}$ letting $h_{0}=b$, and for $\beta>0$, for any $k\in I$,
$h_{\beta}(k)=\begin{cases}b(k),\ \text{if}\ k\in\{j\in I\mid a(j)=c_{\beta}(j)\}\\ c_{\beta}(k),\ \text{otherwise}\end{cases}$
We now show $h:[a]_{\mathcal{U}}\to[b]_{\mathcal{U}};c_{\alpha}\mapsto h_{\alpha}$ is an injection: for $c_{\delta},c_{\gamma}\in[a]_{\mathcal{U}}$ and $\delta\neq\gamma$, clearly, $h_{\delta},h_{\gamma}\in[b]_{\mathcal{U}}$. Suppose $h_{\delta}=h_{\gamma}$. By construction, that means
$\{j\in I\mid c_{\delta}(j)\neq c_{\gamma}(j)\}\subseteq\{j\in I\mid a(j)=c_{\delta}(j)\}$,
(otherwise $c_{\delta}(t)=h_{\delta}(t)\neq h_{\gamma}(t)=c_{\gamma}(t)$ for some $t\not\in\{j\in I\mid a(j)=c_{\delta}(j)\}$) and likewise
$\{j\in I\mid c_{\delta}(j)\neq c_{\gamma}(j)\}\subseteq\{j\in I\mid a(j)=c_{\gamma}(j)\}$,
so that
$\{j\in I\mid c_{\delta}(j)\neq c_{\gamma}(j)\}\subseteq(\{j\in I\mid a(j)=c_{\delta}(j)\}\cap\{j\in I\mid a(j)=c_{\gamma}(j)\})$.
However,
$(\{j\in I\mid a(j)=c_{\delta}(j)\}\cap\{j\in I\mid a(j)=c_{\gamma}(j)\})\subseteq\{j\in I\mid c_{\delta}(j)=c_{\gamma}(j)\}$,
and thus
$\{j\in I\mid c_{\delta}(j)\neq c_{\gamma}(j)\}\subseteq\{j\in I\mid c_{\delta}(j)=c_{\gamma}(j)\}$.
But that means $\{j\in I\mid c_{\delta}(j)\neq c_{\gamma}(j)\}=\varnothing$, and therefore $c_{\gamma}=c_{\delta}$. Using an adequately modified construction we may build a similar injection in the other direction. ∎
Corollary 5.20.
Let $\{\mathbb{F}_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ be generated by a family of ultrafilters $\{\mathcal{U}_{\alpha}\}_{\alpha<\beta}$ respectively over $\{\mathcal{F}_{\alpha}(A)\}_{\alpha<\beta}$, and $\kappa_{\alpha}$ be the size of the equivalence classes of $\mathbb{F}_{\alpha}(\mathcal{A})$. If, for each $\alpha<\beta$, $\mathcal{U}_{\alpha}$ is $\kappa_{\alpha}$-complete, then there is a family of ultrafilters $\{\mathcal{U}^{\prime}_{\alpha}\}_{\alpha<\beta}$ respectively over the index sets $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha<\beta}$ such that if an ultrapower hierarchy $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ is generated by those parameters, then for any $\gamma\leq\beta$, $\Upsilon_{\gamma}(\mathcal{A})\cong\mathbb{F}_{\gamma}(\mathcal{A})$.
Proof.
Since $\mathcal{U}_{\alpha}$ is $\kappa_{\alpha}$-complete, by Lemma 5.18, there is a choice function $\theta:\mathbb{F}_{\alpha}(\mathcal{A})\to\mathcal{F}_{\alpha}(A)$ such that $\theta[\mathbb{F}_{\alpha}(\mathcal{A})]\in\mathcal{U}_{\alpha}$. The rest follows from Theorem 5.17. ∎
Therefore, the above result shows that if an ultrapower hierarchy of a structure is such that each successor level is constructed by having the preceding level as its index set, then there is a hierarchy of cumulative powers with each successor level also constructed by having the preceding level as its index set which is isomorphic to it. Likewise, if the cumulative power hierarchy generated with such choice of index sets quotiented by $\sim$ is such that the equivalence relation of each successor level is constructed by taking an ultrafilter $\mathcal{U}$ such that there is a choice function $\theta$ on the equivalence classes of $\mathbb{F}_{\beta}(\mathcal{A})$ with $\theta[\mathbb{F}_{\beta}(\mathcal{A})]\in\mathcal{U}$, then there is an ultrapower hierarchy isomorphic to it – that is, there is an equivalent construction of the so quotiented cumulative power hierarchy by the ultrapower hierarchy. The above results thus pose the following question:
  • Open Problem 2: Let $\{\mathbb{F}_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ be generated by a family of ultrafilters $\{\mathcal{U}_{\alpha}\}_{\alpha<\beta}$ respectively over $\{\mathcal{F}_{\alpha}(A)\}_{\alpha<\beta}$, and $\kappa_{\alpha}$ be the size of the equivalence classes of $\mathbb{F}_{\alpha}(\mathcal{A})$. Let also $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\leq\beta}$ be generated by a family of ultrafilters $\{\mathcal{U}^{\prime}_{\alpha}\}_{\alpha<\beta}$ respectively over the index sets $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha<\beta}$. Can there be $\gamma<\beta$ such that $\mathcal{U}^{\prime}_{\gamma}$ is $\kappa_{\gamma}$-incomplete, while $\mathbb{F}_{\beta}(\mathcal{A})\cong\Upsilon_{\beta}(\mathcal{A})$?
We conjecture that, in fact, such restriction on the ultrafilters is not necessary, and so that the similarities between both constructions is such that the quotiented $\mathcal{F}$-hierarchies and ultrapower hierarchies constructed by having the previous levels as the index sets of the successor levels are equivalent, so that any structure obtained by one of the constructions may be obtained by the other in a straightforward manner.
Now, given an ultrapower hierarchy, notice regardless of the choice of family of index sets and ultrafilters, by Łos’s theorem we immediately have:
Proposition 5.21.
For any $\alpha\leq\beta\in\mathbf{On}$, $\Upsilon_{\alpha}(\mathcal{A})\preceq\Upsilon_{\beta}(\mathcal{A})$, and therefore $\mathcal{A}\preceq\Upsilon_{\beta}(\mathcal{A})$.
Of course, the properties of each level of this hierarchy depend on the choice of generating structure and ultrafilter of each level. Consider the case we want saturation of the structures.
For a set $X$, let $\mathcal{P}^{\omega}(X)$ denote the set of cofinite subsets of $X$.
Definition 5.22 ($\kappa$-good ultrafilter).
Let $\mathcal{U}$ be an ultrafilter and $\kappa$ a cardinal. $\mathcal{U}$ is said to be $\kappa$-good if for every $\tau<\kappa$ and monotonic function $H:\mathcal{P}^{\omega}(\tau)\to\mathcal{U}$ there is a multiplicative function $H^{\prime}:\mathcal{P}^{\omega}(\tau)\to\mathcal{U}$ such that for any $x\in\mathcal{P}^{\omega}(\tau)$, $H^{\prime}(x)\subseteq H(x)$.
A filter is countably incomplete if it is not $\aleph_{1}$-complete. In the context of saturation, countably incomplete $\kappa$-good ultrafilters are of interest for the following result.
Proposition 5.23 ([15], p. 180).
Let $\mathcal{U}$ be a countably incomplete $\kappa^{+}$-good ultrafilter over $I$. Then, for any structure $\mathcal{A}$, $|\mathcal{A}^{I}/\mathcal{U}|=2^{\kappa}$ and $\mathcal{A}^{I}/\mathcal{U}$ is $\kappa^{+}$-saturated.
Chang and Keisler proved in [3] (p. 387), for any set $I$ of cardinality $\kappa\geq\aleph_{0}$, the existence of a countably incomplete $\kappa^{+}$-good ultrafilter over $I$ under the assumption of the Generalized Continuum Hypothesis (GCH). Later, Kunen improved the result by showing the existence of such an ultrafilter without GCH.
Proposition 5.24 ([16], p. 304).
For any $I$ with $|I|=\kappa\geq\aleph_{0}$ there is a countably incomplete $\kappa^{+}$-good ultrafilter over $I$.
Take now the hierarchy previously presented, and let $|A|\geq\aleph_{0}$. By Propositions 5.24 and Proposition 5.23, we may therefore obtain the following:
Corollary 5.25.
Let $|A|=\aleph_{\alpha}$, and for each $\alpha\in\mathbf{On}$ let $\Upsilon_{\alpha^{+}}(\mathcal{A})$ be constructed by taking a countably incomplete $|\Upsilon_{\alpha}(\mathcal{A})|^{+}$-good ultrafilter over $\Upsilon_{\alpha}(\mathcal{A})$. Then, for each $\beta>0$, $|\Upsilon_{\beta^{+}}(\mathcal{A})|=\beth_{\beta^{+}}(|A|)$ and $\Upsilon_{\beta^{+}}(\mathcal{A})$ is $\aleph_{\alpha+\beta^{+}}$-saturated, and for a limit $\lambda$, $|\Upsilon_{\lambda}(\mathcal{A})|=sup\{2^{|\Upsilon_{\alpha}(\mathcal{A})|}\mid\alpha<\lambda\}$ and $\Upsilon_{\lambda}(\mathcal{A})$ is $\aleph_{\alpha+\lambda}$-saturated. Particularly, if $A$ is countable, that means $\Upsilon_{\beta}(\mathcal{A})$ is $\aleph_{\beta}$-saturated and $|\Upsilon_{\beta}(\mathcal{A})|=\beth_{\beta}$.
Corollary 5.26.
Assume GCH. If $A$ is countable, for each $\beta>0$, $\mathcal{A}_{\beta}$ is $\aleph_{\beta}$-saturated and $|A_{\beta}|=\aleph_{\beta}$.
Once again, given Theorem 4.25, we have:
Corollary 5.27.
Let $|A|=\aleph_{\alpha}$. If each ultrafilter in $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ defining $\sim_{\alpha^{+}}$ is a tails ultrafilter respectively over $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\in\mathbf{On}}$, then for each $\beta>0$, $\mathbb{F}_{\beta}(\mathcal{A})$ is $\aleph_{\alpha+\beta^{+}}$-saturated, and for a limit $\lambda$, $|\mathbb{F}_{\lambda}(\mathcal{A})|=sup\{2^{|\mathbb{F}_{\alpha}(\mathcal{A})|}\mid\alpha<\lambda\}$ and $\mathbb{F}_{\lambda}(\mathcal{A})$ is $\aleph_{\alpha+\lambda}$-saturated.
More specifically, if $\mathcal{A}$ is a real closed field, $|A|=\aleph_{\alpha}$, and for each $\mathcal{U}_{\beta}$ over $\Upsilon_{\beta}(\mathcal{A})$ is countably incomplete and $|\Upsilon_{\beta}(\mathcal{A})|^{+}$-good, then each $\Upsilon_{\beta^{+}}(\mathcal{A})$ – and its correspondent $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ – is a $\aleph_{\alpha+\beta^{+}}$-saturated hyperreal field. Notice, however, that by Theorem 5.16 we may see there is a always a straightforward choice of ultrafilter over $\mathcal{F}_{\beta}(\mathcal{A})$ that makes $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ a hyperreal field; and furthermore, that by Theorem 5.13 and Corollary 5.14, it is straightforward to make $\mathbb{F}_{\lambda}(\mathcal{A})$ into a $\aleph_{\lambda}$-saturated hyperreal field, and likewise, $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ into a maximally saturated real closed field without the need of countably incomplete good ultrafilters. In fact, consider the following notion.
Definition 5.28 (Universally embedding).
An ordered field (or ordered abelian group, or ordered class) $\mathcal{A}$ is $\kappa$-universally embedding if for each subfield (or abelian subgroup, or subclass) $\mathcal{B}\subseteq\mathcal{A}$ and extension field (or abelian group, or class) $\mathcal{B}^{\prime}\supseteq\mathcal{B}$ such that $|B|,|B^{\prime}|<\kappa$, there is $\mathcal{B}^{\prime\prime}\subseteq\mathcal{A}$ such that $\mathcal{B}^{\prime}\cong\mathcal{B}^{\prime\prime}$ and the isomorphism is an extension of identity on $\mathcal{A}$.
Introduced by Conway [5], the surreal field $\mathbf{No}$ is an universally embedding real closed field that is a proper class. Besides the operations of multiplication and addition, the operations of division and exponentiation may be further defined [5], [6]. The following results make the assumption of Global Choice, which we shall denote by writing its acronym GC.
Proposition 5.29 ([5], p. 43).
(GC) $\mathbf{No}$ is the unique (up to isomorphism) $\aleph_{\mathbf{On}}$-universally embedding ordered field.
Alling similarly notes that $\mathbf{No}$ also generalizes $\eta_{\beta}$-sets in being an $\eta_{\mathbf{On}}$-class, and thus maximally saturated [1]. Given the saturation properties that may be obtained by either ultrapower or quotiented cumulative power hierarchies previously observed, we may therefore investigate the similarities between $\mathbf{No}$, $\Upsilon_{\mathbf{On}}(\mathcal{A})$, and $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$.
Proposition 5.30 ([7], p. 10).
(GC) For $\beta\in\mathbf{On}$, $\mathcal{A}$ is an $\aleph_{\beta}$-universally embedding ordered field (or ordered abelian group, or ordered class) iff $\mathcal{A}$ is a real-closed field (or ordered abelian group, or ordered class, respectively) that is a $\eta_{\beta}$-set. Any such structure is unique up to isomorphism.
Corollary 5.31.
(GC) If $\mathcal{A}$ is an $\eta_{0}$ real-closed field (or ordered abelian group, or ordered class) and for each $\alpha<\beta$, $\Upsilon_{\alpha^{+}}(\mathcal{A})$ is constructed by taking a countably incomplete $|\Upsilon_{\alpha}(\mathcal{A})|^{+}$-good ultrafilter over $\Upsilon_{\alpha}(\mathcal{A})$, then $\Upsilon_{\beta}(\mathcal{A})$ is the unique (up to isomorphism) $\aleph_{\beta}$-universally embedding field (or ordered abelian group, or ordered class, respectively).
Proof.
By Proposition 5.21, $\Upsilon_{\beta}(\mathcal{A})$ is a real-closed field. By Corollary 5.25, $\mathcal{A}_{\beta}$ is $\aleph_{\beta}$-saturate, so by Proposition 5.9, it is an $\eta_{\beta}$-set. By Proposition 5.30, $\Upsilon_{\beta}(\mathcal{A})$ is then a $\aleph_{\beta}$-universally embedding real-closed field. The argument is similar for the other cases. ∎
Given the above result, with Proposition 5.29 we therefore obtain the following broader result:
Corollary 5.32.
(GC) Let $\mathcal{A}$ be an $\eta_{0}$ real-closed field (or ordered abelian group, or ordered class) and for each $\beta\in\mathbf{On}$:
  • (1)
    let each $\Upsilon_{\beta^{+}}(\mathcal{A})$ be constructed by taking $\Upsilon_{\beta}(\mathcal{A})$ as the index set and a tails ultrafilter over it. Then:
    • (1.a)
      if $\aleph_{\lambda}$ is strongly inaccessible then $\Upsilon_{\lambda}(\mathcal{A})$ is the unique (up to isomorphism) $\aleph_{\lambda}$-universally embedding field (or ordered abelian group, or ordered class, respectively);
    • (1.b)
      $\Upsilon_{\mathbf{On}}(\mathcal{A})$ is the unique (up to isomorphism) $\aleph_{\mathbf{On}}$-universally embedding field (or ordered abelian group, or ordered class, respectively).
  • (2)
    let each $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ be constructed by taking an appropriate choice of ultrafilter over $\mathcal{F}_{\beta}(\mathcal{A})$. [^23] That is, such as constructed in Theorem 5.16. Then:
    • (2.a)
      if $\aleph_{\lambda}$ is strongly inaccessible then $\mathbb{F}_{\lambda}(\mathcal{A})$ is the unique (up to isomorphism) $\aleph_{\lambda}$-universally embedding field (or ordered abelian group, or ordered class, respectively);
    • (2.b)
      $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ is the unique (up to isomorphism) $\aleph_{\mathbf{On}}$-universally embedding field (or ordered abelian group, or ordered class, respectively).
That is, if $\mathcal{A}$ is an $\eta_{0}$ real-closed field, then there are simple choices of ultrafilters defining the ultrapower or quotiented cumulative power such that $\mathbb{F}_{\mathbf{On}}(\mathcal{A})\cong\mathbf{No}\cong\Upsilon_{\mathbf{On}}(\mathcal{A})$. Particularly, that means the ultrapower, or cumulative power, hierarchy generated by the reals or the algebraic numbers quotiented in the described manner is a surreal field.